### v1

# A Survey of Multimodal Large Language Model from A Data-centric Perspective

Abstract.
Human beings perceive the world through diverse senses such as sight, smell, hearing, and touch. Similarly, multimodal large language models (MLLMs) enhance the capabilities of traditional large language models by integrating and processing data from multiple modalities including text, vision, audio, video, and 3D environments. Data plays a pivotal role in the development and refinement of these models. In this survey, we comprehensively review the literature on MLLMs from a data-centric perspective. Specifically, we explore methods for preparing multimodal data during the pretraining and adaptation phases of MLLMs. Additionally, we analyze the evaluation methods for datasets and review benchmarks for evaluating MLLMs. Our survey also outlines potential future research directions. This work aims to provide researchers with a detailed understanding of the data-driven aspects of MLLMs, fostering further exploration and innovation in this field.
Multimodal Large Language Models; Data-centric Artificial Intelligence
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACM
1. introduction
In recent years, we have witnessed rapid advancement of large language models (LLMs) and multimodal large language models (MLLMs) (Zhao et al., 2023d; Wu et al., 2023b). MLLMs, such as GPT-4 (OpenAI, 2023), Flamingo (Alayrac et al., 2022), BLIP2 (Li et al., 2023f) and X-InstructBLIP (Panagopoulou et al., 2023), integrate multiple modality information, demonstrating impressive comprehension and generation capabilities. These models achieve competitive performance in traditional multimodal tasks, such as visual recognition (Zhang et al., 2024a), video understanding (Xu et al., 2021a; Tang et al., 2023), speech recognition (Min and Wang, 2023) and 3D understanding (Guo et al., 2023; Hong et al., 2023). Moreover, their excellent language understanding capacity enables strong performance in text-rich tasks, such as question answering (Hu et al., 2024), multi-dialog conversation and logical reasoning (Xu et al., 2024; Li et al., 2023l).
Most existing MLLMs focus on modifying model architecture to explore the use of information from multiple modality (Longpre et al., 2023; Xie et al., 2024a; Soboleva et al., 2023; Xie et al., 2024b; Kaddour, 2023; Ye et al., 2024). While model effectiveness is crucial, data also significantly impacts the success of MLLMs. For example, Hoffmann et al. (2022) demonstrate that in order to scale up models, it is necessary to increase the training data size. Beyond data quantity, data quality is equally important. Previous research (Sorscher et al., 2022) indicates that carefully curated datasets can enable smaller models to achieve comparable performance to larger ones. However, comprehensive studies on data curation and utilization for MLLMs are still lacking. Therefore, this study aims to provide a comprehensive understanding of MLLMs from a data-centric perspective.
In contrast to model-centric approaches that prioritize architectural enhancements while relying on fixed datasets, data-centric perspectives emphasize iterative improvements to datasets to enhance performance. Within the scope of data-centric MLLMs, our focus lies on leveraging the heterogeneous nature of data modalities, enhancing data structure, increasing data quantity, and elevating data quality to augment MLLMs (Zha et al., 2023). Our discussion answers three key questions from a data-centric perspective at different stages of MLLMs:
  • Q1: How to collect, select, and manage data for MLLMs? The substantial data quantity requirements and the heterogeneity of multimodal data pose challenges in gathering, selecting, and effectively managing data for model training. Different training stages of MLLMs also lead to varying data type requirements.
  • Q2: How data affects the performance of MLLMs? Understanding the relationship between data characteristics and the performance of MLLMs is crucial for optimizing datasets and enhancing model capabilities.
  • Q3: How to evaluate the data for MLLMs? It is necessary to develop comprehensive evaluation benchmarks to assess the performance and robustness of MLLMs across diverse tasks.
Difference between this survey and the existing one. Several existing surveys have focused on LLMs (Zhao et al., 2023d; Hadi et al., 2023; Naveed et al., 2023) and MLLMs (Wu et al., 2023b; Zhang et al., 2024c) from a model-centric perspective, but they lack an in-depth analysis of data-centric aspects. Recently, some surveys have started to focus on data preparation for LLMs, such as data management methods (Wang et al., 2023b), data selection methods (Albalak et al., 2024), and comprehensive reviews of LLM datasets (Liu et al., 2024a). However, these surveys primarily focus on data management and selection methods for text-only LLMs and do not provide a thorough analysis of the data processing pipeline for MLLMs. Although Zhang et al. (2024c) summarize datasets for MLLMs, they fall short of providing a comprehensive analysis of these datasets. The most closely related work to ours is data-centric artificial intelligence (DCAI) (Zha et al., 2023; Jarrahi et al., 2022; Jakubik et al., 2022; Polyzotis and Zaharia, 2021; Whang et al., 2023), which also focuses on data-centric views of AI research but does not specifically analyze LLMs and MLLMs.
With the rapid growth of MLLMs and the increasingly important role of data in this large model era, we believe it is crucial to provide a comprehensive overview of data-centric approaches for MLLMs. This survey aims to thoroughly review the literature on the advances of MLLMs from a data-centric perspective and discusses open issues or future directions in this field.
Contributions. In this survey, we reviewed the literature on the advances of MLLMs from data-centric perspectives. We provide researchers and developers with a general and comprehensive understanding of the latest developments in the field of MLLM data side. The key contributions of this survey are summarized as follows:
  • New data-centric perspective. We provide a comprehensive review of MLLMs from a data-centric perspective, considering modalities such as text, image, video, and audio.
  • Data preparation and management pipeline. We summarize the data preparation and management pipeline for MLLMs in both the pre-training and adaptation phases.
  • Data evaluation benchmarks. We outline commonly used evaluation benchmarks from a data-centric perspective.
  • Open issues and future directions. We discuss open issues in current research on data-centric LLMs and propose several future research directions.
Figure 1. Overview of the data pipeline for MLLMs.
[…]
[Overview of data pipeline for MLLMs]Overview of data pipeline for MLLMs
The remainder of this article is organized as follows: Section 2 introduces the preliminaries for LLMs and MLLMs, and discusses the motivations for analyzing them from a data-centric perspective. Sections 3 to 5 summarize the main stages of collecting, processing and selecting data for MLLMs training. Section 6 summarizes data evaluation methods and existing evaluation datasets for MLLMs according to evaluation tasks. Section 7 discusses open issues and highlights several future research directions in this field. Finally, we conclude the survey in Section 8. Our Github repository can be found in https://github.com/beccabai/Data-centric_multimodal_LLM.
2. backgrounds and categorization
2.1. Large Language Models
The field of natural language processing (NLP) has witnessed a remarkable evolution in language modeling techniques, culminating in the development of LLMs. The journey began with statistical methods (Jelinek, 1998; Gao and Lin, 2004; Rosenfeld, 2000), which laid the foundation for language modeling by capturing the probabilistic distributions of words and phrases. Subsequently, neural network approaches emerged (Bengio et al., 2000; Mikolov et al., 2010; Kombrink et al., 2011), leveraging the power of deep learning to learn more complex and abstract representations of language. However, such models were typically designed for specific tasks and encountered challenges related to scalability and generalization.
A significant breakthrough came with the introduction of pre-trained language models, which aimed to capture general language representations by training on vast amounts of unlabeled text data. Models such as BERT (Kenton and Toutanova, 2019) and GPT (Radford et al., [n. d.]) demonstrated the effectiveness of pre-training and fine-tuning for various NLP tasks. Building upon these advancements, LLMs have taken the concept of pre-training to an unprecedented scale, with models containing billions of parameters and trained on massive corpora spanning hundreds of gigabytes to terabytes of text data (Brown et al., 2020; Raffel et al., 2020)
LLMs exhibit several notable properties that distinguish them from previous language models. One fundamental characteristic is the scaling law, which describes how the performance of large language models improves as they scale in terms of model size, training data, and computational resources. These laws provide insights into the relationships between the number of parameters, the amount of training data, the computational power used, and the resulting model performance. Empirical evidence has shown that the performance of LLMs on downstream tasks improves as the model size and the amount of pre-training data increase (Kaplan et al., 2020; Hoffmann et al., 2022). This power-law relationship suggests that larger models trained on more data tend to capture more complex patterns and generalize better to new tasks.
Another intriguing property of LLMs is the emergence of abilities that were not explicitly trained for, often referred to as emergent abilities (Wei et al., 2022). This ability suggests that LLMs can capture and leverage complex linguistic patterns and knowledge from the pre-training data, enabling them to perform tasks beyond their original training objective. Furthermore, LLMs have demonstrated the capability of in-context learning (Brown et al., 2020), where they can perform tasks based on a few examples provided in the input prompt, without the need for explicit fine-tuning. This highlights the models’ ability to rapidly adapt to new tasks and generalize from limited examples.
The development and deployment of LLMs typically involve three key stages: training, adaptation, and evaluation (Zhao et al., 2023d). The training stage focuses on learning general language representations from large-scale unlabeled corpora, capturing the underlying patterns and structures of natural language. The adaptation stage involves tailoring the pre-trained model to specific domains or tasks, often through techniques such as domain adaptation or task-specific fine-tuning (Raffel et al., 2020). This step is crucial for optimizing the model’s performance on the target application and ensuring its outputs are aligned with the desired objectives. Finally, the evaluation stage involves assessing the performance of the adapted model through various metrics and benchmarks to ensure that it meets the required standards and performs effectively in real-world scenarios.
From a data-centric perspective, the performance and generalization capabilities of large language models (LLMs) are highly dependent on the quality and diversity of the data used throughout the different stages of development (meta llama, 2024). During the training stage, the pre-training data is crucial in shaping the model’s language understanding and generation abilities. Efforts have been made to curate high-quality datasets (Gao et al., 2020) that cover a wide range of domains and languages. In the adaptation stage, domain-specific data is often used to fine-tune the model and enhance its performance on targeted tasks. The evaluation stage involves using real-world data for inference and assessment, underscoring the importance of data representativeness and robustness.
However, challenges persist in addressing data bias, privacy concerns, and the need for more diverse and representative datasets that capture the richness and nuances of human language across different domains and cultures. Overcoming these challenges requires ongoing research in data collection, curation, and augmentation techniques (Raffel et al., 2020; Perez et al., 2022).
The development of LLMs has revolutionized the field of NLP, enabling significant advances in language understanding and generation. The scaling law, emergent capabilities, and in-context learning capabilities of LLMs have opened up new avenues for research and application.
2.2. Multimodal Large Language Models
Multimodal large language models (MLLMs) extend the capabilities of traditional LLMs by leveraging their comprehension, generation, and task-solving abilities across different modalities, including natural language (text and audio) and visual information (videos, images, and 3D models). By integrating LLMs with multimodal projectors, MLLMs can process various types of information, demonstrating strong understanding and analytical capabilities to address downstream tasks across various modalities (Wu et al., 2023b).
The development of MLLMs is driven by the need to address real-world problems that often involve multiple modalities. For instance, in robotics, an MLLM can process visual input from cameras, interpret natural language instructions, and analyze structured data from sensors to perform complex tasks (Zeng et al., 2023). Similarly, in healthcare, MLLMs can analyze medical images, process electronic health records, and interpret patient-doctor conversations to aid in diagnosis and treatment planning.
The training of MLLMs typically involves three distinct stages: pre-training, adaptation, and evaluation. First, the pre-training stage involves training the model on large-scale multimodal datasets to learn general representations and cross-modal interactions. Second, the adaptation stage fine-tunes the pre-trained MLLM on task-specific datasets, enabling it to specialize its knowledge and skills for particular applications. Finally, the evaluation stage assesses the adapted MLLM’s performance by processing multimodal inputs and generating appropriate outputs, ensuring its effectiveness in solving real-world problems (Li et al., 2022b; Liu et al., 2024b, 2023c; Zhang et al., 2024c).
From a data-centric perspective, the performance of MLLMs heavily relies on the quality and diversity of the multimodal datasets used throughout their development stages (Liu et al., 2024b, 2023c). In the pre-training stage, large-scale datasets encompassing multiple modalities are essential for learning robust and generalizable representations (McKinzie et al., 2024). During the adaptation stage, task-specific datasets with high-quality annotations are crucial for effective fine-tuning (Liu et al., 2023c). Lastly, in the evaluation stage, MLLMs must be assessed on their ability to handle diverse and potentially noisy multimodal inputs encountered in real-world scenarios (Zhang et al., 2024c).
MLLMs represent a significant advancement in AI by extending the capabilities of traditional LLMs to process and analyze multiple modalities. By leveraging large-scale multimodal datasets and specialized training techniques, MLLMs can solve complex real-world problems that require understanding and reasoning across different types of information. As research in this field progresses, we can expect MLLMs to play an increasingly important role in various domains, from robotics and healthcare to education and entertainment.
Data-Centric MLLMsData Colloctingand ProcessingData CollectingSourcesCommonWebpages CommonCrawl Project (Raffel et al., 2020), Self-Crawled Data: General Approches (Sharma et al., 2018; Changpinyo et al., 2021), Focused Approaches (Wikipedia (Gao et al., 2020; Srinivasan et al., 2021; Schamoni et al., 2018), Flickr (Young et al., 2014), FreeSound (Mei et al., 2023; Fonseca et al., 2022)), Querying search engines-based approaches (Wu et al., 2017; Gu et al., 2022) Social Media Stack Exchange (Soboleva et al., 2023; Chen et al., 2023g), Reddit (Baumgartner et al., 2020; Desai et al., 2021; Overbay et al., 2023), YouTube (Miech et al., 2019; Zhou et al., 2017), X (Barbieri et al., 2020) Academic Papers ArXiv (Computer, 2023; Gao et al., 2020; Liu et al., 2023d), S2ORC (Lo et al., 2020; Soldaini and Lo, 2023; Soldaini et al., 2024) Books Project Gutenberg (Rae et al., 2019), Smashwords (Zhu et al., 2015), Bibliotik (Gao et al., 2020), Amazon.com (Mishra et al., 2019), Old Photo Books (Okamoto et al., 2023) Codes BigQuery (Nijkamp et al., 2022), GitHub (Kocetkov et al., 2022; Nijkamp et al., 2022; Gao et al., 2020), The Stack (Li et al., 2023a) ProfessionalSources Legal Domain: FreeLaw (Gao et al., 2020); Math Domain: Khan Academy, DeepMind (Saxton et al., 2018); Medical Domain: Online Medical Websites (Li et al., 2023j; Ben Abacha and Demner-Fushman, 2019), Knowledge Bases (Li et al., 2023j; Welbl et al., 2018), In-Hospital Database Systems (Johnson et al., 2023, 2019; Bustos et al., 2020); Financial Domain: Financial Platforms and Forums (Ok, 2023; Lu et al., 2023b) Data ProcessingMethodsFiltering Textual Filtering (Joulin et al., 2016; Grave et al., 2018; Raffel et al., 2020, 2020; Yuan et al., 2021; Gunasekar et al., 2023), Image-Level Filtering (Zhu et al., 2024; Guo et al., 2021; Yang et al., 2022), Video-Level Filtering (Wang et al., 2023a; Blattmann et al., 2023; Farnebäck, 2003; Baek et al., 2019) Deduplication Exact Duplication (Raffel et al., 2020; Suárez et al., 2019; Lee et al., 2022), Approximate Duplication (Computer, 2023; Lee et al., 2022; Penedo et al., 2023), Sentence-Level and Document-Level Semantic Duplication (Reimers and Gurevych, 2019; Song et al., 2020; Silcock et al., 2022; Abbas et al., 2023) DataEnhancement Enhance X Modality Data (Li et al., 2023k; Luo et al., 2024; Wei et al., 2023), Enhance Text Data (Nguyen et al., 2023; Fan et al., 2024; Lai et al., 2023a; Chen et al., 2023e) Commonly-Used DatasetsImage datasets Image-Caption Datasets (Schuhmann et al., 2021, 2022, 2022), Content Descriptive Image-Caption Dataset (Chen et al., 2015; Young et al., 2014; Chen et al., 2022), Interleaved Image-Text Documents (Alayrac et al., 2022; Zhu et al., 2024; Laurençon et al., 2024), Visual Question Answer Image-Caption Datasets (Goyal et al., 2019; Zhu et al., 2016; Biten et al., 2019) Video datasets Video Captioning Datasets (Xu et al., 2016; Bain et al., 2021; Alayrac et al., 2022), Video Question Answering Datasets (Chen and Dolan, 2011; Xu et al., 2016; Jang et al., 2017), Interleaved Video-Text Dataset (Sun et al., 2023b) Audio datasets Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022c), Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) 3D datasets 3D Captioning Datasets (Dai et al., 2017; Armeni et al., 2017; Zheng et al., 2020a) Data-CentricPre-trainingDomain Mixture Pure Text Domain Mixture (Xie et al., 2024a, b; Du et al., 2022), Multimodal Domain Mixture (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a) Modality Mixture Image Caption Data Mixture (McKinzie et al., 2024; Lai et al., 2023b), Image and Video Caption Data Mixture (Lin et al., 2023b; Zhang et al., 2023; Luo et al., 2023; Jin et al., 2023; Li et al., 2023g; Han et al., 2023; Li et al., 2023c, h; Wang et al., 2023a; Chen et al., 2023c) Quality Selection Active Learning-Based Selection (Xu et al., 2023a), Selection Before Training (Gadre et al., 2024), Distribution-Agnostic Selection (Gadre et al., 2024; Wang et al., 2024d; Mahmoud et al., 2024), Distribution-Aware Selection (Wang et al., 2024a; Gadre et al., 2024) Data-Centric AdaptationSupervisedFinetuningData Processing SFT Data Processing (Zhu et al., 2023; Chen et al., 2023e) Data Generation Captioning Instruction-Response Datasets Generation (Zhu et al., 2023; Xu et al., 2024; OpenAI, 2023), Question Answer Instruction-Response Datasets Generation (Liu et al., 2023c; Singh et al., 2019; Li et al., 2023l), Reasoning Instruction-Response Datasets Generation (Chen et al., 2023j, h, m), Other Instruction-Response Datasets Generation (Li et al., 2023l; Panagopoulou et al., 2023; Dai et al., 2024) Data Selection Coreset-Based Selection (Chen et al., 2023k; Das and Khetan, 2023; Sorscher et al., 2022), LLMs-Based Selection (Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e), gradient-Based Selection (Paul et al., 2021; Attendu and Corbeil, 2023; Xia et al., 2024), Self-Instruction-Based Selection Methods (Li et al., 2023m, d; Kung et al., 2023; Liu et al., 2024c) Alignment Human Preference Alignment: Pure Text (Ouyang et al., 2022; Li et al., 2023g), Multimodal (Liu et al., 2023b; Sun et al., 2023a; Chen et al., 2023i; Ouyang et al., 2022) Data-CentricEvaluationData Evaluation Dataset Diversity (Heusel et al., 2017; Friedman and Dieng, 2023); Dataset Quality (Honovich et al., 2022a; Rohrbach et al., 2018; Jing et al., 2023); Dataset Similarity (Sun and Saenko, 2016; Jiang et al., 2022; Pillutla et al., 2023) Evaluation DatasetsCaptioningTasks Image Captioning Datasets (Goyal et al., 2019; Hudson and Manning, 2019; Marino et al., 2019), Video Captioning Datasets (Xu et al., 2017; Chen and Dolan, 2011; Yu et al., 2019), Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022c) QuestionAnswering Tasks Image Question Answering Datasets (Chen et al., 2015; Young et al., 2014; Karpathy and Fei-Fei, 2015), Video Question Answering Datasets (Chen and Dolan, 2011; Venugopalan et al., 2015; Xu et al., 2016), Audio Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) Perception &Reasoning Tasks Perception and Reasoning Datasets (Singh et al., 2019; Masry et al., 2022; Kembhavi et al., 2016), Spatial Reasoning Datasets (Liu et al., 2023b; Kazemzadeh et al., 2014; Mao et al., 2016) Other NoteworthyTasks Classification Datasets (Kiela et al., 2020; Deng et al., 2009; Piczak, 2015; Wu et al., 2015; Goyal et al., 2017), Multimodal Dialogue Datasets (Bai et al., 2023b)
Figure 2. Overview of Data-Centric MLLMs
[…]
[Overview of Data-Centric MLLMs]Categorization of Data-Centric MLLMs work
2.3. Data-Centric AI and Why Data-Centric LLMs
The field of artificial intelligence (AI) has experienced a paradigm shift towards data-centric approaches in recent years (Zha et al., 2023). Data-centric AI emphasizes the critical role of data quality, diversity, and representativeness in building robust and effective AI systems. This shift acknowledges that the performance of AI models is heavily based on the quality and characteristics of the data used for training, rather than solely focusing on algorithmic improvements.
In the context of LLMs and MLLMs, data play a pivotal role in determining their capabilities and limitations. LLMs, such as GPT-3 (Brown et al., 2020) and BERT (Kenton and Toutanova, 2019), are trained on vast amounts of textual data to learn the intricacies of language and generate coherent, contextually relevant outputs. Similarly, MLLMs, which integrate multiple modalities such as text, images, and speech, require diverse and well-aligned datasets in different modalities to learn meaningful cross-modal representations (Gadre et al., 2024). The quality, diversity, and representativeness of the training data directly impact the models’ ability to understand and generate language, as well as their capacity to reason and perform tasks across multiple modalities (meta llama, 2024; Liu et al., 2023c; Chen et al., 2023e).
Analyzing LLMs and MLLMs from a data-centric perspective offers several advantages. First, it enables researchers to identify and address potential biases and limitations in the training data that can propagate to the models’ outputs. By carefully curating and augmenting datasets to ensure diversity and representativeness, researchers can mitigate biases and improve the fairness and generalizability of the models (Chen et al., 2024a). Second, a data-centric approach allows for a deeper understanding of the models’ capabilities and limitations based on the characteristics of the training data (McKinzie et al., 2024). By systematically varying data properties and evaluating model performance, researchers can gain insights into the specific data attributes that contribute to the models’ success or failure in various tasks. This understanding can guide the development of more efficient and effective data collection and curation strategies.
Moreover, adopting a data-centric perspective in the development of LLMs and MLLMs opens up opportunities for data-efficient learning and adaptation (Liu et al., 2024b, 2023c; Chen et al., 2023e). Techniques such as few-shot learning and prompt engineering enable models to adapt to new tasks or domains with limited task-specific data. This adaptability is particularly valuable in scenarios where labeled data is scarce or expensive to obtain. Additionally, data-centric approaches facilitate the development of more interpretable and explainable models by linking model behaviors to specific data characteristics, allowing for the identification of relevant data subsets for debugging and error analysis.
The shift towards data-centric AI has underscored the crucial role of data in the development and evaluation of LLMs and MLLMs. By prioritizing data quality, diversity, and representativeness, researchers can build more robust, unbiased, and adaptable models. Analyzing these models from a data-centric perspective offers valuable insight into their capabilities, limitations, and potential for improvement. As the field continues to advance, a data-centric approach will be essential to develop LLMs and MLLMs that can effectively understand, generate, and reason across multiple modalities while ensuring fairness, interpretability, and generalizability.
2.4. Categories of Data-Centric MLLMs
In this survey, we review previous work on MLLMs from a data-centric perspective. Our taxonomy is primarily based on the different stages of MLLM development, as illustrated in Figure 2. Specifically, we organize our article around a data pipeline, shown in Figure 1, where the data pipeline for MLLMs comprises three stages: pre-processing (including data collection and data processing in Section 3), pre-training data processing (Section 4), and adaptation data processing (Section 5).
Along this pipeline, we discuss data-related model performance and analyze how dataset curation affects model outcomes. During the pre-processing stage, large-scale multimodal data are collected from various sources and constructed into datasets. Processing steps such as filtering, deduplication, and data enhancement are applied to improve data quality. In the pre-training stage, data selection is performed to choose appropriate data for model training. In the adaptation stage, we focus on data processing and generation for supervised fine-tuning and human preference alignment. Furthermore, we discuss data evaluation metrics and evaluation datasets for MLLMs, examining their construction and main characteristics. Finally, we suggest possible future directions for further research on data-centric MLLMs.
3. Data collecting and processing
The first and fundamental step in training MLLMs is to collect and process sufficient data from various sources. In this section, we first introduce the sources for data collection. Next, we summarize the processing steps for multimodal data, including commonly used filtering, deduplication, and data enhancement methods. Finally, we provide an overview of commonly used datasets for MLLMs. These datasets are utilized in the pre-training, adaptation, and evaluation stages of MLLMs.
3.1. Data Collecting Sources
We first introduce the sources where researchers typically find raw data for MLLM pre-training. Commonly used data sources can be divided into six categories, as shown in Figure 3: common webpages, social media, academic papers, books, code repositories, and professional sources. Data from these different sources have various characteristics, each contributing uniquely to improving different abilities of the models.
Figure 3. Commonly used data sources.
[…]
[Commonly used data sources.]Commonly used data sources.
Common Webpages.
Webpages are the main source for collecting large-scale training corpus. CommonCrawl project serves as the most commonly used start-point for large-scale webpages. It is instrumental in generating large-scale pre-training corpora like C4 (Raffel et al., 2020) for LLM pre-training. Beyond textual content, CommonCrawl’s vast archive of web pages, which includes numerous image-text pairs, has also become a vital resource for constructing multimodal pre-training datasets such as LAION-5B (Schuhmann et al., 2022).
Apart from CommonCrawl, several works focus on selecting specific webpages and crawling data independently. For textual datasets, initiatives like WuDaoCorpora (Yuan et al., 2021) crawl thousands of web pages to construct vast textual corpora. In the multimodal dataset sector, there are three main self-crawling approaches: general, focused on specific platforms, and querying search engines for images. Examples include AI Challenger Captions (Wu et al., 2017) and Wukong (Gu et al., 2022). General approaches, such as CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021), involve crawling billions of web pages without targeting specific platforms. Focused approaches, on the other hand, involve selecting specific webpages for crawling. For example, Wikipedia is a commonly used resource that serves as a rich and accurate source for text-based single-modal pre-training datasets like the Pile (Gao et al., 2020). It also plays a crucial role in creating multimodal pre-training datasets such as WIT (Srinivasan et al., 2021) and WikiCaps (Schamoni et al., 2018).
For the vision modality, Flickr (Young et al., 2014) serves as an important source, containing photos and videos shared by online users. This contributes to image-text datasets such as Flickr30k (Young et al., 2014). For the audio modality, platforms like BBC Sound Effects, FreeSound, and SoundBible are commonly used, contributing to datasets such as WavCaps (Mei et al., 2023) and FSD50K (Fonseca et al., 2022).
Social Media.
Social media enriches MLLM training datasets with real-time, varied, and colloquial text data, as well as multimodal content, capturing human expression and interaction. Stack Exchange is a vital source of high-quality textual and visual content, hosting extensive forums with rich dialogues and images. This contributes to text-only datasets such as SlimPajama (Soboleva et al., 2023) and image-text datasets such as VQAonline (Chen et al., 2023g). Beyond Stack Exchange, Reddit features user-shared content with discussions and voting on a wide array of topics, contributing to text-rich datasets such as Pushshift Reddit (Baumgartner et al., 2020) and visual datasets such as RedCaps (Desai et al., 2021) and MREDDITSUM (Overbay et al., 2023).
For video data, YouTube stands out as a video-sharing platform with over 500 hours of video uploaded per minute. It offers a vast dataset of multimodal information, contributing to datasets such as HowTo100M (Miech et al., 2019), which leverages instructional videos for multimodal learning, and YouCook2 (Zhou et al., 2017), focused on cooking videos. X, commonly known by its former name, Twitter, is a popular social media platform for sharing text messages, images, audio, and videos. Datasets derived from X include TweetEval (Barbieri et al., 2020), designed for evaluating sentiment and stance.
Academic Papers.
Academic papers offer high-quality, authoritative content, enriching pre-training datasets with specialized knowledge and formal language, which is critical for developing language models with expertise in academic and professional domains. Currently, academic papers are primarily utilized as sources for text-based pre-training datasets.
arXiv serves as a crucial data source, offering a vast array of prepublished scientific research for training specialized language models. This is evidenced by its role in creating datasets like RedPajama-Data-1T (Computer, 2023), Pile-arXiv (Gao et al., 2020), and image-text datasets such as MMC (Liu et al., 2023d).
S2ORC (Lo et al., 2020) is a corpus containing 81.1 million academic papers, offering metadata, abstracts, references, and full texts for 8.1 million open-access works. It compiles materials from various sources, forming the largest publicly accessible academic dataset, such as peS2o (Soldaini and Lo, 2023) and Dolma (Soldaini et al., 2024).
Books.
Books provide a rich source of high-quality textual content for language model training and serve as fertile ground for multimodal pre-training datasets. They offer diverse visual data from book covers and illustrations, enriching models with a blend of literary depth and visual context.
Project Gutenberg acts as a pivotal source for language modeling and analysis with its extensive library of more than 70,000 free eBooks, fueling datasets such as PG-19 (Rae et al., 2019). The BookCorpus (Zhu et al., 2015), sourced from Smashwords, encompasses a diverse collection of 11,038 self-published novels that span genres such as romance, science fiction, and fantasy. This extensive corpus has been instrumental in training landmark models such as GPT (Radford et al., [n. d.]) and BERT (Kenton and Toutanova, 2019).
Bibliotik consists of a mix of fiction and non-fiction books. Books3 (Gao et al., 2020), part of the Pile dataset, is derived from the Bibliotik private tracker made available by Shawn Presser in 2020, containing about 197,000 books. This dataset is essential for context modeling and narrative research, accounting for 2.1% of the RedPajama-Data-1T dataset.
Meanwhile, books also serve as an important source for constructing multimodal datasets. Book covers hold vast multimodal information, combining visual elements with metadata such as author names, titles, and genres. This fusion of data supports tasks like genre classification and visual question-answering. The OCR-VQA dataset (Mishra et al., 2019), sourced from Amazon.com, exemplifies how book covers can be leveraged to build rich pre-training datasets for MLLMs, demonstrating the value of covers in enhancing model understanding of visual and textual content. Old photo books offer a unique advantage for constructing multimodal pre-training datasets by providing historical visual and textual contexts, enriching models’ understanding of past societies and landscapes. For example, previous work collected 9,516 image-text pairs from 175 Japanese old photo books (Okamoto et al., 2023).
Code.
Training LLMs on extensive code datasets significantly boosts their capability to generate high-quality, accurate programs, as evidenced by their enhanced performance in expert-designed tests and complex programming challenges. The most common sources for codes are BigQuery and GitHub. BigQuery, a Google data warehouse, contributes to code datasets such as BIGQUERY, which encompasses open-source code in six programming languages: C, C++, Go, Java, JavaScript, and Python. This dataset is aimed at training multilingual CODEGEN models (CODEGEN-MULTI) (Nijkamp et al., 2022). GitHub, a premier platform for software collaboration, supports datasets like The Stack (Kocetkov et al., 2022), BigPython (Nijkamp et al., 2022), and the Pile (Gao et al., 2020) with its vast, open-source repositories and global developer community. The Stack (Kocetkov et al., 2022) leverages GitHub’s permissively licensed code and governance tools to support the StarCoder (Li et al., 2023a) project with a diverse programming language dataset.
Domain-Specific Sources.
To enhance the performance of LLMs in specific domains, general pre-training is supplemented with incremental, domain-specific pre-training. In the legal domain, pre-training data is primarily textual. FreeLaw stands out by offering free access to judgments from India’s higher courts, exemplified by the Pile-FreeLaw (Gao et al., 2020) dataset, which structures these judgments for machine learning.
In the math domain, textual datasets primarily originate from Khan Academy’s exercises for foundational concepts and DeepMind’s Pile-DeepMind Mathematics (Saxton et al., 2018) for advanced, algorithmically generated problems. Additionally, DVQA (Kafle et al., 2018) introduces a multimodal aspect by combining bar charts with question-answer pairs generated through Matplotlib.
In the medical domain, pre-training data comes from online medical websites, knowledge bases, and in-hospital database systems. Online medical websites such as Qianwen Health and PubMed contribute to datasets like Huatuo-26M (Li et al., 2023j) and MedQuAD (Ben Abacha and Demner-Fushman, 2019). Knowledge bases like Wikipedia also contribute to Huatuo-26M (Li et al., 2023j) and MedHop (Welbl et al., 2018). In-hospital database systems, including electronic health record systems (EHR) and ICU-specific clinical information systems, are used in MIMIC-IV (Johnson et al., 2023). Multimodal datasets in radiology, such as MIMIC-CXR-JPG (Johnson et al., 2019) and PADCHEST (Bustos et al., 2020), are generated from in-hospital radiology reports.
In the financial domain, pre-training primarily relies on textual data sourced from both English and Chinese platforms. English contributions come from EDGAR and the SEC Financial Statement and Notes Data Sets, which enhance the FinTree project (Ok, 2023). Chinese sources include major sites like Sina Finance and Tencent Finance, as well as Eastmoney for specific documents, and forums like Guba and Xueqiu, culminating in the comprehensive BBT-FinCorpus (Lu et al., 2023b). This assortment of sources underscores the rich textual foundation for financial model pre-training.
3.2. Data Processing
3.2.1. Filtering
Data filtering in MLLM training is critical for enhancing model reliability and efficiency. Filtering out undesirable content is essential, as even minimal exposure to hate speech can negatively influence model behavior (Luccioni and Viviano, 2021; Gunasekar et al., 2023). For multimodal datasets, previous work has focused on filtering out unwanted data from different modalities separately. Specifically, for X-Text datasets, they consider filtering data based on both the text and the X modality independently.
Textual filters primarily encompass language and content filtration. In language filtering, documents or sentences that fall below a certain threshold for a specific language are removed. For English-only datasets, commonly used tools like Langdetect and FastText operate at the document level. For instance, C4 (Raffel et al., 2020) utilizes Langdetect to filter out any non-English pages with a probability below 0.99, reducing the Common Crawl April 2019 corpus from 6TB to 20GB (c4/en.noclean). Another renowned English language filter, FastText, employed by Dolma (Soldaini et al., 2024), eliminates documents classified as English with a probability less than 0.5, filtering out over 60% of documents. RefinedWeb (Penedo et al., 2023) utilizes a higher FastText threshold (0.65) to obtain purer English documents. For Chinese datasets, WuDaoCorpora (Yuan et al., 2021) implements heuristic rules to filter out web pages containing fragments with more than ten successive non-Chinese characters at the document level. CLUECorpus2020 (Xu et al., 2020) adopts a sentence-level filtering approach, selecting sentences whose language type is Chinese if a language is mentioned. For multilingual datasets, FastText, trained on Wikipedia data to classify 176 languages, remains a preferred choice, capable of processing 1,000 documents per second on a single CPU core. ROOTS (Laurençon et al., 2022) utilizes FastText for document-level language classification, resulting in a 1.6TB dataset containing 59 languages. For code datasets, detection methods are more elementary, often based on file extensions. For instance, in Palm (Chowdhery et al., 2023), files are filtered based on filename extensions to restrict to one of 24 common programming languages, resulting in 196GB of source code.
Another type of textual filtering is content filtering. Content filtering includes removing toxic and distracting content. Toxic content filtering targets text that is deemed rude, disrespectful, or unreasonable, employing heuristic and rule-based methods. For instance, in C4 (Raffel et al., 2020), pages containing words listed as ”Dirty, Naughty, Obscene, or Otherwise Bad Words” are removed. Distracting content includes short or incomplete sentences, ”lorem ipsum” text, and useless identifiers like HTML, CSS, and JavaScript tags. Filtering methods may involve multiple rule-based filters, such as punctuation-based segmentation in WudaoCorpora (Yuan et al., 2021), or machine learning approaches, as seen in phi-1 (Gunasekar et al., 2023). In phi-1, GPT-4 annotates a small portion of high-quality data, which is then used to train a random forest classifier for high-quality content classification.
For image-level filtering, the most fundamental step is to remove images with excessively low resolution, as these images often fail to convey effective information. Additionally, it is necessary to filter out images with inappropriate aspect ratios, as such unconventional images frequently resemble banner-like advertisements (Zhu et al., 2024). Furthermore, similar to removing harmful content in text filtering, such as NSFW material, it is necessary to filter out such content from images. This can be achieved by training a binary classification model using appropriate datasets. Additionally, any potential occurrences of human faces or other sensitive elements in the images should be detected using a face detector and blurred accordingly (Guo et al., 2021). These filtering processes have a minimal effect on the overall model performance (Yang et al., 2022).
For video-level filtering, previous work has leveraged image or image-text filtering methods to handle static content and applied specific techniques for dynamic content. Video filtering involves four main components: scene transition detection, video quality and integrity improvement, refinement and coherence evaluation, and modality completeness. For scene transition detection, algorithms are used to remove abrupt scene changes, ensuring a cleaner dataset free from disruptive transitions (Wang et al., 2023a; Blattmann et al., 2023). To improve video quality and integrity, three aspects are considered: motion, textual information in the video, and resolution and frame rate. For motion analysis, dense optical flow (Farnebäck, 2003) is typically utilized to measure motion complexity and filter out static or repetitive sequences with limited educational value. To detect textual information in videos, OCR (Baek et al., 2019) is applied to identify and remove clips with substantial text, which can otherwise misguide the model’s visual interpretation. The resolution and frame rate of training video datasets are standardized, enabling VLMs to learn from uniformly structured data (Wang et al., 2023a). In constructing video-text datasets, ensuring the relevance and clarity of annotations is paramount. Focused efforts are made to refine the textual content of video-text pairs, particularly subtitles, to ensure they are contextually meaningful and devoid of promotional or irrelevant material (Xu et al., 2023c). Models such as CLIP are employed to evaluate the coherence between video frames and the accompanying text (Chen et al., 2024b; Xu et al., 2023c; Wang et al., 2023a), ensuring that annotations accurately reflect the video content. In some studies, clips missing any modalities—vision, audio, or subtitles—are excluded to meet the requirements for comprehensive understanding across all modalities (Chen et al., 2024b).
3.2.2. Deduplications
Previous work has found there are great amount of duplicate data in various training datasets. For example, the BOOK Corpus (Zhu et al., 2015) contains thousands of duplicated books (Bandy and Vincent, 2021). The commonly used curated web crawl dataset C4 (Raffel et al., 2020) contains a single 61 word English sentence that repeated more than 6 hundred thousands times (Lee et al., 2022). More repeated data in training datasets can increase the rate of emitting memorized training data verbatim (Carlini et al., 2022; Kandpal et al., 2022), and deduplication of training datasets can prevent the memorization problem, thus alleviating privacy concerns (Kandpal et al., 2022). Moreover, duplication of training data can also cause the performance degradation of the pre-trained models (Hernandez et al., 2022). Training on deduplicated datasets can save the training cost while does not hurt the model perplexity (Lee et al., 2022).
Existing deduplication methods consider exact duplication, approximate duplication and semantic duplication in sentence-level (or sequence-level) and document-level. Exact deduplication is the most simple way to remove duplication. For sentence or sequence level, exact deduplication consider to remove exact string matching between sentence (or sequence) (Raffel et al., 2020; Suárez et al., 2019). To improve computational and memory efficiency, Suffix Arrays (Manber and Myers, 1993; Lee et al., 2022) and Bloom Filters (Computer, 2023; Soldaini et al., 2024) are used to achieve parallelized linear time complexity. For document level exact deduplication, URL deduplication is considered to remove exactly same web pages (Penedo et al., 2023). Approximate deduplication methods mainly focus on document-level duplicates. Usually, Locality Sensitive Hashing (LSH)-based MinHash (Broder, 1997) or SimHash (Charikar, 2002) methods are used to remove approximately duplicate documents (Computer, 2023; Lee et al., 2022; Penedo et al., 2023). These methods can achieve document-level deduplication with linear time and space complexity of the document number and can be implemented in a highly distributed setting (Lee et al., 2022). Apart from hashing-based approximate deduplication, recently there are some work consider to leverage pre-trained foundation models as semantic embedding metric for document-level embedding (Silcock et al., 2022; Kaddour, 2023; Abbas et al., 2023; Tirumala et al., 2023). These Semantic deduplication methods consider using Sentence-BERT (Reimers and Gurevych, 2019) MPNET (Song et al., 2020) for embedding (Silcock et al., 2022), E5-Large (Wang et al., 2022b) for embedding (Kaddour, 2023), and OPT-125M (Zhang et al., 2022) for embedding (Abbas et al., 2023; Tirumala et al., 2023). These semantic deduplication processes are usually after the exact deduplication and approximate deduplication processes, removing the semantic duplicates by clustering the embedding points and keeping representative data in each clusters. However, how to choose an appropriate pre-trained foundation models and whether this models can embed semantic information efficiently are still open questions.
For image-text pairs or interleaved image-text documents, common methods of image deduplication include using the image URL (Zhu et al., 2024) or employing pHash (Zauner, 2010) algorithms. It’s worth noting that in the realm of interleaved image-text documents, there is currently no universally accepted, reliable method for deduplication based on both image and text elements simultaneously.
For Video multimodal data, video fingerprinting technology is uniquely applied to videos to identify and remove exact duplicates, addressing the challenge of video re-uploads and mirrored content, which is more prevalent and complex in videos than in static images(Xu et al., 2023c).
3.2.3. Data Enhancement
Data enhancement of multimodal datasets usually focus on two aspects, enhancing the X modality data, and enhancing the text data. Traditional data augmentation methods for single modality data have been discussed thoroughly in previous work (Xu et al., 2023d; Cauli and Reforgiato Recupero, 2022; Ko et al., 2015; Wei et al., 2020; Zhao et al., 2021; Feng et al., 2021). In this work, we only consider data enhancement methods for MLLMs datasets.
For vision-language models, enhancing the quality of image-caption datasets is crucial for training MLLMs. Improving the captions not only enhances their alignment with the images but also prevents the discarding of highly informative images due to poor text quality. Rewriting text using BLIP2 has been shown to effectively improve training results (Nguyen et al., 2023). However, for large-scale image-text datasets (approximately 1.28 billion), this method experiences diminishing returns in terms of ImageNet accuracy, possibly due to the lack of diversity in the generated captions (Nguyen et al., 2023). Approaches like LaCLIP (Fan et al., 2024) and VeCLIP (Lai et al., 2023a) employ LLMs with carefully designed prompts to generate more diverse captions. ShareGPT4V (Chen et al., 2023e) employed the 100K high-quality captions generated by GPT4-Vision to fine-tune an alternative caption model and named it as Share-Captioner. The Share-Captioner is capable of generating highly content-related descriptions with unified instruction for the pre-train dataset. MLM (Wang et al., 2024d) used GPT-4 or GPT-4V to constructing multimodal instruction tuning data on proposed quality scoring tasks to fine-tune MLM to realize accurate quality assessment. Then they adopt the fine-tuned MLM Filter to generate quality scores for each data point in the data pool and then select the high-quality data.
Improving image resolution can enhance the performance of MLLMs. Initially, MLLMs primarily processed fixed, lower-resolution inputs, typically around 224 pixels (Liu et al., 2024b; Chen et al., 2023j; Zhu et al., 2023). Recent models such as LLaVA-1.5 (Liu et al., 2023c) and BLiVA (Hu et al., 2024) have improved performance by increasing the input resolution to 336 pixels and integrating task-specific global features. Furthermore, models like Qwen-VL (Bai et al., 2023a) and OtterHD (Li et al., 2023n) have pushed resolution support to 448 pixels, incorporating fine-tuning in the visual encoder during training, while maintaining the original image size during inference, leading to more precise segmentation recognition. Notably, Monkey (Li et al., 2023k) has significantly raised the resolution to 896 pixels by employing multiple visual encoders and leveraging fine-tuning techniques from Qwen-VL. Models like LLAVA HR (Luo et al., 2024) and Vary (Wei et al., 2023) have introduced additional visual encoders to capture more complex features, requiring extensive pre-training tasks. LLAVA UHD (Xu et al., 2024) and Ureader (Ye et al., 2023a) have improved the model’s capacity to understand detailed image features through adaptive region segmentation. However, a fundamental challenge remains within the MLLM architecture: models using lower-resolution inputs struggle to detect fine details, whereas those with higher resolutions may underperform in tasks requiring a broader global understanding.
3.3. Commonly-used Datasets for MLLMs
This section we will briefly introduce commonly used multimodal datasets from different modality. The comprehensive summary of text-only datasets has been elaborately discussed in previous literature (Liu et al., 2024a); therefore, this part will not be extensively covered within the scope of this discussion. For the details of datasets introduced in this section, please refer to A.1.
3.3.1. Image Datasets
For vision langauge models, the image-text datasets commonly used can be categorized into three distinct types: image-caption datasets, interleaved image-text datasets, and visual question answering datasets.
Image-Caption Datasets.
Image-caption pairs datasets are the most commonly used datasets in the pre-training stage of MLLMs. By leveraging the captions of each image, one can align the image modality with the text modality, enabling LLMs to understand image information. General image-caption datasets contain images with short captions, typically one sentence or a few words, that describe the key features of the image. These datasets vary significantly in size and scope, making them adaptable for various research and application needs. They are vital for the training and research of multimodal language-vision models, as they provide a corpus of image-text pairings that facilitate the understanding and generation of textual descriptions from visual inputs. Hence, collecting high-quality image-caption pairs is crucial in this phase. Fortunately, previous research in computer vision has amassed plenty of high-quality image-caption pairs.
Notably, the LAION series (comprising LAION-400M (Schuhmann et al., 2021) and LAION-5B (Schuhmann et al., 2022)) represents some of the largest compilations, aggregating billions of pairs from the CommonCrawl archive. These datasets emphasize the maintenance of high quality through meticulous filtering based on the relevance of image-text pairs and the exclusion of inappropriate content. Derived from the LAION-5B subset, the LAION-COCO (Schuhmann et al., 2022) dataset is specifically designed to explore the impact of synthetic captions on model training. Also curated from CommonCrawl project, the COYO-700M (Minwoo Byeon, 2022) dataset emphasizes the informative correlation between images and alt-texts. The DataComp (Gadre et al., 2024) challenge introduces COMMONPOOL, a substantial multimodal dataset constructed from 12.8 billion image-text pairs that encourages innovation in dataset design.
Self-crawled captioning datasets including SBU (Ordonez et al., 2011), CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), Wukong (Gu et al., 2022), ALT200M (Hu et al., 2022) and LTIP (Alayrac et al., 2022). Among these datasets, SBU (Ordonez et al., 2011) is the smallest containing only 1 million image-caption pairs curated from Flickr. Several datasets are curated from images with alt attributes from webpages. For example, Conceptual Captions (CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021)) use generalized alt-text from the web and are processed with multiple filters, containing 3.3 million and 12 million image-text pairs respectively. Also leveraging alt-text for images on webpages, ALT200M (Hu et al., 2022) dataset collects 200 million images with their alt attributes, aiming for understanding scaling of vision-language models. For multilingual image-caption datasets, a representative one is the Wukong (Gu et al., 2022) dataset, which addresses the scarcity of large-scale datasets in the Chinese language, offering 100 million quality-assured pairs that are instrumental for the development of Chinese vision-language pre-training models. Larger dataset such as Long text & image pairs (LTIP) (Alayrac et al., 2022) dataset provides a unique collection of 312 million images with lengthier textual descriptions, enhancing models like Flamingo that are tailored for complex multimodal tasks. Collectively, these datasets advance the understanding of the complex interactions between visual and textual data, supporting the development of sophisticated artificial intelligence models.
Content Descriptive Image-Caption Dataset
Content descriptive image-caption datasets contain longer caption with more descriptive information. Usually each image in content description image-caption datasets contain at least 5 sentences of captions. Content descriptive image-caption datasets are essential tools for advancing the interface between visual content and linguistic descriptions, enhancing the ability of MLLMs to interpret and generate nuanced narratives from images. These datasets vary considerably in scope and methodology but share the goal of improving the semantic association between images and accompanying text. For example, MS-COCO (Chen et al., 2015) employs images with multiple reference captions to refine evaluation metrics, while Flickr30K (Young et al., 2014) extends the variety of descriptions to bolster semantic inference capabilities.
In addition, datasets like Visual Genome (Krishna et al., 2017) provide comprehensive annotations beyond simple captions, including objects, attributes, and relationships, to facilitate complex scene understanding. On the other hand, the AI Challenger Captions (Wu et al., 2017) dataset specifically addresses the need for non-English language representation in image captioning, offering extensive annotations in Chinese to bridge the semantic gap between low-level visual features and high-level conceptual descriptions.
Narrative and textual comprehension are further introduced by datasets like VIST (Huang et al., 2016), which emphasizes storytelling through visual sequences. TextCaps (Sidorov et al., 2020) incorporates reading comprehension into captioning. Collectively, these datasets not only propel advancements in automated image captioning but also tackle broader challenges in AI’s capability to process and generate meaningful visual and textual data, significantly enhancing context-aware machine understanding.
Interleaved Image-Text Documents.
A fundamental contrast between image-text caption datasets and interleaved image-text document datasets lies in their composition: image-caption pair datasets typically consist of a single image accompanied by multiple closely related captions, whereas interleaved image-text datasets consist of a text document interspersed with several illustrative images, wherein the correlation between the images and the accompanying text tends to be relatively lower. Interleaved image-text document datasets commonly used include M3W (Alayrac et al., 2022), MMC4 (Zhu et al., 2024), and OBELICS (Laurençon et al., 2024), which play crucial roles in advancing MLLMs (McKinzie et al., 2024). M3W (Alayrac et al., 2022) leverages data from 43 million web pages, integrating images with text to train models in few-shot learning scenarios. MMC4 (Zhu et al., 2024), an extension of the C4 corpus, aligns images with texts using CLIP features, resulting in a dataset containing over 101.2 million documents aimed at enhancing vision and language models within context. OBELICS (Laurençon et al., 2024), sourced from Common Crawl, comprises 141 million web pages and concentrates on high-quality multimodal content, filtering data to improve model training in comprehending and generating complex multimodal interactions.
Visual Question Answer (VQA) Image-Caption Datasets
Visual question answer (VQA) image-caption datasets facilitate advanced research in MLLMs by allowing models to interpret images and answer related questions. These datasets typically consist of images paired with corresponding questions and answers, aiming to develop and evaluate models’ ability to understand visual content and provide accurate responses. For example, VQAv2.0 (Goyal et al., 2019) expands on its forerunner by introducing complementary images that prompt different answers to the same question, effectively addressing biases and increasing the dataset’s diversity. Visual-7W (Zhu et al., 2016) extends the scope by incorporating questions across multiple dimensions—what, where, when, who, why, how, and which—linked to specific objects within images. ST-VQA (Biten et al., 2019) integrates scene text into the VQA framework, making it possible to answer text-based questions from visual data. Shikra-RD (Chen et al., 2023j) leverages advanced language models to annotate images with relational descriptions, enhancing image comprehension. OCR-VQA focuses on reading text from book covers to answer related questions, thus combining elements of optical character recognition and VQA. DocVQA (Mathew et al., 2021) targets document images for extractive question answering, emphasizing precision in answers derived from visible text. A-OKVQA (Marino et al., 2019) introduces questions that demand commonsense and world knowledge, pushing the envelope on reasoning capabilities required from AI systems. TextVQA (Singh et al., 2019) specifically challenges models to read and understand text within images to respond accurately, and GQA (Hudson and Manning, 2019) promotes advanced reasoning over detailed visual scenes annotated with comprehensive scene graphs. These datasets collectively advance the field by enabling more nuanced interactions between AI models and the rich content within images, aiming for greater depth in the understanding and contextual integration of visual and textual data.
3.3.2. Video Datasets
Video datasets are comprehensive collections of video clips accompanied by associated annotations or labels. These datasets are meticulously designed to facilitate the training and evaluation of models for a wide range of video-related tasks, including video classification, object detection, video question answering, action detection and video captioning.
In the realm of video datasets, MSR-VTT (Xu et al., 2016), WebVid-2M, WebVid-10M (Bain et al., 2021), and VTP (Alayrac et al., 2022) are three prominent examples commonly utilized for pre-training MLLMs. Specifically, the MSR-VTT (Bain et al., 2021) dataset features 10,000 video clips totaling 41.2 hours, each annotated with multiple descriptive sentences. Similarly, the WebVid series offers millions of web-sourced video clips with accompanying captions, expanding from 2.5 million pairs in WebVid-2M to 10 million pairs in WebVid-10M (Bain et al., 2021), covering a vast array of 13,000 hours of video content. The VTP (Alayrac et al., 2022) dataset further enriches the field with 27 million short video-text pairs sourced from a select few high-quality websites. Moreover, the recently introduced Panda-70M (Chen et al., 2024c) dataset contains a massive 70 million high-quality video-caption pairs, designed to enhance the training of high-performance MLLMs. Collectively, these datasets are integral to advancing the capabilities of MLLMs in understanding and interacting with video content.
3.3.3. Audio Datasets
Audio datasets are curated collections of sound recordings along with associated annotations or labels. These datasets are specifically designed to support the training and evaluation of models for various audio-related tasks, such as speech recognition, sound event detection, music classification, and speaker identification. Audio datasets vary significantly in size, quality, and scope, with some focusing on specific languages or acoustic environments, while others strive for diversity to ensure generalizability across different audio processing scenarios. Notable examples include AISHELL-2 (Du et al., 2018), a Mandarin Chinese speech corpus with over 1000 hours of data from multiple regions; WavCaps (Mei et al., 2023), an extensive English audio captioning dataset featuring approximately 400,000 clips; and VSDial-CN (Chen et al., 2023b), a multi-modal dataset derived from VisDial, encompassing visual data alongside related dialogues and captions, tailored for Automatic Speech Recognition (ASR) systems. These datasets are instrumental in advancing the capabilities of language models, enhancing their ability to understand and generate human-like responses in various audio-visual contexts.
3.3.4. 3D Datasets
The development of MLLMs benefits from the utilization of diverse 3D datasets, which provide comprehensive environmental and structural data critical for tasks like scene understanding and semantic segmentation. Notably, the ScanNet (Dai et al., 2017) dataset offers an extensive collection of RGB-D video data across 1,513 indoor scenes, annotated with 3D camera poses, surface reconstructions, and semantic segmentations, totaling over 2.5 million views. Similarly, the S3DIS (Armeni et al., 2017) dataset includes point clouds from six large indoor areas, encompassing 271 rooms, with detailed semantic annotations for each point. The Structured3D (Zheng et al., 2020a) dataset provides a vast repository of 3,500 home designs, encompassing 21,835 rooms detailed with elements such as object geometry, materials, and textures, tailored for analyses in 3D reconstruction and interior design. Together, these datasets present a rich framework for training and evaluating the capabilities of LLMs in interpreting and generating three-dimensional spatial data.
4. Data-centric pre-training
The pre-training stage of multimodal large language models (MLLMs) is crucial for developing models that can effectively process and generate information across multiple modalities. This stage can be divided into two distinct phases, each focusing on specific aspects of the model’s architecture and training objectives.
The first phase involves two parallel processes: pre-training the LLM backbone (such as Vicuna (Chiang et al., 2023) and LLaMA2 (Touvron et al., 2023)) using text-only datasets, and pre-training the modality encoders (such as ViT (Dosovitskiy et al., 2020) and CLIP-ViT (Radford et al., 2021) for visual encoding, C-Former (Chen et al., 2023b) and HuBERT (Hsu et al., 2021) for audio encoding, and ULIP-2 (Xue et al., 2022, 2023) for 3D point cloud encoding) using pairs of data from different modalities, such as image-text or video-text pairs. This phase aims to establish a strong foundation for the model’s understanding of both textual and non-textual information.
The second phase builds upon the knowledge acquired in the first phase by further training the LLM and modality encoders using a mixture of multimodal data. During this phase, the input projector (such as a linear projector, cross-attention mechanism, Q-Former (Li et al., 2022b), or P-Former (Jian et al., 2024)) is trained to effectively map the features extracted by the modality encoders into the LLM’s embedding space. This mapping allows for a unified representation of textual and non-textual information within the model’s architecture.
Some researchers consider employing a selective training approach during the second phase, which involves keeping certain components of the model frozen while training specific parts. This approach helps to preserve the knowledge acquired during the first phase, allows for targeted adaptation to specific multimodal tasks, and reduces computational costs.
4.1. Domain Mixture
The performance of a language model (LM) is significantly influenced by the composition of its pre-training data, which often includes sources such as Wikipedia, books, and web text (Xie et al., 2024a, b).
Previous methods typically select domain weights heuristically or optimize them using downstream tasks (Du et al., 2022). However, these approaches can be suboptimal or costly, requiring the training of LMs for different sets of domain weights and potentially leading to overfitting to specific downstream tasks. DoReMi (Xie et al., 2024a) offers a solution by optimizing domain weights using proxy models that are 30 times smaller than the target LLM, without requiring knowledge of downstream tasks.
In the realm of video understanding, there is a consensus on the necessity for diverse datasets (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a). Xu et al. (2023c) highlight the use of hierarchical multi-label classification models (Giunchiglia and Lukasiewicz, 2020) to ensure dataset balance and comprehensiveness across multiple dimensions, thereby optimizing the model’s ability to understand and process various video domains. Such diversity is instrumental in training more versatile and universally applicable video-language models.
4.2. Modality Mixture
When pre-training MLLMs, determining the optimal proportions of multimodal data is crucial for enhancing the model’s performance across different tasks.
MM1 (McKinzie et al., 2024) explores the balance between image-caption pairs, interleaved image-caption documents, and text-only data for vision-language pre-training. They found that a ratio of 5:5:1 for caption/interleaved/text data yields the best overall performance on text-only tasks as well as zero-shot and few-shot image-text tasks. Additionally, incorporating synthetic data (Lai et al., 2023b) improves the model’s few-shot learning capabilities for image-text tasks.
For the development of MLLMs tailored for video understanding, the strategic integration of both video-text and image-text pairs during the pre-training phase is pivotal. This approach, highlighted in recent research (Lin et al., 2023b; Zhang et al., 2023; Luo et al., 2023; Jin et al., 2023; Li et al., 2023g; Han et al., 2023; Li et al., 2023c, h; Wang et al., 2023a; Chen et al., 2023c), enhances the model’s capabilities by leveraging the complementary nature of these data types.
Integrating labeled image datasets like COCO with video data broadens the variety of training datasets and addresses the scarcity of high-quality video-text resources. Studies (Lin et al., 2023b; Jin et al., 2023) indicate that training on both images and videos enhances models’ ability to understand static and dynamic visual information without needing task-specific adaptations. Adding a temporal modeling module to the vision encoder bridges the temporal dynamics of videos with the static nature of images, fostering cohesive visual understanding across different types of visual media.
However, treating images as single-frame videos (Zhang et al., 2023; Luo et al., 2023) or pseudo-videos may weaken the model’s ability to grasp the temporal aspects of video sequences. An imbalanced mix could also bias the model towards one modality, complicating its learning process. Therefore, determining the optimal mix ratio of image-text pairs and video-text pairs is crucial for improving the model’s comprehensive understanding capabilities, ensuring a balanced approach to learning from both static images and dynamic videos.
Additionally, video LLMs are increasingly being trained through multiple branches (Chen et al., 2024b; Zhang et al., 2023; Lyu et al., 2023; Shu et al., 2023b; Han et al., 2023; Sun et al., 2023c; Chen et al., 2023c), such as vision-language, audio-language, and subtitle-language. These methods leverage the correlations among different modalities to broaden the models’ understanding and learning capabilities. Recent studies (Shu et al., 2023b; Chen et al., 2024b; Han et al., 2023) have shown that combining audio, subtitles, and visual data during training enhances performance across various video understanding benchmarks. This modular training approach also offers the flexibility to pre-train with partial data (Han et al., 2023; Sun et al., 2023c), such as visual-only or audio-only, particularly when comprehensive multimodal data is not available.
4.3. Quality Selection
Due to the diverse distribution of data, training a large model with all available data is not always optimal. Therefore, data selection becomes essential, offering benefits such as reduced training time and energy consumption (Gadre et al., 2024). Unlike pure text datasets, data selection for multimodal datasets must consider the alignment between different modalities, in addition to metrics like perplexity used for single-modality evaluation (Wang et al., 2024d).
Data selection methods can be categorized into two types: active learning-based and pre-training selection. Active learning-based methods, such as CiT (Xu et al., 2023a), use data proxies to dynamically select training data during the training process, achieving significant acceleration effects. Pre-training selection methods, like Datacomp (Gadre et al., 2024), evaluate and select all data before training begins.
Regarding data selection criteria, methods can be distribution-agnostic, focusing solely on individual data point quality, or distribution-aware, considering the overall data distribution. Distribution-agnostic methods include training a model with the top 30% of data ranked by CLIP score, which can significantly improve results (Gadre et al., 2024). More comprehensive metrics have also been developed, showing better performance than using CLIP score alone (Wang et al., 2024d). For instance, Mahmoud et al. (2024) use the difference between original and synthetic captions generated by a small model to evaluate image-text pair alignment.
Given the limited scope of purely distribution-agnostic work, combining distribution-agnostic and distribution-aware methods is often more effective. For example, using CLIP-score-based filters alongside image-based filters can outperform either method alone on large datasets (Wang et al., 2024a; Gadre et al., 2024).
5. data-centric adaptation
Adaptation is crucial for aligning pre-trained multimodal large language models (MLLMs) with specific tasks and user preferences.
Self-supervised pre-training provides LLMs with a broad understanding of textual and multimodal information, while supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) are essential for adapting these models to excel in targeted applications and adhere to user preferences and societal values.
Data-centric SFT involves training models on carefully curated datasets across multiple modalities. In contrast, data-centric RLHF focuses on collecting human judgments to guide the model toward generating desirable and ethically aligned responses. Emphasizing data-centric approaches highlights the importance of high-quality, domain-specific, and multimodal datasets in the successful adaptation of MLLMs.
Table 1. A detailed list of datasets and processing methods for different models with image modality during fine-tuning stage. * indicates the dataset is newly generated using certain method within a respective model, while the other datasets (without *) serve as the original data sources for that model. - denotes directly using original data sources without additional processing. #Examples denotes the statistics for each dataset.
Models Time Finetuning Datasets Processing Method #Examples
Flamingo(Alayrac et al., 2022) 2022.4 VQAv2 - #images: 265K    #questions: 1.4M
VizWiz #image-question pairs: 33.8K
TextVQA #images: 28.4K    #questions: 45.3K
BLIP-2(Li et al., 2023f) 2023.01 COCO - #images: 330K    #captions: 5 per image
NoCaps #images: 15.1K    #captions: 11 per image
VQAv2 #images: 265K    #questions: 1.4M
Flickr30K #images: 31.7K    #captions: 5 per image
LLaVA(Liu et al., 2024b) 2023.04 COCO GPT-4, Prompt, Manual examples #images: 330K    #captions: 5 per image
LLaVA-Instruct-150K* #image-text pairs: 158K
MiniGPT-4(Zhu et al., 2023) 2023.04 CC12M Manual verifying and refining, ChatGPT, Random selecting #image-text pairs: 12M
cc_sbu_align* #image-text pairs: 3.5K
mPLUG-Owl(Ye et al., 2023b) 2023.04 LLaVA-Instruct-150K - #image-text pairs: 158K
InstructBLIP(Dai et al., 2024) 2023.05 ScienceQA - #image-question pairs: 10.3K
OCR-VQA #images: 207K    #image-question pairs: 1M
OKVQA #images: 14K    #questions: 14K
A-OKVQA #images: 23.6K    #questions: 24.9K
PaLI-X(Chen et al., 2023a) 2023.05 COCO - #images: 330K    #captions: 5 per image
NoCaps #images: 15.1K    #captions: 11 per image
TextCaps #images: 28.4K    #captions: 5 per image
VizWizCap #images: 39.1K    #captions: 5 per image
Screen2Words #images: 22.4K    #captions: 5 per image
WidgetCap #images: 21.7K    #captions: 162K    #widgets: 61.2K
Shikra(Chen et al., 2023j) 2023.06 LLaVA-Instruct-150K GPT-4, Prompt, Sampling ratio: 0.5 #image-text pairs: 158K
Flickr30K #images: 31.7K    #captions: 5 per image
Shikra-RD* #qustion-answer pairs: 5.9K
DLP(Jian et al., 2024) 2023.07 COCO - #images: 330K    #captions: 5 per image
ChatSpot(Zhao et al., 2023c) 2023.07 Visual Genome Unification, GPT-4, Prompt #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
MiniGPT-5(Zheng et al., 2023) 2023.10 VIST Flexible framework for various task, Placeholders for images -¿ Text prompts #images: 210K    #stories: 50K
MMDialog #images: 1.53M    #texts: 1.08M    #turns: 4.92M
LLaVA-1.5(Liu et al., 2023c) 2023.10 RefCOCO Merging and concatenating, Splitting, Sampling #images: 20K    #captions: 142K
GQA #images: 113K    #questions: 22M
OCR-VQA #images: 207K    #image-question pairs: 1M
TextCaps #images: 28.4K    #captions: 5 per image
Visual Genome #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
llava_v1_5_mix665k* #image-text pairs: 665K
MiniGPT-v2(Chen et al., 2023m) 2023.10 LLaVA-Instruct-150K Object parsing and grounding, Selecting captions #image-text pairs: 158K
Flickr30K #images: 31.7K    #captions: 5 per image
CogVLM(Wang et al., 2024c) 2023.11 LLaVA-Instruct-150K Translation into Chinese, Correcting and retranslation #image-text pairs: 158K
cc_sbu_align #image-text pairs: 3.5K
CogVLM-SFT-311K* #images: 155K    #image-text pairs: 311K
DRESS(Chen et al., 2023i) 2023.11 LLaVA-Instruct-150K Partition the multi-turn into separate turns, LLM filtering, LLM-Human-in-the-Loop process #image-text pairs: 158K
COCO #images: 330K    #captions: 5 per image
VLSafe* #image-text pairs: 5.8K
VILA(Lin et al., 2023a) 2023.12 llava_v1_5_mix665k - #image-text pairs: 665K
ShareGPT4V(Chen et al., 2023e) 2023.11 COCO GPT-4, Data-specific prompt, Data replacement #images: 330K    #captions: 5 per image
LAION2B-en #image-text pairs: 2.32B
CC3M #image-text pairs: 3.3M
SBU Captions #image-text pairs: 1M
SAM #images: 11M    #segmentation masks: 1.1B
TextCaps #images: 28.4K    #captions: 5 per image
WikiArt #images: 81.4K    #artist classes: 129    #genre classes: 11    #style classes: 27
llava_v1_5_mix665k #image-text pairs: 665K
ShareGPT4V-1.2M* #image-text pairs: 1.2M
GLaMM(Rasheed et al., 2023) 2023.11 RefCOCO Re-purposing of data sources, Manually annotation #images: 20K    #captions: 142K
RefCOCO+ #images: 20K    #captions: 141K
RefCOCOg #images: 25.8K    #captions: 95K
Visual Genome #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
LLaVA-Instruct-150K #image-text pairs: 158K
GranD-f* #image-text pairs: 214K
PVIT(Chen et al., 2023h) 2023.08 GQA ChatGPT, Prompt, Multi-turn data #images: 113K    #questions: 22M
VCR #images: 110K    #questions: 290K
COCO #images: 330K    #captions: 5 per image
Visual Genome #images: 108K    #region descriptions: 5.4M    #image-question pairs: 1.7M
COCO-Text #images:63.6K    #labeled text regions: 173K
PVIT* #image-text pairs: 13.7M (Stage 1)    #image-question pairs: 10.4K (Stage 2)
TextMonkey(Liu et al., 2024d) 2024.03 COCO-Text Structured data (documents, tables, charts), 5% of pre-train data, #images:63.6K    #labeled text regions: 173K
TextOCR #images:28.1K    #texts: 903K
HierText #images:11.6K    #texts: 1.2M
TextVQA #images: 28.4K    #questions: 45.3K
MLT #image-text pairs: 10K
ChartQA #image-text pairs: 20.8K    #tables: 20.8K
DocVQA #images: 12K    #questions: 50K
InfoVQA #images: 5K    #questions: 30K
Monkey_Data* #image-text pairs: 409.1K
Table 2. A detailed list of datasets and processing methods for different models with video modality during fine-tuning stage. * indicates the dataset is newly generated using certain method within a respective model, while the other datasets (without *) serve as the original data sources for that model. - denotes directly using original data sources without additional processing. #Examples denotes the statistics for each dataset.
Models Time Finetuning Datasets Processing Method #Examples
VideoChat(Li et al., 2023c) 2023.05 WebVid GPT-4, Prompt, Randomly choosing #video-text pairs: 10M
VideoChat* #video-description pairs: 7K    #video-conversation pairs: 4K
Video-LLaMA(Lin et al., 2023b) 2023.01 VideoChat - #video-description pairs: 7K    #video-conversation pairs: 4K
Video-ChatGPT(Maaz et al., 2023) 2023.01 ActivityNet Human-assisted annotations, GPT, Semi-automatic annotation #videos: 20K    #texts: 100K
VideoInstruct100K* #video-instruction pairs: 100K
Table 3. A detailed list of datasets and processing methods for different models with audio modality during the fine-tuning stage. * indicates the dataset is newly generated using certain methods within a respective model, while the other datasets (without *) serve as the original data sources for that model. The datasets in other modals used for generating multi-modal datasets (including audio) are also shown in this table. - denotes directly using original data sources without additional processing. #Examples denote the statistics for each dataset.
Models Time Finetuning Datasets Processing Method #Examples
BuboGPT(Zhao et al., 2023a) 2023.07 Clotho GPT-4, Prompt #audios: 5K    #captions: 24K
VGGSS #video-audio pairs: 5K
Clotho-Detail* #audio-caption pairs: 3.9K    #tokens: 207K
VGGSS-Instruction-Tuning* #audio-image-caption pairs: 5.1K
X-LLM(Chen et al., 2023b) 2023.05 AISHELL-2 Manually selecting, ChatGPT #reading-speech audios: 1000 hours
VSDial-CN #automatic speech recognition (ASR) samples: 1.2M
cc_sbu_align (Image-Text) #image-text pairs: 3.5K
ActivityNet (Video-Text) #videos: 20K    #texts: 100K
NExT-GPT(Wu et al., 2023a) 2023.09 WebVid (Video-Text) External resources, GPT-4, Prompt #video-text pairs: 10M
AudioCaps #audio-text pairs: 46K
CC3M (Image-Text) #image-caption pairs: 3.3M
T2M* #images: 4.9K    #videos: 4.9K    #audios: 4.9K    #instances: 14.7K
MosIT* #images: 4K    #videos: 4K    #audios: 4K    #instances: 5K
X-InstructBLIP(Panagopoulou et al., 2023) 2023.11 AudioCaps Automatic generation, Prompt #audio-text pairs: 46K
AudioCapsQA* #audio QA pairs: 25.4K    #unique questoins: 10.8K
DisCRn* #audio-video pairs: 8.8K
Qwen-Audio(Bai et al., 2023a) 2023.11 Self-constructed Manual annotation, GPT-3.5, Data Mixing #audio-text pairs: 20K
5.1. Data-Centric Supervised Finetuning
Supervised fine-tuning (SFT) has become an indispensable technique for adapting MLLMs to specific domains and tasks. By leveraging carefully curated instruction-formatted datasets that encompass multimodal information, these models can be guided to acquire the necessary knowledge and abilities to excel in targeted applications.
During the SFT stage, various strategies have been explored, including fine-tuning one or all of the three primary pre-trained components: the LLM backbone, the modality encoders, and the input projector. This selective approach allows for a balance between adapting the model to specific tasks and preserving the generalized knowledge acquired during pre-training.
To facilitate effective supervised fine-tuning, diverse datasets that align with the target domain and task are used. These datasets typically include a combination of multimodal instruction-response pairs and text-only SFT data. By exposing the MLLMs to this rich variety of data, the model can develop a comprehensive understanding of the target domain and acquire the necessary skills to perform the desired tasks effectively.
The curation of self-supervised fine-tuning datasets usually includes four steps:
  • Collect X-text pairs from various data sources, similar to the pre-training stage as introduced in Section 3.1.
  • Process the data for further use, including filtering, deduplication, translation, etc.
  • Construct the processed X-text pairs into instruction-response form, which involves designing the instruction and converting captions into answers.
  • Select high-quality instruction-response pairs based on the fine-tuning needs. This step is not always necessary but can enhance the dataset’s quality.
Following these steps, supervised fine-tuning can significantly enhance the model’s performance in specific domains and tasks.
5.1.1. SFT Data Collection and Processing
The sources of supervised fine-tuning (SFT) datasets are similar to those of pre-training datasets, consisting of X-text pairs, as introduced in Appendix A.1. We list commonly used SFT models and their data constitution in Table 1, Table 2, and Table 3 for image, video, and audio modalities, respectively.
After collecting these data, several general approaches are employed to improve data quality, including filtering, deduplication, and data enhancement. These steps are similar to the data processing methods used in the pre-training stage described in Section 3.2.
In addition to these general data processing approaches, specific methods aim to improve data quality for supervised fine-tuning tasks. The main goal is to add more textual information to the datasets or improve the quality of captions. These methods, while akin to the data enhancement techniques mentioned in Section 3.2, are more focused on SFT data generation. For example, MiniGPT-4 (Zhu et al., 2023) curated detailed image description datasets for vision-language adaptation by leveraging their pre-trained model with instructions such as ”Describe this image in detail,” ensuring generated sentences contained more than 80 tokens. They also used ChatGPT to refine the descriptions and manually checked for quality, obtaining 3,500 high-quality image-text pairs. Similarly, ShareGPT4V (Chen et al., 2023e) used GPT-4 Vision to generate 100,000 high-quality captions that included world knowledge, spatial information, aesthetic evaluation, and more. These captions were directly used to generate SFT instruction-response pairs.
5.1.2. Instruction-Response Pairs Generation
After collecting and processing the original X-Text datasets, the next step is to generate the instruction-response pairs datasets based on the original datasets. The supervised fine-tuning stage is typically aiming to enhance the model’s ability on downstream tasks. For this reason, previous work consider to generate instruction-response pairs for different downstream tasks, including descriptive tasks, question-answer tasks, reasoning tasks, and classification tasks. Based on task categorization, previous work selects different source datasets and uses various methods to generate the instruction-response pairs. In this section, we first introduce the differences between each downstream task, and then we summarize representative methods of generating instruction-response datasets for each type of downstream task.
Captioning Instruction-Response Datasets.
The simplest tasks are the captioning ones. For this type of task, images (or other modality data) are usually provided, and the MLLMs are asked to give out description of the data. The instruction-response datasets designed to enhance the model’s captioning ability are relatively easy to construct. These kinds of instructions are usually one question asking for the description of the X modality data. Specifically, for image data, MiniGPT-4 (Zhu et al., 2023) construct description instruction tuning by randomly sample instructions such as ”Describe this image in detail”, and their modified high-quality captions described in Section. 5.1.1 are directly used as the response. While LLaVA (Xu et al., 2024) considers designing two types of description instructions, brief description and detailed description and uses GPT-4 (OpenAI, 2023) to generate the answers. For brief descriptions, they add keywords such as ”concisely”, ”brief”, and ”short” in their questions, and for detailed descriptions, they add keywords such as ”descriptive” and ”comprehensive”. ShareGPT4V (Chen et al., 2023e) replaces the detailed description data in LLaVA with its GPT4-Vision generated high-quality captions. X-InstructBLIP (Panagopoulou et al., 2023) further uses these descriptive prompts not only on image-text datasets but also on audio-text, video-text, and 3D-text datasets as a description instruction.
Question Answer Instruction-Response Datasets.
Question answer task requires the model to answer questions based on the given information on the X modality. This kind of task includes simple questions answers, and multiple choices. The Most well-known question-answer task is the visual question answer (VQA) task for visual modality. For VQA tasks, previous research on computer vision has developed a large amount of VQA datasets, and recent work on MLLMs leverages these VQA datasets for constructing instruction-response datasets. For example, LLaVA-1.5 (Liu et al., 2023c) leverage GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), OKVQA (Marino et al., 2019) and VQAv2 (Goyal et al., 2019) to enrich their instruction response datasets. They format these datasets in a uniform template as ”Answer the question using a single word or phrase.” TextVQA (Singh et al., 2019) is another source used for constructing instruction-response datasets which contain text information in images. M3IT (Li et al., 2023l) utilize VQA datasets, such as VQA-v2 (Goyal et al., 2019), Shapes VQA (Andreas et al., 2016) and DocVQA (Mathew et al., 2021), to develop a multimodal instruction tuning dataset that contains 2.4 million instances. In addition to visual question-answer tasks, instruction tuning to improve model performance in the video question answer task also leverages previous video question answer datasets, such as MSRVTT-QA (Xu et al., 2017), iVQA (Yang et al., 2021), MSVD-QA (Xu et al., 2017), and ActivityNet-QA (Yu et al., 2019).
Reasoning Instruction-Response Datasets
Shikra-RD (Chen et al., 2023j) focuses on enhancing the models’ ability for referential dialogue problems, i.e., users point to specific areas and ask questions. They collect public VQA datasets, image captioning data, and several datasets with existing positional annotations to build instruction tuning datasets. Leveraging captions from the Flickr30K (Young et al., 2014) image-caption dataset and using GPT-4, they obtain high-quality RD annotations. PVIT (Chen et al., 2023h) converts the traditional VQA datasets such as GQA and OCR into instruction-response datasets with region-level information. MiniGPT-V2 (Chen et al., 2023m) designs a multi-task instruction template, following the conversation template from LLaMA-2 (Li et al., 2023g). They add task identifier tokens and spatial location representations to identify the task and spatial information in the image. They build five types of instruction-response datasets, including LLaVA-Instruct-150K, grounded image captions from Flickr30K (Young et al., 2014) with direct instruction and object parsing, multi-round conversations, and text-only unnatural instruction datasets (Honovich et al., 2022b). CogVLM (Wang et al., 2024c) provides an IT dataset called CogVLM-SFT-311K. To construct this dataset, they manually select high-quality IT data from MiniGPT-4 and integrate it with LLaVA-Instruct-150K, translating it into Chinese. They then manually correct any noise in their collected instruction tuning datasets and translate them back into English. In their model training, they also leverage VQA datasets and bounding box instruction tuning datasets generated from four types of datasets: grounded captioning (GC) datasets with box-noun phrase pairs in each image, referring expression generation (REG) datasets with box and explanation, referring expression comprehension (REC) datasets with boxes annotated for text, and grounded visual question answering (GroundedVQA) datasets. LLaVA1.5 (Liu et al., 2023c) leverage region level VQA datasets such as Visual Genome (Krishna et al., 2017) and RefCOCO (Kazemzadeh et al., 2014; Mao et al., 2016) in their IT datasets construction.
Other Instruction-Response Datasets
Another type of downstream task is the classification task, where the data is classified into either a label from a candidate pool or an open-label. M3IT (Li et al., 2023l) leverages various datasets for image classification, including ImageNet (Russakovsky et al., 2015), Grounded Object Identification (COCO-GOI)(Lin et al., 2014), COCO-Text(Veit et al., 2016), Image Text Matching (COCO-ITM)(Lin et al., 2014), e-SNLI-VE(Kayser et al., 2021), Multi-modal Fact Checking (Mocheg)(Yao et al., 2023), and IQA(Duanmu et al., 2021). Additionally, for image classification tasks, InstructBLIP (Dai et al., 2024) uses HatefulMemes (Kiela et al., 2020), a binary classification dataset for detecting hateful content in memes, to generate an instruction-response dataset. X-InstructBLIP (Panagopoulou et al., 2023) further extends this approach to audio classification by using AudioSet (Gemmeke et al., 2017). The instructions used are straightforward sentences asking for the classification of the audio, such as ”Classify the following audio”. And the response is the class of the original data.
5.1.3. SFT Data Selection
Supervised fine-tuning is crucial for adapting LLMs with downstream applications. Zhou et al. (2024) found that the primary knowledge of LLMs is acquired during the pre-training stage, and the purpose of instruction tuning is to enable LLMs to learn how to perform well on specific tasks and interact with humans. They determined that only a small set of carefully crafted, high-quality instructions is sufficient to endow LLMs with powerful instruction-following capabilities. Subsequently, various methods have been proposed for data selection to identify high-quality data, improving performance and reducing training costs. These methods are categorized into four types: coreset-based, LLMs-based, gradient-based, and self-instruction-based methods.
Coreset-Based Methods
Coreset provides a compact representation of a larger dataset while preserving its essential characteristics. From a geometric perspective, similar data points in the feature space are close to each other. For example, Sener and Savarese (2017) employs the greedy k-center algorithm to select a coreset and applies it to CNN image classification. In addressing the multi-class classification task, MODERATE CORESET (Xia et al., 2022) scores data points based on their distance to the class center. Data points with scores close to the score median are selected as the coreset, which achieves a balance between discriminative, compressible, and diverse. SIMILAR (Kothawade et al., 2021) picks data points according to the submodular information measures (SMI). A properly selected SMI function can maintain the diversity of selected data and handle imbalance classes, out-of-distribution data, and redundancy. Some algorithms select data points based on model performance, focusing on ”important” samples. For example, Bachem et al. (2017) perform importance sampling based on the upper bound of the sensitivity score to generate a coreset. Generally speaking, data points with high costs are more likely to be selected. Other mertics of performance include GraNd (Paul et al., 2021), least confidence (Wang and Shang, 2014), and etc.
Given that Coreset techniques have demonstrated promising results both theoretically and empirically, it’s feasible to apply these algorithms to the data used for training LLMs. Chen et al. (2023k) employ the K-greedy algorithm to identify just 0.5% of the core data for fine-tuning a pre-trained language model, achieving performance only 1-2% lower than that obtained using the entire dataset. Similarly, the approach proposed by Das and Khetan (2023) involves selecting data that represents the easiest and most challenging examples, a strategy proven effective in (Sorscher et al., 2022). By utilizing only 32.5% of carefully selected data, they achieved state-of-the-art results.
LLMs-Based Methods
External model-based methods frequently utilize external models to evaluate the quality, diversity, and complexity of data. These models serve various functions, from scoring data to enhancing its characteristics. For instance, Du et al. (2023b) leverage DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023f) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023e) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023c) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
For Multimodal instruction tuning datasets, MLLMs are commonly used to obtain or select high-quality instruction tuning data. LLaVA1.5 (Liu et al., 2023c) is the pioneer in leveraging text-only GPT-4 to expand COCO (Lin et al., 2014) bounding box and caption dataset to a multimodal instruction-following dataset. Following LLaVA, ShareGPT-4v (Chen et al., 2023e) use GPT-4v to improve part of LLaVA’s instruction tuning data.
Gradient-Based Methods
Neural Networks are typically trained through gradient-based optimization techniques. Consequently, researchers have been exploring gradient-based methods to use a subset of data to approximate the gradient effectively. The EL2U score was introduced as a method to evaluate how removing a single data point affects the gradient (Paul et al., 2021). However, the EL2U score is limited to single-task learning scenarios. To address multi-task learning, particularly in NLP tasks, another approach has been developed (Attendu and Corbeil, 2023). This method applies the EL2U score to each individual task and then combines these scores using the L2 norm to accommodate the complexities of multi-task learning. Another notable work is (Xia et al., 2024), which adapts existing influence formulations to work with the Adam optimizer. It only uses a few steps of gradients to compute gradient features efficiently and then stores the compressed features in a gradient datastore for efficient data selection.
Self-Instruction-Based Methods
For self-instruction-based methods, the evaluation and selection of data do not require the involvement of any external models. Li et al. (2023m) proposes a self-instruction method for LLMs to autonomously identify and select challenging data. Li et al. (2023d) identifies high-quality data as those that, when included as part of a few-shot example, can enhance the model’s performance. Kung et al. (2023) proposes a novel task-level uncertainty metric that measures the sensitivity of LLMs to instruction perturbations for a task, to identify high-quality data, and combines this with the concept of active learning to iteratively select data. Liu et al. (2024c) scores data at three granular levels: token, sentence, and model. Each level’s scoring is calculated based on the previous level, with the model-level score serving as the final score.
5.2. Data-Centric Human Preference Alignment
To align human preferences with language models, previous work has considered constructing instruction-response datasets that respect human preferences. These datasets usually involve three main elements. RLHF aims to use reinforcement learning (RL) to align the model with human preferences (Ouyang et al., 2022). After training the LLM, RLHF collects human feedback to rank Q&A text data, trains a reward model (RM), and then uses RL to fine-tune the LLMs.
The most impactful work of RLHF is InstructGPT (Ouyang et al., 2022). To collect human feedback, InstructGPT first designs prompts manually and from users who use the InstructGPT playgrounds. Then, they process the prompts using heuristic de-duplication methods, such as checking for prompts that share a long common prefix and limiting the number of prompts to 200 per user ID, to protect the user’s sensitive information and improve the quality of prompts.
These prompts encompass multiple tasks such as text generation, question answering, dialogue, summarization, and more. Human taggers are then asked to rank the responses to each prompt, constructing a Q&A dataset used to train the reward model (RM). Subsequently, the fine-tuning stage of the LLM is framed as a reinforcement learning (RL) problem using the trained reward model. In LLaMa2 (Li et al., 2023g), human preferences are further divided into fine-grained aspects, with two separate reward models trained to rank helpfulness and safety, respectively. Further work also aims to improve the models’ generation abilities from these aspects.
For MLLMs, several works have focused on aligning models with human preferences. Based on LLaVA (Liu et al., 2023b), LLaVA-RLHF (Sun et al., 2023a) designed a reinforcement learning from human feedback (RLHF) method to align LLaVA with human feedback. DRESS (Chen et al., 2023i) aims to improve models’ response quality from the perspective of human values, particularly the 3H criteria (Ouyang et al., 2022) (helpfulness, honesty, and harmlessness). They provide two datasets: the Large Vision Language Model with Natural Language Feedback (LVLM_NLF) dataset and the Vision-Language Safety (VLSafe) dataset. To construct the VLSafe dataset, they adopt the LLM-Human-in-the-Loop approach, iteratively creating and filtering the data based on the COCO dataset. The LLM they used is GPT-3.5 Turbo. The final dataset contains 5,874 samples, meeting most requirements for harmlessness alignment and evaluation.
6. evaluation
To offer a thorough data-centric viewpoint for assessing MLLMs, this section starts with two key aspects. First, we review widely-used methods for evaluating data quality. Then, we provide an overview of MLLMs evaluation datasets, highlighting their collection, processing, and characteristics.
6.1. Data Evaluation
Evaluating datasets is essential to ensure their quality and reliability. This process helps identify and correct errors and biases, enhancing the accuracy of model training and predictions. In this section, we summarize recent dataset evaluation metrics from the following perspectives.
6.1.1. Dataset Diversity
Dataset diversity is crucial for developing robust and generalizable machine learning models. A diverse dataset enables the model to manage various scenarios and minimizes the risk of bias. Some studies (Heusel et al., 2017; Sajjadi et al., 2018) require a reference distribution or dataset. For instance, Heusel et al. (2017) measures the Wasserstein-2 distance between two Gaussian distributions: one fitted to the embeddings of the reference sample and the other to the embeddings of the sample being evaluated for diversity.
Other papers use similarity scores to define diversity. For example, Shen et al. (2019); Fomicheva et al. (2020) utilize the average pairwise similarity score or its complement, the average dissimilarity. Fomicheva et al. (2020) propose a two-metric evaluation paradigm using precision and recall, with precision measuring quality and recall assessing diversity in terms of coverage of the reference distribution. The Vendi Score (Friedman and Dieng, 2023; Pasarkar and Dieng, 2023; YEH et al., 2023) stands out from existing diversity evaluation metrics by being a reference-free, flexible, and interpretable metric that measures internal diversity without comparing it to a reference distribution. Its reliance on a user-defined similarity function allows for broad applicability across domains, incorporating correlations between features while being computationally efficient and unsupervised. This makes it a valuable tool for diverse machine learning applications.
6.1.2. Dataset Quality
Dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enables models to learn effectively and make precise predictions. Evaluating data quality can be approached from multiple perspectives. TRUE, as presented by Honovich et al. (2022a), provides a comprehensive study evaluating factual consistency metrics in grounded text generation systems. By standardizing datasets and introducing a meta-evaluation protocol, the study demonstrates the robust performance of large-scale NLI and QG-QA methods across various tasks.
Ensuring dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enables effective learning and precise predictions. There are several approaches to evaluate data quality. The study TRUE (Honovich et al., 2022a) offers a thorough evaluation of factual consistency metrics in grounded text generation systems. By standardizing datasets and introducing a meta-evaluation protocol, it highlights the robust performance of large-scale NLI and QG-QA methods across different tasks. Object Hallucination (Rohrbach et al., 2018) introduces CHAIR (Caption Hallucination Assessment with Image Relevance), which measures the proportion of words in a generated caption that correspond to objects present in the image, using ground truth sentences and object segmentations. This method helps evaluate the hallucination of the data. FAITH SCORE (Jing et al., 2023) assesses the faithfulness of generated answers from large vision-language models (LVLMs). It identifies descriptive statements, extracts atomic facts, and checks their consistency with input images. This structured process ensures that generated answers align closely with the visual content, providing a comprehensive evaluation of faithfulness in LVLM outputs.
6.1.3. Dataset Similarity
In machine learning theory, performance improves when the training data distribution closely matches the testing data distribution. Therefore, it is essential to evaluate data distribution distances both theoretically and empirically. Common metrics for measuring distribution similarity include Euclidean distance, KL-divergence, CORAL Loss (Sun and Saenko, 2016), Wasserstein distance, and MMD distance (Jiang et al., 2022). MAUVE Scores (Pillutla et al., 2023) serve as a comparison measure between pairs of distributions, such as those encountered in the generative modeling of text or images. This metric provides statistical bounds and extensive experiments to demonstrate its effectiveness.
6.2. Evaluation Datasets for MLLMs
In this section, we summarize and compare the commonly used evaluation datasets for MLLMs based on the evaluation tasks.
6.2.1. Captioning Tasks
Captioning datasets act as benchmarks to evaluate a model’s fundamental ability to understand information from different modalities. They can be used to assess the zero-shot or few-shot learning capabilities of MLLMs in multimodality tasks, offering insights into their ability to generalize and adapt to new scenarios with limited training examples.
For image captioning tasks, numerous datasets are available to assess a model’s capability in image understanding. MS-COCO (Chen et al., 2015) is one of the most frequently used datasets for both training and evaluating image captioning. Flickr30K (Young et al., 2014) is also extensively utilized to evaluate models’ capabilities across Zero/Few/Full-shot scenarios. The Karpathy split for MS-COCO and Flickr30K (Karpathy and Fei-Fei, 2015) is commonly adopted for dividing training and testing data in the training and evaluation of MLLMs. Nocaps (Agrawal et al., 2019) was developed to overcome the limitations of MS-COCO by focusing on objects not present in MS-COCO captions, serving as a valuable complement to the MS-COCO dataset.
For video captioning, the MSVD dataset (Chen and Dolan, 2011) is a popular choice for evaluating action recognition and video description tasks (Venugopalan et al., 2015), comprising 2,089 video segments and 85,550 English descriptions, with a standard split provided by Venugopalan et al. (2015). The MSRVTT dataset (Xu et al., 2016) includes 10,000 web video clips totaling 41.2 hours, each annotated with approximately 20 natural sentences. The DiDeMo dataset (Anne Hendricks et al., 2017) contains over 10,000 unedited personal videos, each associated with 3-5 pairs of descriptions pinpointing specific moments. The VATEX dataset (Wang et al., 2019) includes more than 41,250 videos paired with 825,000 captions in both English and Chinese, featuring over 206,000 English-Chinese parallel translation pairs. Additionally, the TVC dataset (Lei et al., 2020) includes both validation and private test components for further evaluation. These datasets are instrumental in evaluating both video captioning and video information retrieval capabilities.
For evaluating audio captioning, the two most commonly used datasets are AudioCaps and Clotho. The AudioCaps dataset (Kim et al., 2019) features 46,000 pairs of audio clips and human-written text descriptions. Meanwhile, the Clotho dataset (Drossos et al., 2020) includes 4,981 audio samples, each lasting between 15 to 30 seconds, accompanied by 24,905 captions that range from 8 to 20 words in length. Both datasets are instrumental in assessing a model’s audio understanding capabilities.
6.2.2. Question Answering Tasks
Evaluation datasets for question answering (QA) are crucial for assessing the performance of MLLMs that handle diverse data types like text, images, and audio. These datasets require models to analyze and combine information from multiple sources to respond to relevant questions accurately.
Visual Question Answering (VQA) evaluation dataset tests the MLLM’s ability to integrate and interpret both text and image data. These datasets contain images paired with questions that require visual understanding to answer. Here are several commonly used VQA evaluation datasets: VQAv2 (Goyal et al., 2019) contains 13 million answers associated with roughly 200,000 images from the COCO dataset. It also contains multi images question answer pairs that can test models’ multi image understanding abilities. GQA (Hudson and Manning, 2019) dataset is designed to address deficiencies in previous visual question answering (VQA) datasets by fostering advanced visual reasoning and reducing biases. It utilizes over 22 million questions generated from 113K real-world images annotated with detailed scene graphs, detailing objects, attributes, and relationships. OKVQA (Marino et al., 2019) was also created to address the limitations of existing VQA benchmarks by focusing on knowledge-based visual question answering. This dataset consists of more than 14,000 questions that require external knowledge to answer, covering various categories such as science & technology, history, and sports. Vizwiz (Gurari et al., 2018) focuses on collecting visual questions from blind users, presenting unique challenges such as image quality and conversational question styles. POPE (Li et al., 2023b) focuses on evaluating object hallucination in vision-language models, using 6,136 binary questions to test if models hallucinate non-existent objects in images from MS-COCO and other sources.
Video question-answering evaluation datasets are developed to assess the capabilities of models in video-to-text translation and overall video comprehension. Key datasets include MSVD-QA (Xu et al., 2017) and MSRVTT-QA (Xu et al., 2017), derived from the MSVD (Chen and Dolan, 2011) and MSRVTT (Xu et al., 2016) video captioning datasets, respectively, which are frequently utilized for video QA tasks. For more intricate web videos, the ActivityNet-QA (Yu et al., 2019) dataset, containing 58,000 QA pairs from 5,800 complex videos, tests models’ abilities to understand complex video content. The TGIF-QA (Jang et al., 2017) dataset, with 165,165 QA pairs from 71,741 animated GIFs, evaluates spatio-temporal reasoning and visual question-answering skills in videos. LSMDC (Rohrbach et al., 2017), which includes 118,114 sentences aligned with clips from 202 movies, is used for assessing video description generation and movie QA abilities. Additionally, MoVQA (Rohrbach et al., 2017), featuring 21,953 manually annotated QA pairs from 100 diverse movies, is aimed at evaluating the understanding of long-form videos over various temporal durations, focusing on complex and extended video content comprehension.
Audio question-answering evaluation datasets are designed to test audio-to-text translation and the ability to understand audio. The ClothoAQA (Lipping et al., 2022) dataset, derived from the Clotho dataset, includes 1,991 audio files, each lasting between 15 to 30 seconds, with six different questions per audio file collected. The MUSIC-AVQA (Li et al., 2022c) dataset contains over 45,000 question-answer pairs extracted from more than 9,000 videos, and over 150 hours of content. It uses 33 question templates across 9 types of questions to support spatio-temporal reasoning in audio-visual scenarios.
6.2.3. Perception and Reasoning Tasks
Perception and reasoning are crucial tests for assessing the general abilities of machine learning language models (MLLMs). TextVQA (Singh et al., 2019) consists of 28,408 images with questions that require reading and reasoning about text in the image. ChartQA (Masry et al., 2022) offers diverse questions based on real-world charts, testing visual and logical reasoning. AI2D (Kembhavi et al., 2016) includes over 5,000 grade school science diagrams with extensive annotations to evaluate diagram interpretation skills. ScienceQA(Lu et al., 2022) presents around 21,000 multimodal questions from science curricula, emphasizing multi-hop reasoning. MathVista (Lu et al., 2023a) comprises 6,141 image-text pairs to benchmark mathematical reasoning in visual contexts. MMVet (Yu et al., 2023) includes 200 images with questions testing large multimodal models across six capabilities. MMBench (Liu et al., 2023a) assesses 20 distinct abilities with 3,000 questions, ranging from object localization to social reasoning. LLaVAW (Liu et al., 2024b) features 24 images with diverse visual content and complex reasoning tests. MME (Fu et al., 2024) focuses on 14 subtasks covering perception and cognition, such as text translation and arithmetic. Lastly, MVBench (Li et al., 2023h) contains 20 video tasks that challenge comprehension of spatial and temporal dynamics, designed to evaluate advanced understanding skills.
There are several datasets designed to evaluate MLLMs’ spatial reasoning abilities. VSR (Liu et al., 2023b) comprises over 10,000 natural text-image pairs with 66 types of spatial relations, making it an excellent resource for evaluating spatial reasoning. The datasets RefCOCO, RefCOCO+, and RefCOCOg (Kazemzadeh et al., 2014; Mao et al., 2016) are crucial for assessing how well models understand and ground natural language expressions within visual contexts, featuring images paired with expressions that describe specific objects or areas. These datasets, which vary in expression collection and annotations, challenge models to integrate vision and language by capturing detailed spatial relationships and contextual nuances. Additionally, the GRIT (Gupta et al., 2022) benchmark addresses seven vision tasks using multiple data sources to enhance referring expression grounding. DocVQA (Mathew et al., 2021) focuses on visual question answering within document images, utilizing a web-based tool to generate its 50,000 questions based on over 12,000 document images. OCR-VQA (Mishra et al., 2019) contains 207,572 images of book covers, with over a million question-answer pairs, testing models’ ability to understand text in document images and perform visual question answering. Lastly, SEED-Bench (Li et al., 2023i) evaluates spatial and temporal comprehension in multimodal contexts, serving as a platform for testing generative comprehension capabilities.
For video reasoning, there exist numerous evaluation datasets. EgoSchema (Mangalam et al., 2024) features over 5,000 very long-form video language understanding questions derived from 250 hours of diverse egocentric video data. VideoChatGPT (Maaz et al., 2023), employed to tackle challenges in video-based conversation models, addresses aspects like temporal understanding, spatial consistency, and contextual comprehension. Both EgoSchema (Mangalam et al., 2024) and VideoChatGPT (Maaz et al., 2023) offer the means to assess models on longer and more intricate video content. Charades-STA (Gao et al., 2017) comprises approximately 10,000 videos, each annotated with temporal activity details across 157 activity categories and multiple video-level descriptions. Additionally, QVHighlights (Lei et al., 2021) contains over 10,000 YouTube videos covering various topics such as everyday activities, travel, and social and political events, facilitating the evaluation of MLLMs’ abilities in detecting moments and generating highlights.
6.2.4. Other Noteworthy Tasks
There are also several noteworthy tasks not mentioned above. One of the traditional tasks is the classification task. For the classification task, InstructBLIP uses HatefulMemes (Kiela et al., 2020), a binary hateful content classification dataset for memes classification. ImageNet-1K (Deng et al., 2009) is also a widely used dataset to test models’ image classification ability. While X-InstructBLIP extends this to ESC50 (Piczak, 2015) for audio classification and ModelNet40 (Wu et al., 2015) for 3D classification. Furthermore, M3IT (Li et al., 2023l) use Something-Something (Goyal et al., 2017) for video action classification. Multimodal dialogue task is different from question answering task because it requires multiple rounds of conversion. TouchStone (Bai et al., 2023b) explores multimodal dialogue through detailed image annotations, addressing the challenge of evaluating LLMs in open-ended dialogues.
7. future direction
In this section, our attention is directed towards research avenues for MLLMs from a data-centric viewpoint. We categorize forthcoming directions based on the structure of our article. We begin by addressing future directions concerning data collecting and processing, followed by pre-training, adaptation, and evaluation aspects.
7.1. Data Processing System for MLLMs.
The processing of data for MLLMs entails numerous intricate steps across various modalities, as outlined in Section 3.2. Recent efforts in preparing data for MLLM training involve processing data autonomously through specialized data pipelines and processing operators (Alayrac et al., 2022; Changpinyo et al., 2021; Bain et al., 2021; Du et al., 2018). While previous work has focused on designing data processing systems for LLMs, such as Data-Juicer (Chen et al., 2023d) and Oasis (Zhou et al., 2023), there remains a gap in the availability of data processing systems tailored specifically for MLLM data. Such systems have to be equipped to handle multi-modal data types, encompassing not only textual data but also images, videos, audio, and 3D data formats.
7.2. Data Quantity Analysis for MLLM Pre-training.
Large language models have shown specific properties such as Large language models exhibit distinct characteristics such as emergent abilities and scaling laws. Previous research has examined emergent phenomena concerning model scale, measured by factors like training compute and the number of model parameters, as well as in relation to evaluation metrics (Wei et al., 2022; Schaeffer et al., 2024). Scaling laws have been investigated concerning model size, data scale, and data quality. OpenAI initially explored the power-law relationship of pre-trained loss with respect to model size, dataset size, and training compute in neural language models (Kaplan et al., 2020). Subsequently, DeepMind introduced a new scaling law demonstrating the optimal allocation of compute resources (Hoffmann et al., 2022). Building on this, Goyal et al. (2024) investigated the trade-offs between dataset quality and quantity, while Ye et al. (2024) examined scaling laws associated with data mixtures. Furthermore, Zhang et al. (2024b) analyzed the impact of these properties, particularly focusing on text-only LLMs. However, there remains a gap in understanding how data quantity influences emergent abilities in MLLMs. Exploring the scaling laws of MLLMs, particularly regarding the evaluation of data quantity, represents an area that warrants further investigation.
7.3. Data Quality Analysis for MLLM Pre-training.
Adjusting the ratio and quantity of pre-training data holds significant potential for enhancing model performance. However, with the abundance of pre-training data available, there arises a critical necessity for efficient data selection algorithms. Evaluating data for large models typically demands the utilization of the models themselves, which can impose significant computational burdens. Hence, the development of Proxy Models becomes crucial, as they offer substantial reductions in computational costs. Moreover, Proxy Models prove instrumental in optimizing the mixing ratio of data from diverse sources. While several approaches like Doremi (Xie et al., 2024a) and Doge (Fan et al., 2023) have been proposed in this regard, there remains a need for the refinement of proxy models and further exploration into the relationship between proxy models and their original counterparts.
7.4. MLLM Data Evaluation.
As elaborated in Section 6.1, despite the introduction of various data evaluation metrics, there remains a notable absence of comprehensive metrics tailored specifically for evaluating multimodal data. The assessment of multimodal data presents heightened challenges owing to its diverse array of data types, often encompassing multiple tasks and modalities (Liu et al., 2023c; Chen et al., 2024c). The quality of such data can be impacted by the individual models corresponding to each modality and their alignments. Hence, it becomes imperative to scrutinize the quality of multimodal data from multiple dimensions.
In addition to relying solely on human-defined features, enhancing data quality can be achieved by leveraging statistical metrics and methodologies, such as scrutinizing distributions. The notion of domain adaptation offers theoretical underpinnings for matching distributions (Jiang et al., 2022). Consequently, assessing distributions emerges as a promising avenue, both in terms of theoretical frameworks and empirical validation.
7.5. Data Quality Improving for MLLM Supervised Fine-Tuning.
Harnessing Instruction Tuning data can substantially bolster a model’s proficiency in adhering to instructions and enhancing its performance across specific tasks. Given the pivotal role of both quality and quantity in optimizing model performance, the exploration of this area emerges as a crucial research endeavor. The evaluation of data for large-scale models necessitates metrics that are model-agnostic, effectively capturing the distinct characteristics of the data. As highlighted in the preceding paragraph, there is a pressing need to devise metrics tailored specifically for instruction tuning data.
An alternative pragmatic approach involves employing Large Language Models (LLMs) for data evaluation, leveraging the extensive knowledge accrued during their pre-training phase (Du et al., 2023b; Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e; Lu et al., 2023c; Parkar et al., 2024). However, utilizing LLMs for data assessment proves to be cost-effective. Despite the success of GPT-based methods for automated data quality evaluation, such approaches often lack interpretability. Hence, further investigation is warranted to gain a deeper understanding of GPT’s efficacy and to delineate the boundaries of employing GPT for data evaluation.
7.6. MLLM Lifelong Learning.
MLLMs often learn from diverse data types sequentially across various training phases (Li et al., 2022b; Liu et al., 2024b, 2023c). Typically, this sequential training starts with initializing the model using pre-trained weights for each modality, followed by pre-training, instruction tuning, and concluding with reinforcement learning from human feedback. It is critical throughout these stages to prevent catastrophic forgetting, ensuring the model retains its foundational language capabilities while acquiring multimodal skills. Therefore, integrating lifelong learning strategies with MLLMs represents a significant and necessary research direction.
8. conclusion
In this survey, we review recent advancements in data-centric multimodal large language models (MLLMs), introducing key concepts, findings, and techniques essential for processing training data for these models. Specifically, our discussion focuses on three crucial aspects of data handling: pre-training, adaptation, and evaluation. For each aspect, we highlight pivotal techniques and insights for data processing that are critical for effectively training MLLMs. Additionally, we provide a comprehensive summary of available data resources for training these models and discuss strategies for utilizing machine learning to enhance the performance of MLLMs. This survey aims to encapsulate the most recent literature on data-centric MLLMs, serving as a valuable reference for both researchers and engineers in the field.
Appendix A appendix
A.1. Commonly Used X-Text Datasets
X-Test datasets are the most commonly used evaluation datasets. Here, we summarize Image-Text, Video-Text, Audio-Text and 3d-Text datasets.
A.1.1. Image-Caption Datasets
In this subsubsection, we provide a comprehensive summary of image caption datasets to help people find the most suitable datasets for MLLMs training.
General Image-Caption Datasets.
General image-caption datasets are datasets that contains image with short caption. Usually these captions are only one sentence or several words, describing the key character of the image.
  • LAION-400M (Schuhmann et al., 2021). LAION-400M is a pioneering dataset for multimodal language-vision model training, comprising CLIP-filtered 400 million image-text pairs sourced from the Common Crawl web archive. These pairs are accompanied by CLIP embeddings and k-Nearest Neighbors (kNN) indices for efficient similarity searches. The dataset was processed through distributed parsing of Common Crawl’s WAT files, extracting images with alt-texts as captions. Images and captions were then filtered based on character length, image size, and cosine similarity of CLIP embeddings to ensure relevance and quality. Additionally, duplicates were removed, and illegal contents were filtered out using CLIP embeddings. Less than 1% of the images were tagged as Not Safe For Work (NSFW), allowing for easy exclusion. LAION-400M enables the broad research community to train and research large-scale multimodal models, previously limited to proprietary datasets. This dataset signifies a crucial step towards democratizing research in multimodal machine learning by providing a vast, diverse, and accessible resource for developing advanced language-vision models.
  • LAION-5B (Schuhmann et al., 2022). LAION-5B is the largest public image-text dataset to date, comprising over 5.8 billion examples derived from the Common Crawl web archive. It includes 2.32 billion English image-text pairs, 2.26 billion pairs in over 100 other languages, and 1.27 billion samples where a language couldn’t be clearly identified, often depicting products or places with captions that, while semantically clear, may contain noise such as SEO keywords. The dataset was assembled using an existing CLIP model to filter CommonCrawl data, focusing on images with alt-text for creating pairs. Preprocessing involved removing pairs with low cosine similarity between text and image embeddings, leading to a significant reduction from the initial collection. LAION-5B facilitates training state-of-the-art language-vision models, offering unprecedented scale for research into zero-shot transfer, robustness, and multimodal learning.
  • LAION-COCO (Schuhmann et al., 2022). LAION-COCO is a dataset featuring 600 million synthetic captions for images from the LAION-5B English subset, utilizing models like BLIP L/14 and CLIP. It was developed to explore the potential of generated captions in training models. The dataset includes the original and top captions alongside alternative ones, all available in parquet files. This unique assembly allows for the examination of synthetic captions’ impact on model training, offering a vast resource for multimodal learning.
  • COYO-700M (Minwoo Byeon, 2022). The COYO-700M dataset comprises 747 million image-text pairs, meticulously curated from CommonCrawl data spanning from October 2020 to August 2021, designed to enhance multimodal learning in vision-and-language research. This extensive collection targets the informative association between alt-text and corresponding images in HTML documents, undergoing rigorous image and text-level filtering to boost quality, reduce redundancy, and eliminate inappropriate content. Tailored for training large-scale foundation models, COYO-700M supports diverse image formats and offers a variety of meta-attributes, making it a crucial resource for advancing machine learning models and applications, particularly in tasks like image captioning and visual question answering. This dataset represents a significant stride in understanding the complex relationship between visual and textual data, facilitating the development of more sophisticated vision-and-language models.
  • DataComp (Gadre et al., 2024). The DataComp challenge provides a testbed for image-text datasets. The dataset proposed in this challenge, known as COMMONPOOL, is a large-scale multimodal dataset constructed from 12.8 billion image-text pairs sourced from Common Crawl, designed to foster innovation in dataset design within the machine learning community. This challenge allows researchers to participate in two tracks: filtering from the provided pool COMMONPOOL datasets or using external data in a ”Bring Your Own Data” (BYOD) approach, ensuring data does not overlap with evaluation sets. This approach aims to enhance the understanding of how dataset curation impacts machine learning model performance, placing a strong emphasis on rigorous dataset design over algorithmic competition.
  • SBU (Ordonez et al., 2011). The SBU captioned photo dataset features over 1 million Flickr images, each with a user-generated caption, developed to bridge visual content with natural language descriptions. This extensive collection is the result of targeted Flickr queries and careful filtering to ensure captions accurately reflect the visual elements of the images. The dataset facilitates the challenging task of automatic image description using simple matching techniques and the incorporation of various image content estimates like objects, attributes, and scenes. This approach has shown promising results in non-parametric image captioning methodologies. Moreover, the SBU dataset introduces a novel performance measure for evaluating image captions, providing a valuable resource for research in automated image understanding and description generation. With its focus on leveraging web-sourced captions for detailed image annotation, the SBU dataset represents a significant advancement in the field of image captioning.
  • Conceptual Captions 3M (CC3M) (Sharma et al., 2018). The Conceptual Captions (CC3M) dataset is constructed by programmatically harvesting and processing image-caption pairs from billions of web pages, focusing on alt-text associated with images. This process includes rigorous filtering based on image quality, relevance, and caption appropriateness, followed by text transformations where specific names and details are generalized to create more universally learnable captions. Comprising approximately 3.3 million pairs, the CC3M dataset is designed to be significantly larger and more diverse than traditional datasets like MS-COCO.
  • Conceptual Captions 12M (CC12M) (Changpinyo et al., 2021). The CC12M dataset was created by relaxing filters used in the CC3M dataset, resulting in a larger set of 12 million image-text pairs for vision-and-language pre-training. This dataset exhibits a longer-tail distribution, capturing a wider range of visual concepts compared to CC3M. By including more diverse and complex concepts, CC12M aims to improve transferability and performance on long-tail recognition tasks in vision-and-language models. The dataset’s creation process involved steps such as image-based and text-based filtering, as well as text transformations to enhance privacy and maintain data quality. Overall, CC12M serves as a valuable resource for addressing challenges in vision-and-language tasks, offering a rich and diverse set of training examples for improved model generalization and performance.
  • Wukong (Gu et al., 2022). The Wukong dataset, containing 100 million Chinese image-text pairs from the web, is significant for Chinese vision-language pre-training (VLP) models. Its preprocessing includes image and text-based filters to ensure quality: images are screened for size and aspect ratio, while texts are checked for language, length, and relevance, discarding overly common or irrelevant descriptions. This careful curation addresses the lack of large-scale Chinese datasets, facilitating the development and benchmarking of VLP models across diverse concepts and applications.
  • ALT200M (Hu et al., 2022). The ALT200M dataset comprises over 200 million image-text pairs sourced from the web, using image alt attributes. Constructed to explore scaling in vision-language pre-training (VLP) for image captioning, the dataset aims to assess the impact of increased model and data sizes on VLP performance. Images and corresponding alt-texts were collected via a pipeline with minimal rule-based filtering, focusing on image size and aspect ratio, while alt-texts were curated based on length and vocabulary derived from English Wikipedia. Named entities in texts were anonymized to maintain privacy. The dataset is designed to facilitate experiments on various model sizes and configurations, particularly in handling large-scale, noisy data, with the goal of improving recognition of diverse and rare visual concepts and achieving new benchmarks in image captioning.
  • LTIP (Alayrac et al., 2022). The Long Text & Image Pairs (LTIP) dataset, consisting of approximately 312 million image and text pairs, is meticulously assembled from various web sources to ensure high-quality and lengthy textual descriptions, surpassing typical datasets in detail. This dataset serves as a fundamental component in training the Flamingo models, which are tailored for multimodal tasks involving both visual and textual data interpretation. Through rigorous quality checks and careful formatting, LTIP is optimized for integration into training systems, significantly enhancing the capability of models like Flamingo to perform complex vision and language tasks. The primary purpose of the LTIP dataset is to enrich these models’ understanding and generation of contextually relevant textual descriptions from visual inputs, thereby boosting their efficacy across diverse vision and language applications.
Content Descriptive Image-Caption Datasets.
Content descriptive image-caption datasets contain longer caption with more descriptive information. Usually each image in content description image-caption datasets contain at least 5 sentences of caption.
  • MS-COCO (Chen et al., 2015). This dataset contains 328k images with a total of 2.5 million labeled instances, the creation of this dataset drew upon extensive crowd worker involvement via novel user interfaces for category detection, instance spotting and instance segmentation. The MS-COCO dataset was created by collecting images from Flickr that contain multiple objects in their natural context, with captions gathered through human subjects on Amazon’s Mechanical Turk. Each image has multiple reference captions, with a subset having 40 reference sentences to improve evaluation metrics’ correlation with human judgment. The dataset includes training, validation, and testing sets. This dataset’s characteristics make it suitable for training and evaluating algorithms for image captioning tasks, addressing the challenge of generating accurate and descriptive captions for diverse and complex visual content of existing LLMs.
  • Flickr30K (Young et al., 2014). The Flickr30K dataset in the article was created by collecting 31,783 images depicting everyday activities and events from Flickr, along with 158,915 captions obtained through crowd-sourcing. This dataset extends a previous corpus and follows similar annotation guidelines to ensure quality control. Each image is described by five annotators, providing diverse descriptions ranging from general situations to specific actions. The dataset’s unique feature lies in the variety of descriptions associated with each image, enabling the induction of denotational similarities between expressions not directly related by syntactic rules. This dataset can be utilized to address challenges in semantic inference tasks by leveraging the visual denotations of linguistic expressions to define novel similarity metrics and enhance understanding of textual descriptions in relation to visual content.
  • WEBLI (Chen et al., 2022). The WEBLI dataset is a vast multilingual image-language dataset comprising 10 billion images and approximately 12 billion alt-texts, constructed from publicly available internet resources to support the training of the PaLI multimodal model. It includes text in over 100 languages and features OCR annotations to enhance text recognition capabilities in images. To ensure high quality, only the top 10% of image-text pairs, totaling around 1 billion, were retained for training the PaLI model. The dataset’s primary purpose is to facilitate the pre-training of large-scale models on a variety of multimodal tasks across multiple languages, thereby improving the models’ performance on complex tasks such as image captioning and visual question answering, and advancing research in image and language processing.
  • ALIGN (Jia et al., 2021). The ALIGN dataset represents a significant expansion over CC3M by utilizing over one billion image-text pairs, compared to CC3M’s 3.3 million. This increase is achieved by simplifying the data collection process; while CC3M underwent rigorous image-based, text-based, and image&text-based filtering to ensure high-quality data, ALIGN limits its preprocessing to only image-based and text-based filtering. By omitting the image&text-based filtering step, ALIGN trades some data cleanliness for scale, demonstrating that the volume of data can outweigh the impact of noise. This approach facilitates the training of more robust models capable of understanding complex visual and linguistic tasks, highlighting a strategic shift towards leveraging the vast, albeit imperfect, data available on the web for advanced representation learning.
  • RefCOCOg (Yu et al., 2016; Mao et al., 2016). RefCOCOg is a dataset created through Amazon Mechanical Turk, featuring 85,474 natural language referring expressions for 54,822 objects across 26,711 MSCOCO images. Images were chosen to have 2 to 4 objects of the same category, ensuring complexity in object referencing. The dataset’s creation involved two sets of workers: one to write expressions for objects in images and another to identify objects based on these expressions. Validity was confirmed through successful object identification, making RefCOCOg a robust resource for tasks requiring detailed understanding of complex, referential language in visual contexts.
  • RefCOCO & RefCOCO+ (Kazemzadeh et al., 2014; Mao et al., 2016; Yu et al., 2016). RefCOCO and RefCOCO+ are datasets generated through the ReferItGame, a two-player game designed to collect natural language expressions referring to objects within images. RefCOCO contains 142,209 expressions for 50,000 objects in 19,994 images. RefCOCO+ includes 141,564 expressions for 49,856 objects across 19,992 images, specifically banning location words in expressions to focus on appearance-based descriptions. This approach aims to gather expressions independent of viewer perspective, providing a rich resource for computer vision tasks requiring detailed object identification based on descriptive language.
  • Visual Genome (Krishna et al., 2017). The Visual Genome dataset vastly enhances the scope of image-text pair research, offering over 108,000 images enriched with detailed annotations, a leap from CC3M’s 3.3 million pairs. It diverges from CC3M’s simpler image-caption pairs by integrating images from the YFCC100M and MS-COCO datasets, supplemented with extensive crowd-sourced annotations. Beyond mere captions, it details over 3.8 million object instances, attributes, and object relationships within images through structured scene graphs. This innovation not only expands the dataset’s volume but also its utility for advanced image understanding, providing a comprehensive framework for analyzing complex visual scenes. Visual Genome aims to advance the field by enabling models to interpret nuanced visual contexts, making it a pivotal resource for progressing in object detection, scene comprehension, and visual reasoning. In essence, Visual Genome represents a significant evolution in image-text datasets, designed to mirror the complexity of the visual world.
  • CC595k (Liu et al., 2024b). The CC595k dataset is derived from the CC3M dataset, refined to include around 595,000 image-text pairs. This selection process involved extracting noun-phrases from captions, focusing on those with a frequency of at least 3 to ensure concept coverage while omitting rare phrases. For highly frequent noun-phrases (over 100 occurrences), a random subset of 100 captions was chosen. This preprocessing aims to balance concept variety against the dataset’s size, optimizing for broad yet manageable representation of concepts within the constraints of fewer image-text pairs.
  • TextCaptions (Sidorov et al., 2020). The TextCaps dataset was created with the aim of exploring the task of image captioning with reading comprehension. The dataset collection involved annotators describing images in one sentence that required reading the text present in the image. The dataset contains 145,329 captions for 28,408 images, with a focus on integrating OCR tokens into image descriptions. Annotators were tasked with generating captions that relate text to visual context, requiring spatial, semantic, and visual reasoning. This dataset presents new technical challenges for existing image captioning models, as it requires understanding and incorporating text from images into descriptive sentences. The TextCaps dataset can be used to address the challenge of teaching image captioning models how to ”read” and process information from text within images, potentially benefiting visually impaired individuals and advancing AI research in image understanding tasks.
  • NoCaps (Agrawal et al., 2019). The Nocaps dataset enhances image captioning with 166,100 captions for 15,100 images from the Open Images Dataset V4, focusing on objects not in COCO Captions, including nearly 400 novel objects. Images, primarily from Flickr, are categorized into in-domain, near-domain, and out-of-domain, based on their object classes’ relation to COCO, to test model generalization across familiar and novel concepts. Developed with AMT captions, the dataset introduces a ”priming” method for more accurate descriptions by hinting at object categories during annotation. This large-scale, diverse dataset aims to bridge the gap between existing captioning datasets and the vast array of real-world visual concepts, providing a challenging benchmark for novel object captioning. Nocaps dataset pushes image captioning models towards broader visual comprehension and linguistic description capabilities, aiming for applicability in more generalized, real-world scenarios.
  • VizWizCap (Gurari et al., 2020). The VizWizCap dataset, also known as VizWiz-Captions, comprises over 39,000 images taken by people who are blind, each annotated with five captions to facilitate research in image captioning algorithms that better serve blind users. Constructed using crowdsourced captions from the Amazon Mechanical Turk platform, this dataset includes metadata on text presence and image quality, reflecting real-world variability in the images. The primary purpose of the VizWizCap dataset is to advance the development of image captioning technology by providing a resource that addresses the specific needs of blind individuals, challenging existing computer vision models with images that significantly differ from those in controlled environments typically found in other datasets. This initiative aims to bridge technological gaps and enhance the practical utility of assistive technologies for visually impaired users.
  • AI Challenger Captions (Wu et al., 2017). The AI Challenger dataset was created to address the limitations of existing image caption datasets labeled in English. It is the largest image captioning dataset with Chinese annotations, consisting of three sub-datasets: human keypoint detection, large-scale attribute dataset, and image Chinese captioning. The dataset aims to bridge the semantic gap between low-level images and high-level concepts by providing rich annotations such as class labels, keypoint coordinates, bounding boxes, attributes, and captions in Chinese. This dataset not only serves as an effective benchmark for evaluating and improving computational methods in image understanding but also offers a valuable resource for researchers and practitioners working on tasks related to image understanding.
  • VIST (Huang et al., 2016). The VIST dataset pioneers sequential vision-to-language learning by offering 81,743 unique photos across 20,211 sequences, aimed at visual storytelling. Derived from Flickr, it’s structured into three tiers: isolated image descriptions (DII), sequential image descriptions (DIS), and image sequence stories (SIS), enriched by crowdsourced annotations from Amazon Mechanical Turk. This tiered approach enables the study of narrative and temporal context effects on visual comprehension. The dataset underwent careful preprocessing, including tokenization and anonymization, to ensure consistency and privacy. An automatic evaluation metric, grounded in METEOR scores, benchmarks storytelling performance, aligning closely with human judgment. ”VIST” significantly advances AI’s capability to not just describe but narrate visual sequences, enhancing machines’ understanding and generation of human-like stories from visual data, marking a leap towards more nuanced and context-aware AI storytelling abilities.
Interleaved Image-Text Document Datasets
Interleaved image-text document datasets consist of a text document interspersed with several illustrative images, wherein the correlation between the images and the accompanying text tends to be relatively lower.
  • M3W (Alayrac et al., 2022). The M3W dataset (MultiModal MassiveWeb dataset), derived from about 43 million webpages, is crafted for training Flamingo on interleaved text and image data. It features data from the Document Object Model (DOM) of webpages, with ¡image¿ tags inserted in plain text at their original positions, followed by a special ¡EOC¿ token for delineation. To ensure manageability and diversity, from each document, a sequence of 256 tokens and up to the first 5 images are selected, discarding the rest. The dataset emphasizes clean, non-English, and explicit content-free data, using custom scrapers for content extraction that maintains the hierarchical structure of HTML. It undergoes rigorous preprocessing, including text filtering to eliminate low-quality documents and image filters to remove unsuitable visuals. M3W is instrumental for developing few-shot learning in models, enabling them to perform complex image and video tasks with minimal examples, thereby serving as a key resource for advancing multimodal machine learning. This dataset is not open-source, luckily, the Open Flamingo provide an alternative open-source dataset called MMC4.
  • MMC4 (Zhu et al., 2024). The Multimodal C4 (MMC4) dataset is constructed as a significant enhancement of the well-known text-only C4 corpus, incorporating images alongside text to foster advancements in in-context vision and language models. The construction of the dataset involves several meticulous steps, leveraging the public web pages contained in the cleaned English C4 corpus as the primary source. The process includes standard preprocessing methods like deduplication, NSFW (not safe for work) image removal, ad filtering, and more. Images are placed into text sequences using a linear assignment algorithm based on CLIP features, a method that showed superiority over alternatives in aligning images to relevant textual content. After these filtration processes, the final MMC4 dataset comprises over 101.2 million documents, featuring 571 million images interleaved within 43 billion English tokens. The Multimodal C4 (mmc4) dataset is designed to advance the capabilities of in-context vision and language models, enabling them to understand and process interleaved sequences of images and text.
  • OBELICS (Laurençon et al., 2024). The OBELICS dataset is a meticulously constructed, open web-scale filtered dataset comprising 141 million web pages, 353 million images, and 115 billion text tokens derived from the Common Crawl dumps. Designed to advance the training of large multimodal models, it addresses the scarcity of publicly available datasets by providing a rich collection of interleaved image-text documents. The construction process encompasses several key steps: collecting and simplifying HTML files, extracting and filtering multimodal web documents, and implementing responsible filtering and deduplication measures to ensure quality and minimize inappropriate content. By preserving the natural structure of web documents and emphasizing the contextual interplay between text and images, OBELICS aims to support the development of next-generation models capable of sophisticated understanding and generation of multimodal content, thereby filling a critical gap in resources available for multimodal research and model training.
Visual Question Answer Image-Caption Datasets.
These datasets are usually driven by asking question on the image and giving out the answer according to the character of the image. Usually each image in this kind of datasets contains one question and one short answer as its caption.
  • VQAv2.0 (Goyal et al., 2019). VQAv2.0 is a new dataset aiming to help with the problem in understanding pictures and answering questions about them, known as Visual question answering(VQA) (Antol et al., 2015) problem. The VQA2.0 dataset contains approximately 1.1 million image-question pairs, almost twice the size of the original VQA dataset. It includes around 13 million answers associated with roughly 200,000 images from the COCO dataset. The dataset is divided into train, validation, and test splits, with complementary images and answers collected for all splits. The VQA2.0 dataset was constructed by adding ”complementary” images to the existing VQA dataset. For each question-answer pair associated with an original image, a similar yet distinct image was selected, where the same question would have a different answer. This approach resulted in a dataset where each question is linked with two images, leading to two different answers, thereby doubling the number of image-question pairs. This method aimed to balance the dataset and reduce the inherent biases towards certain answers. The complementary images were identified through a data collection protocol involving human annotators who selected images from a set of nearest neighbors (based on deep features) that were visually similar to the original image but would lead to a different answer to the same question.
  • Visual-7W (Zhu et al., 2016). Visual-7W enriches COCO images with 327,939 QA pairs, linked via object-level annotations to 47,300 images, covering what, where, when, who, why, how, and which questions. Questions and answers were generated through Amazon Mechanical Turk, where workers were asked to create natural language expressions for objects and another set to identify objects based on these expressions. This dual-step process ensures high-quality, relevant textual descriptions, directly tied to image regions. Visual-7W is a leap towards models that understand both the visual and textual components of an image, offering a robust platform for advanced AI research.
  • ST-VQA (Biten et al., 2019). The ST-VQA dataset creation process involved collecting images from various datasets containing scene text and other computer vision data sources. The ST-VQA dataset, incorporating 23,038 images from diverse sources like ICDAR, ImageNet, VizWiz, Visual Genome, and COCO-Text, focuses on integrating textual information within images for answering questions. Questions and answers were crowd-sourced through Amazon Mechanical Turk, ensuring unambiguous answers based on text in the images. The dataset comprises 23,038 images with 31,791 question-answer pairs for training and testing. ST-VQA emphasizes leveraging high-level semantic information from text in images for visual question answering tasks. It addresses the challenge of reasoning and generating answers based on scene text, providing a unique dataset for exploring the integration of textual cues in the VQA process.
  • Shikra-RD (Chen et al., 2023j). Shikra-RD is a dataset generated using GPT-4 to annotate Flickr30K Entities images with relational descriptions (RD), focusing on objects identified within bounding boxes. Despite GPT-4’s inability to view images, it interprets them through detailed descriptions and bounding box formats to create Q&A pairs, emphasizing questions that leverage known information for answers. This process resulted in 5,922 QA pairs, including coordinate data. Shikra-RD, leveraging advanced language models for image understanding, aims to enrich RD datasets by incorporating spatial annotations and natural language questions and answers.
  • OCR-VQA (Mishra et al., 2019). OCR-VQA consists of 207,572 book cover images, 1 million QA pairs, designed for reading text in images through OCR. The dataset emphasizes visual question answering (VQA) by incorporating optical character recognition (OCR) to interpret text on book covers for answering questions. QA pairs are derived from metadata and human paraphrasing of template questions, addressing titles, authors, genres, etc. The dataset, with its unique focus on text reading within VQA, poses challenges in text detection and recognition, aiming to advance both document analysis and VQA research fields.
  • DocVQA (Mathew et al., 2021). The DOC-VQA dataset, encompassing over 50,000 questions on more than 12,000 document images, was annotated through a meticulous three-stage process using a web-based tool and the efforts of remote workers. Workers initially generated question-answer pairs directly from text present in the images, aiming for extractive QA akin to NLP and VQA tasks. Subsequent stages involved verification and refinement of these QAs, including assigning question types and ensuring the precision of answers. This process underscores DOC-VQA’s commitment to high-quality, relevant, and diverse annotations for advancing document-based VQA research.
  • A-OKVQA (Marino et al., 2019). The A-OKVQA dataset was created by generating a diverse set of about 25K questions that require a broad base of commonsense and world knowledge to answer. Unlike existing knowledge-based VQA datasets, the questions in A-OKVQA cannot be answered simply by querying a knowledge base; instead, they demand some form of commonsense reasoning about the scene depicted in the image. This dataset aims to push the boundaries of Visual Question Answering by requiring models to recognize the image, understand the question, recall relevant knowledge, and use reasoning to arrive at an answer. It serves as a challenging testbed for AI models to reason over visual and natural language inputs, addressing the need for more complex reasoning systems in VQA tasks.
  • TextVQA (Singh et al., 2019). The Text-VQA dataset comprises 28,408 images with 45,336 questions requiring textual information within images to answer. Sourced from Open Images, the dataset spans categories likely to include text, aiming to enhance VQA models’ text-reading capabilities. Questions and answers were collaboratively generated using a web-based annotation tool by remote workers in a three-stage process, ensuring the questions necessitate text interpretation for answers. This dataset pushes the boundaries of VQA by demanding models to not only perceive images but also read and comprehend text within them, marking a significant step towards models assisting in real-world tasks, particularly aiding visually impaired users.
  • GQA (Hudson and Manning, 2019). The GQA dataset is designed to address deficiencies in previous visual question answering (VQA) datasets by fostering advanced visual reasoning and reducing biases. It utilizes over 22 million questions generated from 113K real-world images annotated with detailed scene graphs, detailing objects, attributes, and relationships. Each question is crafted through a robust engine combining linguistic grammar with scene graph data, and is accompanied by a functional program that semantically represents the reasoning steps needed for answering, ensuring complexity and diversity. GQA also introduces new metrics for evaluating consistency, validity, plausibility, and grounding of responses, moving beyond simple accuracy to offer a more comprehensive assessment of AI models’ visual and compositional understanding. This dataset aims to promote the development of AI capabilities in deep semantic understanding and unbiased reasoning in visual contexts.
Domain-Specific Image-Caption Datasets.
Domain-specific image-caption datasets are collections of images paired with captions that are tailored to a particular field or industry. It can enhance the performance of LLMs in specific domains.
  • Screen2Words (Wang et al., 2021). The Screen2Words dataset is designed to advance the automatic summarization of mobile user interfaces (UIs), encompassing 22,417 unique Android UI screens derived from the RICO dataset. This dataset is annotated by professional labelers, yielding 112,085 language summaries aimed at capturing the core functionalities of each screen for applications in screen readers, conversational agents, and other language-based interfaces. It supports multi-modal learning, leveraging text, images, and UI structure to enhance semantic understanding and interaction. The primary goal is to foster research in HCI and machine learning to improve accessibility and usability, making it easier for systems to provide concise, useful descriptions of mobile UIs automatically.
  • WidgetCap (Li et al., 2020). The WidgetCap dataset aims at enhancing mobile UI accessibility by generating descriptions for UI elements. It comprises 162,859 human-annotated captions for 61,285 UI elements across 21,750 screens from 6,470 apps, sourced from the RICO dataset and manually annotated via Amazon Mechanical Turk. This initiative addresses the prevalent issue of missing UI captions, crucial for accessibility and voice interactions. The dataset’s multimodal nature, combining image and structural UI data, enables comprehensive model training for accurate caption generation. WidgetCap’s creation process involved rigorous preprocessing to ensure data quality, focusing on visible and clickable elements essential for interaction. By providing a vast resource for widget captioning, WidgetCap facilitates research and development towards more accessible and interactive mobile applications, representing a significant stride in bridging accessibility gaps in mobile UI design.
A.1.2. Video-Caption Datasets
In this subsubsection, we provide a comprehensive summary of video caption datasets to help people find the most suitable datasets for MLLMs training.
  • MSR-VTT (Xu et al., 2016). Microsoft Research Video-to-Text (MSR-VTT) is a large-scale dataset offering 10,000 video clips totaling 41.2 hours, accompanied by 200K clip-sentence pairs. Derived from 257 queries across 20 categories, the dataset ensures diversity and representativeness for training large parameter models. MSR-VTT is meticulously curated, involving the collection of 30,404 videos from top search results, clip selection based on shot segmentation, and sentence annotation by Amazon Mechanical Turk workers. Noteworthy adopters of MSR-VTT include Git (Wang et al., 2022a), Vatt (Akbari et al., 2021), Videoclip (Xu et al., 2021b), Lavis (Li et al., 2022a), and TVLT (Tang et al., 2022), showcasing its significance in language-video research.
  • WebVid 2M, WebVid 10M (Bain et al., 2021). The WebVid dataset originally encompasses 2.5 million video clips collected from various web sources, each paired with a caption, and subsequently expands to encompass 10 million pairs of video and text totaling 13k hours. Scraped from web sources similar to the image-text dataset Google Conceptual Captions (Sharma et al., 2018) (CC3M), WebVid offers diverse video content with text descriptions, adhering to ethical processing standards for copyright compliance. The textual descriptions in WebVid are typically manually crafted, allowing for better alignment with the content of the video images due to the well-structured sentences, fewer grammatical errors, and richer vocabulary. Adopted by models like Stable Video Diffusion (Blattmann et al., 2023), Next-gpt (Wu et al., 2023a) and Videochat (Li et al., 2023c), WebVid serves as a valuable resource for text-to-video generation task. However, because of the copyright, WebVid is no longer available.
  • VTP (Alayrac et al., 2022). VTP (Video & Text Pairs) consists of 27 million short videos paired with descriptive sentences, collected by crawling less than ten websites known for high-quality and detailed video descriptions. Each instance within the dataset consists of a sequence of UTF-8 bytes encoding the textual content and a compressed video. VTP is processed in a manner including adding image labels and special tokens. Additionally, automated text formatting is applied to eliminate irrelevant information, such as dates and locations, from the captions, thus refining the dataset for training purposes. Notably, deduplication is omitted due to the unique sourcing strategy of this dataset, ensuring minimal overlap with common video evaluation datasets. VTP is leveraged by models such as Flamingo (Alayrac et al., 2022) for few-shot Learning.
  • Panda70M (Chen et al., 2024c). The Panda-70M dataset was created through an automatic captioning pipeline that leverages multimodal information, such as video descriptions, subtitles, and individual static video frames. This dataset contains high-resolution and semantically coherent video samples, aiming to provide precise descriptions of the main objects and actions in the videos. The dataset addresses the limitations of existing datasets by offering a large-scale collection of videos with annotations generated by multiple cross-modality vision-language models. It can be utilized to enhance various downstream tasks, including video captioning, video and text retrieval, and text-to-video generation.
A.1.3. Audio-Caption Datasets
In this subsubsection, we provide a comprehensive summary of audio caption datasets to help people find the most suitable datasets for MLLMs training.
  • AISHELL-2 (Du et al., 2018). AISHELL-2 is a Mandarin Chinese Speech Corpus containing 1000 hours of clean speech data from 1991 Chinese speakers across diverse regions. This dataset covers 12 domains including keywords, voice commands, smart home, autonomous driving and industrial production, etc. The data was recorded in a quiet indoor environment using the iOS-system. Through professional speech annotation and strict quality inspection, the manual transcription accuracy rate of this dataset is above 96%.
  • WavCaps (Mei et al., 2023). WavCaps is a large-scale English audio captioning dataset that offers a vast collection of around 400k audio clips, each paired with corresponding captions. The dataset is sourced from three prominent websites, including FreeSound (Font et al., 2013), BBC Sound Effects  [^1] https://sound-effects.bbcrewind.co.uk/, and SoundBible  [^2] https://soundbible.com/, along with a sound event detection dataset AudioSet Strongly-labelled Subset (Hershey et al., 2021). To ensure high data quality, the dataset undergoes a three-stage processing pipeline, which can filter and convert the raw descriptions obtained from the aforementioned sources into accurate and informative captions using ChatGPT  [^3] https://pypi.org/project/langdetect/.
  • VSDial-CN (Chen et al., 2023b). VSDial-CN is a multi-modal Automatic Speech Recognition (ASR) dataset constructed from VisDial (Das et al., 2017). It comprises around 120,000 images, each accompanied by a caption and 10 rounds of dialogue consisting of questions and answers. To create VSDial-CN, the text from the VisDial dataset was translated into Chinese, and the questions were then synthesized into speech using the FastSpeech2 (Ren et al., 2020) model. The dataset offers approximately 1,200k training samples for multimodal ASR. After removing duplicate questions, there are approximately 370k unique speech utterances available for training.
A.1.4. 3D-Caption Datasets.
  • ScanNet (Dai et al., 2017). With a substantial scale comprising 2.5 million views across 1513 scenes, ScanNet offers RGB-D video data annotated with 3D camera poses, surface reconstructions, and semantic segmentations. ScanNet, which is a snapshot of available data from roughly one month of data acquisition by 20 users at locations in several countries, incorporates annotations from over 500 crowd workers via the Mechanical Turk platform. Its creation involved a meticulous process encompassing data collection, RGB-D scanning and calibration, data upload and processing, semantic annotation, and dataset organization. Notably, a data collection framework was designed to enable untrained users to capture semantically labeled indoor scenes using commercial hardware. The captured RGB-D videos underwent rigorous processing stages, including calibration to align depth data with color data, surface reconstruction, camera attitude estimation, and quality verification. Semantic annotation, performed through crowdsourcing, involved instance-level object class labeling and CAD model alignment tasks. ScanNet has been included in the pytorch library Kaolin and adopted by models such as ENet, FCAF3D, PointNet++.
  • S3DIS (Armeni et al., 2017). S3DIS provides a comprehensive collection of indoor scene point clouds. S3DIS encompasses six large-scale indoor areas with a total of 271 rooms, with each point in the scene point cloud annotated with one of the 13 semantic categories. Point clouds are automatically generated using the Matterport scanner without any human intervention, sets of RGB-D images registered on color point clouds was collected to enrich the dataset. Point cloud data was further processed to parse the building’s 3D point clouds into semantically meaningful spaces like rooms, which were then subdivided into structural and architectural elements like walls and columns. This involved using a 3D sliding window-based approach and machine learning techniques for semantic element detection and labeling. The application of graph models enhanced the context consistency between elements, facilitating effective extraction and understanding of semantic information in complex indoor spatial point clouds. S3DIS has been adopted by models including FCAF3D, PointNet, and PointCNN.
  • Structured3D (Zheng et al., 2020b). Structured3D is a comprehensive collection of 3,500 home designs with 21,835 rooms, derived from a large database of house designs meticulously created by professional designers. It includes detailed information on interior design elements in industry-standard format, detailing object geometry, materials, textures, and functional information like room assignments. An automated program was developed to extract geometric primitives related to room structure, including ceilings, floors, walls, and openings, from a large database of house designs hand-crafted by professional designers. Precise measurements and related information are used to construct detailed models, including all planes, lines, connections, and their relationships. For the creation of photorealistic 2D renderings, the dataset utilizes a specialized rendering engine that employs ray tracing method to approximate the real global lighting effect for RGB rendering. The rendering process uses over a million CAD furniture models from leading manufacturers. Furthermore, default lighting Settings are provided and images with different visual effects are generated by adjusting furniture configurations and lighting Settings. The dataset has been adopted by models such as Structured3D, Ponder V2, and Swin3D.
A.2. Commonly-Used Instruction Tuning Datasets
Instruction tuning datasets are used to fine-tune models to better understand and follow specific instructions, improving their ability to generate accurate and relevant responses. These datasets help enhance the model’s performance in various tasks by providing it with clear and structured examples of desired outputs based on given inputs. We summarize the commonly used multimodal instruction tuning datasets here.
A.2.1. Image instruction tuning datasets
In this subsubsection, we provide a comprehensive summary of image instruction tuning datasets to help people find the most suitable datasets for visual instruction tuning.
  • LLAVA-Instruct-150K (Liu et al., 2024b, 2023c). The LLAVA Instruct 150k dataset is designed to enhance the training of multimodal models that interpret and respond to visual and textual instructions, featuring 150,000 instruction-response pairs derived from image-text pairs from public datasets like COCO and LAION. Utilizing GPT-4, these pairs are transformed into instruction-based queries that require understanding and reasoning about visual content, aligning responses with both direct descriptions and deeper inferential reasoning.
  • MiniGPT-4-IT (Zhu et al., 2023). The MiniGPT-4-IT dataset is composed of 3,500 detailed image description pairs. These pairs were derived from a blend of sources including the Conceptual Captions, SBU, and LAION datasets. This collection was designed to train the MiniGPT-4 model to improve its language generation in response to visual inputs. The primary objective of the dataset is to facilitate the alignment of complex visual features with advanced language model capabilities, thereby enhancing the model’s ability to generate natural and contextually appropriate descriptions for multimodal tasks. Essentially, MiniGPT-4-IT serves as a critical resource for bridging the gap between visual understanding and linguistic expression, aiming to boost the performance of AI models in interpreting and describing visual content.
  • MiniGPT-v2-IT (Chen et al., 2023m). The MiniGPTv2-IT dataset, crafted for instruction tuning, integrates a variety of data types such as weakly-labeled image-text pairs, grounded captions, and fine-grained visual question answering (VQA) data across three training stages. Initial stages build broad vision-language capabilities using diverse sources like GRIT-20M, LAION, CC3M, and SBU. Later stages focus on fine-tuning with high-quality, task-specific datasets including COCO caption and RefCOCO. This structured approach enhances the model’s efficiency in processing complex multimodal instructions and boosts its performance across various vision-language tasks. MiniGPTv2’s training leverages millions of data points to develop robust multimodal interaction capabilities, making it a powerful tool in vision-language applications.
  • Shikra-IT (Chen et al., 2023j). The Shikra-IT dataset for instruction tuning includes two primary types of data: reorganized public datasets and newly generated high-quality RD data from Flickr30K Entities. It incorporates diverse sources such as VQA, Image Captioning datasets, and those with positional annotations like RefCOCO and Visual Genome. Additionally, 5,922 QA pairs were created using GPT-4, explicitly designed to include coordinate information in both questions and answers. This approach enhances the Shikra model’s referential dialogue capabilities, enabling precise multimodal interactions. These interactions are crucial for applications that require accurate spatial references, such as augmented reality interfaces and visual customer support systems, where identifying and discussing specific areas within images are essential. The dataset effectively combines existing resources with AI-generated enhancements to train the model in handling complex, context-rich dialogues.
  • CogVLM-SFT-311K (Wang et al., 2024c). The CogVLM-IT dataset integrates data from visual question-answering sources like VQAv2, OKVQA, TextVQA, and OCRVQA, along with dialogue datasets such as LLaVA-Instruct and LRV-Instruction, into a unified instruction-supervised fine-tuning framework. This collection trains the CogVLM model to enhance its multimodal interaction capabilities, enabling it to handle concise responses and detailed reasoning effectively. The dataset is crucial for ensuring the model’s robust performance across diverse visual and language tasks, significantly aiding in real-world applications that require precise and context-aware multimodal responses. This strategic compilation of varied data sources ensures comprehensive training and adaptability of the CogVLM model.
  • VILA-IT (Lin et al., 2023a). The VILA-IT dataset is designed for instruction tuning with visual language models, blending 1 million text-only instruction data from FLAN with visual language datasets. This integration aims to address the degradation in text-only task performance while enhancing visual language capabilities. The dataset primarily includes visual question answering and caption style data, collected and generated to support joint supervised fine-tuning. This method significantly improves both text-based and visual task accuracies, demonstrating the effectiveness of merging text and visual data for training advanced visual language models. Overall, VILA-IT plays a crucial role in refining the performance of visual language models across diverse tasks.
  • ShareGPT4V (Chen et al., 2023e). The ShareGPT4V dataset comprises 1.2 million detailed captions, initially created from 100K high-quality examples generated by GPT4-Vision. These captions were expanded using a specialized model to enhance multimodal model training, focusing on Supervised Fine-Tuning (SFT) and pre-training stages. The dataset emphasizes comprehensive coverage of topics such as world knowledge, object properties, and spatial relationships to improve modality alignment between visual and textual data. Utilized to develop the ShareGPT4V-7B model, it has shown superior performance across various benchmarks, demonstrating the benefits of detailed image-text pairs in multimodal interactions. ShareGPT4V exemplifies how rich, detailed captions can significantly boost the capabilities of large multimodal models.
  • GranD (Rasheed et al., 2023). The GranD dataset is a large-scale dataset containing 11 million images and 810 million regions with segmentation masks, created for training and evaluating models on visually grounded conversation tasks. This dataset is generated using an automated annotation pipeline that involves multi-level hierarchical structuring and state-of-the-art models to provide detailed object localization, attributes, and relationships. The dataset enables instruction-response pairs by using complex visual scenes structured as queries to generate contextually rich, detailed responses, intertwined with visual prompts that ensure accurate grounding in visual elements. GranD supports the Grounded Conversation Generation (GCG) task, producing responses integrating descriptions with segmentation masks. This focus on dense annotations and complex visual reasoning sets a new benchmark in multimodal datasets, advancing visually grounded language models.
  • ComVint (Du et al., 2023a). The ComVint dataset consists of 32K synthetic visual reasoning instructions designed to enhance the performance of multimodal large language models (MLLMs). Developed through a systematic multi-stage process of synthesis, complication, and reformulation, the dataset utilizes complex visual reasoning tasks to create instructions. The generation process starts with GPT-4 synthesizing an initial instruction from an image’s annotations. Then, through an iterative ”complicate-then-verify” paradigm, the complexity of these instructions is gradually increased, ensuring both the complexity and quality are maintained. The dataset primarily aims to boost MLLMs’ capabilities in complex visual reasoning, demonstrating substantial improvements in model performance on comprehensive benchmarks. Overall, ComVint significantly contributes to advancing AI’s ability to interpret and reason about complex visual inputs within a multimodal context.
  • SparklesDialogue (Huang et al., 2023). The SparklesDialogue dataset is crafted using a GPT-4 assisted process, where GPT-4 simulates realistic dialogues between a user and an assistant, involving multiple images. It is part of the Sparkles project, aimed at enhancing multimodal instruction-following models. Images and textual data are sourced from various image-caption pairs. During its construction, GPT-4 uses detailed image descriptions to generate dialogues, ensuring rich textual interaction while no actual images are sent to GPT-4. This process includes Dialogue Demonstrations, which serve as learning examples for creating coherent and contextually relevant dialogues, and Candidate Image Descriptions, which provide a diverse pool of images for generating discussions. This setup aids in training models like SparklesChat, enhancing their ability to handle intricate multimodal interactions. SparklesDialogue thus plays a crucial role in advancing complex managing, context-aware multimodal conversations and improving multimodal understanding capabilities.
  • TextBind (Li et al., 2023e). The TEXTBIND dataset from GPT-4 consists of 25,629 multi-turn conversations, integrating images and text to train models on multimodal instruction-response tasks. This dataset leverages a three-step pipeline involving topic-aware image sampling, LLM-generated dialogues, and rigorous post-processing. It sources images and text from publicly available image-caption pairs, designed to enhance language models’ ability to handle complex, interleaved image-text content. TEXTBIND’s primary feature is its focus on generating practical, context-rich conversations that reflect diverse real-world scenarios, supporting multimodal outputs. This approach significantly advances the capability of models to navigate open-ended, multimodal interactions, providing a robust platform for developing advanced generative models.
  • PVIT (Chen et al., 2023h). The PVIT dataset consists of region-level instruction data, including 234k image-text pairs. This dataset is generated using three main strategies: 1) Dataset Conversion, where existing Visual Question Answering (VQA) datasets like GQA and VCR are reformatted into region-level instructions using specific templates, producing 146k entries. 2) Task-Specific Instruction Data Generation, which employs ChatGPT to create instructions for predefined multimodal tasks (e.g., object recognition), resulting in 20k single-turn and 66k multi-turn data entries, using detailed image annotations from datasets like MS COCO and Visual Genome. 3) General Instruction Data Generation improves diversity and quality by generating 22k entries with detailed descriptions and visual grounding annotations. These strategies collaboratively enhance the dataset’s utility in training AI models for complex image-based instruction following, improving AI’s understanding and generation capabilities in diverse scenarios.
  • DataEngine-InstData (Zhao et al., 2023b). The DataEngine-InstData dataset is designed to enhance the instruction-following capabilities of AI models using a semi-automated process known as Interactive Prompt Optimization (IPO). This dataset contains QA pairs that are specifically generated using GPT-4, optimized through human collaboration to correct and refine prompts based on identified failures in preliminary tests. These corrections are focused on adhering to the desired question types and accurate use of bounding box annotations to reduce illusions. The process of creating these QA pairs includes manually designing initial prompts, testing, and refining based on failure analysis. The dataset serves to train AI in handling complex queries effectively, particularly improving performance on spatial and imaginative questions. The final dataset aims to provide robust training material that supports advanced AI learning protocols without incorporating inaccurate or inappropriate data.
A.2.2. Video Instruction Tuning Datasets
In this subsubsection, we provide a comprehensive summary of video instruction tuning datasets to help people find the most suitable datasets for video instruction tuning.
  • Video-Chat-Instruction (Li et al., 2023c). This multimodal instruction dataset was created using WebVid-10M as the foundation. This dataset includes 7,000 detailed video descriptions and 4,000 video-based conversations. The detailed descriptions and the multi-turn conversation generations were created by ChatGPT, which utilized the text from VideoChat-Text and several prompts focusing on the spatiotemporal aspects of the videos. To enhance the diversity of the video instruction data, video conversations were introduced, enriching the dataset with temporal and causal dynamics. This enriched dataset was employed to train models such as Videochat (Li et al., 2023c) and Videollama (Zhang et al., 2023).
  • Video-Chat-Instruction (Li et al., 2023h). The MVBench Instruction Tuning dataset comprises 1.9 million samples drawn from 34 diverse datasets across 6 categories. It was constructed by incorporating both image and video data in the instruction set to improve training for Multi-modal Large Language Models (MLLMs). This dataset can be used to address the limited diversity in instruction-tuning data and to evaluate the general capability of MLLMs for open-world temporal understanding, covering tasks from percwption to cognition. This enriched dataset was employed to train models such as Videochat2 (Li et al., 2023h) and InternVideo2 (Wang et al., 2024b).
A.2.3. Audio Instruction Tuning Datasets
  • Qwen-Instruction (Chu et al., 2023). The Qwen-Audio Instruction Tuning dataset consists of 20k data points and is constructed by combining audio-centric instruction data with pure text instruction data. This dataset is designed to facilitate versatile input from both audio and text modalities within multi-turn dialogues. By leveraging this dataset during training, the model can seamlessly handle diverse forms of input, enabling effective interaction following human instructions.
  • Audio-Dialogues Instruction (Goel et al., 2024). The Audio-Dialogues Instruction Tuning dataset comprises 163.8k samples and is designed to address the limitations of existing datasets by focusing on multi-turn dialogues for general sounds and music understanding. The dataset was constructed using a prompting-based approach with caption annotations from the AudioSet strongly labeled dataset and the MusicCaps dataset, guided by the GPT-4 model. This dataset aims to facilitate the training and evaluation of existing audio-augmented large language models on complex interactions regarding audio content, providing a valuable resource for tasks such as audio captioning, sound event classification, and music recommendation systems.
  • HuBERT Instruction (Hsu et al., 2021). The HuBERT tokens Instruction Tuning dataset in the article consists of a substantial amount of data that is used for fine-tuning the model. This dataset is constructed by extracting tokens from audio instructions, providing a valuable resource for training the model to better understand and process spoken instructions. By leveraging this dataset, HuBERT can improve its performance in tasks related to processing and interpreting audio instructions, ultimately enhancing its ability to handle speech-based tasks effectively.
  • LLaSM Instruction (Shu et al., 2023a). The LLaSM Instruction Tuning dataset, named LLaSM-Audio-Instructions, consists of a total of 199k conversation data and 508k samples. This dataset was constructed by generating audio data from language-only datasets and filtering out irrelevant conversations and chatbot responses. By utilizing Microsoft Azure text-to-speech API to convert human responses into speech data, the dataset provides complex instructions for fine-tuning the model. This dataset aims to address the scarcity of cross-modal speech-and-language instruction data and enhance the model’s ability to follow multi-modal instructions effectively.
  • BuboGPT-IT (Zhao et al., 2023a). The BuboGPT-IT dataset consists of three categories for instruction tuning. The image-text portion merges 3,439 pairs from MiniGPT-4 with 158K from LLaVA, focusing on detailed descriptions and complex reasoning. The audio-text dataset, termed Clotho-Detail, contains 3,938 extended audio captions, enhancing descriptive richness. The audio-image-text set, sourced from VGGSS, includes 5,158 triple-modality pairs, aligning sounds with their visual sources and incorporating negative pairs to improve semantic reasoning. These datasets were developed by augmenting existing data with AI-generated content, aiming to support complex multimodal interactions. They equip BuboGPT with the capability to handle sophisticated instruction-based tasks across various modalities, boosting its performance in intricate multimodal environments.
A.2.4. 3D Instruction Tuning Dataset
  • X-InstructBLIP (Panagopoulou et al., 2023). The X-InstructBLIP’s IT dataset is designed to support instruction tuning across multiple modalities including images, audio, video, and 3D models. It includes instruction-response pairs with 24K QA samples for audio and 250K QA samples for 3D. The data is collected by leveraging open-source large language models to generate question-answer pairs from existing captioning datasets, such as Cap3D for 3D models and AudioCaps for audio data. The dataset incorporates individual modality training to form cross-modal reasoning capabilities without extensive modality-specific pre-training. The primary feature of this dataset is its ability to facilitate the fine-tuning of instruction-modality alignment in a scalable and efficient manner, showcasing strong cross-modal reasoning by discerning between different modalities. In summary, the X-InstructBLIP’s IT dataset exemplifies a comprehensive and scalable approach to instruction tuning, effectively handling diverse modalities and enhancing cross-modal reasoning tasks.
  • LAMM (Yin et al., 2024). The LAMM dataset facilitates instruction-response pair training, containing 186,098 image-text pairs and 10,262 3D-text pairs from diverse publicly available datasets. Using the GPT-API, instructions and responses are generated through self-instruction methods based on the original dataset labels. This process involves extensive pre-processing, including the enrichment of visual information and conversion of vision task annotations into more comprehensible formats for MLLMs. LAMM uniquely emphasizes fine-grained information, supports commonsense knowledge through a hierarchical knowledge system, and enhances the understanding of vision-related instructions. It represents a significant stride in bridging modalities for AI training, leveraging detailed instructions to improve multimodal language model responsiveness and adaptability.
A.3. Commonly Used MLLM Evaluation Benchmarks
Given the abundance of research papers on image multimodality, we will separately discuss the evaluation datasets of image multimodality and the evaluation of other forms of multimodality.
A.3.1. Image Data Multimodal Evaluation
In this subsubsection, we summarize commonly used evaluation datasets that are utilized to assess the performance of image-text multimodal large language models (LLMs).
Image Caption and General VQA Evaluation Datasets.
Image captioning and general visual question answering (VQA) datasets serve as benchmarks for evaluating a model’s comprehensive understanding of visual content. These datasets can also be utilized to assess the few-shot learning capabilities of multimodal large language models (LLMs) in the context of vision-language tasks, providing insights into their ability to generalize and adapt to new scenarios with limited training examples.
  • MMMU (Yue et al., 2023). MMMU is a comprehensive multimodal benchmark that assesses models on their proficiency in college-level reasoning across multiple disciplines. Featuring 11,500 questions paired with complex images and text sourced from college exams, quizzes, and textbooks, it spans six core areas: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering. The dataset challenges models with college-level tasks that require deep domain knowledge and sophisticated reasoning, closely mimicking the cognitive challenges faced by professionals. MMMU’s detailed creation process ensures the accuracy and relevance of its content, testing models on perception, knowledge comprehension, and logical reasoning. This dataset is pivotal for advancing model capabilities in handling intricate, multimodal data in specialized academic fields.
  • MME (Fu et al., 2024). The MME dataset, structured into MMEP (Perception) and MMEC (Cognition) components, is designed for evaluating multimodal large language models. It comprises 14 sub-tasks across both segments, focusing on object recognition, optical character recognition, and advanced reasoning tasks like commonsense reasoning and numerical calculations. The dataset features image-text pairs and instructions, all manually created to avoid data leakage by excluding public dataset dependencies. Images are sourced from real photographs and other collections, tailored to test both basic and advanced multimodal understanding. MME challenges models to respond to concise, human-like prompts, pushing their capabilities in visual comprehension and cognitive processing. This benchmark is a comprehensive tool for revealing and enhancing the abilities of multimodal models, emphasizing their performance in complex, real-world tasks that blend visual perception with cognitive operations.
  • Nocaps (Agrawal et al., 2019). The Nocaps dataset features images from the Open Images V4 (Kuznetsova et al., 2020) validation and test sets, carefully selecting images to diversify object representation and avoid bias towards common classes like person, car, or plant. This selection aims to address the gap in COCO Captions (Chen et al., 2015), where nearly 400 out of 500 specific classes (excluding overly broad or infrequent ones) are seldom mentioned. Captions for nocaps, including human-generated baselines, were gathered through a specially adapted COCO collection interface, emphasizing high-quality contributions by employing only US-based Amazon Mechanical Turk workers with over 5,000 tasks and a 95% approval rate. A total of 727 workers contributed, averaging 228 captions each, resulting in a comprehensive set of 166,100 nocaps captions.
  • GQA (Hudson and Manning, 2019). Current VQA benchmarks fail to provide accurate indication of visual understanding capacity. Not only are they severely biased, they also lack semantic compositionality, and do not provide any tools or measures to gain significant insight into models’ performance and behavior. The authors design a new dataset, GQA, to address these shortcomings, featuring compositional questions over real-world images. The dataset consists of 22M questions over 113K various day-to-day images. To construct the GQA dataset, a wide array of real-world images depicting everyday scenes was selected from sources like Visual Genome. For each chosen image, detailed scene graphs were generated, describing the objects present, their attributes, and the relationships among them, utilizing advanced computer vision techniques. Questions probing both the direct content of the images and requiring deeper reasoning about the scenes were then crafted based on these graphs. Each question was provided with a structured semantic representation—a functional program that outlines the logical steps necessary for answering it. Throughout the creation process, the data underwent rigorous cleaning and validation to ensure accuracy and consistency across images, scene graphs, questions, and their semantic frameworks. Many of the GQA questions involve multiple reasoning skills, spatial understanding and multi-step inference, thus are generally more challenging than previous visual question answering datasets used in the community.
  • DVQA (Kafle et al., 2018). The DVQA dataset is designed to evaluate the comprehension of bar charts through a question-answearing framework, aiming to enhance algorithms’ ability to extract and interpret numeric and semantic information from a variety of bar chart styles found in diverse sources like scientific publications and business reports. Comprising over 3 million image-question pairs, DVQA emphasizes not just basic visual recognition but also complex cognitive tasks such as pattern matching, attention, and reasoning. It challenges existing visual question answering models with unique demands, such as handling dynamically generated text and answers specific to each chart, thus pushing forward the capability of AI systems to interact with and understand structured visual data effectively.
  • VSR (Liu et al., 2023b). Visual Spatial Reasoning (VSR) is a dataset comprising over 10,000 natural text-image pairs with 66 types of spatial relations in English, such as under, in front of, and facing. The dataset aims to highlight the challenges in capturing relational information for current vision-and-language models. Research indicates a significant performance gap between humans (achieving above 95% accuracy) and state-of-the-art models (reaching around 70% accuracy) on spatial reasoning tasks. VSR emphasizes the difficulties that current models face in recognizing spatial relations concerning object orientations and notes their poor generalization on unseen concepts. Through comprehensive analysis on popular vision-language models, it underscores the importance of positional encodings, the lack of correlation between model performance and training examples, and the specific challenges posed by certain spatial relations. The findings from VSR provide a realistic evaluation of models’ performance in understanding spatial relations.
  • OKVQA (Marino et al., 2019). The OK-VQA dataset was created to address the limitations of existing VQA benchmarks by focusing on knowledge-based visual question answering. This dataset consists of over 14,000 questions that require external knowledge to answer, covering various categories such as science & technology, history, and sports. The questions in OK-VQA challenge models to not only understand the image and question but also to retrieve and incorporate relevant knowledge from external sources. By providing a diverse and difficult set of questions, the OK-VQA dataset enables researchers to evaluate the reasoning capabilities of VQA models in scenarios where answers cannot be obtained solely from the image. This dataset opens up new research opportunities in the domain of knowledge-based visual question answering.
  • Vizwiz (Gurari et al., 2018). The VizWiz dataset was created by collecting visual questions from blind individuals seeking answers to daily visual challenges. This dataset stands out from others as it includes images taken by blind photographers, presenting challenges like image blur and poor lighting. The questions in VizWiz are spoken, reflecting a more conversational style and potential audio recording imperfections. By addressing real-world interests of blind users, the dataset aims to empower them in overcoming visual-based obstacles. Modern algorithms face difficulties in answering questions from VizWiz due to its unique characteristics, highlighting the need for specialized solutions to assist blind individuals effectively.
  • POPE (Li et al., 2023b) The ”POPE” dataset focuses on evaluating object hallucination in vision-language models, using 6,136 binary questions to test if models hallucinate non-existent objects in images from MSCOCO and other sources. Each image is analyzed to establish a ground truth for yes-or-no queries about specific objects, ensuring each question corresponds accurately to image content. POPE assesses the reliability and accuracy of models in interpreting visual data by highlighting their propensity to generate hallucinatory content. This benchmark is crucial for testing object recognition accuracy and addressing significant challenges in model reliability for practical real-world applications.
Text-Oriented VQA Evaluation Datasets
While image captioning and general VQA datasets primarily focus on evaluating an algorithm’s understanding of visual content, some datasets in these categories also include questions that require reading and comprehending text present in the images, thus providing a test for the text recognition capabilities of the models.
  • TextVQA (Singh et al., 2019). The TextVQA dataset was created by selecting images containing text from the Open Images dataset through a three-stage crowdsourcing process, resulting in 28,408 images forming the basis of the dataset. Each image had 1-2 questions collected, with the first question requiring reading the text in the image to answer and the second question requiring reasoning about the text for a different answer. The dataset’s characteristics include a focus on text-related questions that require reading and reasoning about text in images, encouraging research on improving text detection and recognition in unconstrained environments, and enabling VQA models to read and reason about text in images. The dataset also introduces the LoRRA model architecture, which reads text in images, reasons about it based on the question, and predicts answers from a fixed vocabulary or the text found in the image. This dataset and model facilitate exploration and solutions for questions involving text in images in the VQA domain.
  • DocVQA (Mathew et al., 2021). The DocVQA dataset was created by collecting document images from the UCSF Industry Documents Library, spanning various industries and document types. The dataset consists of 50,000 questions on 12,000+ document images, with questions and answers collected through a web-based annotation tool. The dataset stands out for its focus on VQA on document images, requiring models to interpret text within the layout and structure of the documents. This dataset addresses the challenge of extracting information from document images through natural language questions, encouraging models to understand textual, graphical, and structural elements in the documents. It serves as a valuable resource for developing and evaluating models that can effectively answer questions based on the content of document images.
  • ChartQA (Masry et al., 2022). The ChartQA dataset was created by generating human-authored question-answer pairs based on real-world charts, aiming to address the limitations of existing datasets that rely on template-based questions and automatically generated charts. This dataset stands out for its diverse question types involving visual and logical reasoning, allowing for more complex analysis of chart data. By combining visual features and extracted data tables from chart images, the dataset enables the development of transformer-based QA models that achieve state-of-the-art results in handling questions that require both visual and logical reasoning.
  • AI2D (Kembhavi et al., 2016). The AI2D dataset was created by compiling over 5000 grade school science diagrams with more than 150,000 rich annotations, including ground truth syntactic parses and over 15,000 multiple choice questions. The diagrams were collected by scraping Google Image Search using seed terms from science textbooks. Each image was annotated using Amazon Mechanical Turk in a multi-phase process to ensure high agreement levels among annotators. The dataset’s rich annotations categorize constituents, relationships, and textual elements within the diagrams, making it a valuable resource for tasks such as syntactic parsing, semantic interpretation, and diagram question answering. The AI2D dataset’s unique characteristics and comprehensive annotations provide a solid foundation for advancing research in diagram interpretation and reasoning.
  • OCR-VQA (Mishra et al., 2019). The OCR-VQA dataset creation process involved obtaining book cover images with metadata, generating question-answer pairs, and annotating the dataset. This dataset stands out for its focus on visual question answering by reading text in images, filling a gap in traditional VQA literature. It addresses challenges such as robust layout analysis, OCR for fancy fonts, and identifying book categories based on covers. The dataset’s unique characteristics enable researchers to explore new avenues in document image analysis and VQA, offering a valuable resource for studying text-based visual question answering tasks.
  • ScienceQA (Lu et al., 2022) The ”Science QA” evaluation dataset, sourced from IXL Learning, features approximately 21,000 multimodal questions from science curricula, aimed at testing AI’s multi-hop reasoning abilities. Each question is accompanied by images and detailed explanations, also from IXL Learning, designed to mimic educational standards and reasoning processes. This dataset significantly preprocesses content to ensure relevance and educational alignment, evaluating AI’s capability to integrate multimodal inputs and external knowledge into coherent answers. ”Science QA” serves as a benchmark for AI models, emphasizing their interpretability and reliability in reasoning within educational contexts. It challenges AI systems to not only provide answers but also generate understandable explanations, pushing forward the development of AI technologies that can effectively support educational applications
  • MathV (Lu et al., 2023a) The ”MathV” dataset comprises 6,141 image-text pairs from 31 multimodal datasets, including 28 existing and three new ones (IQTest, FunctionQA, PaperQA). Images are sourced from diverse contexts such as natural scenes and scientific diagrams. Questions and answers, rooted in various mathematical reasoning types (algebraic, arithmetic, statistical), are generated from these visuals. The dataset is meticulously annotated with metadata to categorize each example by task type, grade level, and required reasoning skills. ”MathV” primarily serves as a comprehensive benchmark for evaluating the mathematical reasoning capabilities of AI models in visual contexts. It aims to assess and advance the development of AI models, particularly foundation models, by challenging their ability to interpret and solve visually-based mathematical problems, revealing both their strengths and limitations in real-world applications.
    MMVet (Yu et al., 2023) The ”MMVet” dataset is a benchmark for evaluating large multimodal models (LMMs) on complex tasks. It includes 200 images and 218 questions that span six core capabilities: recognition, OCR, knowledge, language generation, spatial awareness, and math. These components are integrated to assess the models’ ability to handle real-world scenarios effectively, such as interpreting jokes or solving visual arithmetic problems. Questions and images are sourced from diverse online platforms, including specialized domains like medical imaging. MMVet employs an LLM-based evaluator, notably using models like GPT-4, to score responses based on their conformity to varied answer styles and question types, adapting to the multifaceted nature of real-world applications. This benchmark challenges models to synthesize and apply multiple capabilities, providing a realistic assessment of their practical utility.
Refer Expression Comprehension Evaluation Datasets.
Referring Expression Comprehension (REC) evaluation datasets, such as RefCOCO, RefCOCO+, and RefCOCOg, are essential for assessing models’ ability to understand and ground natural language expressions in visual contexts. These datasets consist of images paired with referring expressions that describe specific objects or regions. The task involves accurately identifying the referred object based on the given expression. REC datasets challenge models to bridge the gap between vision and language, requiring the capture of fine-grained details, spatial relationships, and contextual information. They serve as benchmarks for evaluating the performance of models in understanding referring expressions, contributing to advancements in vision-language integration with various applications.
  • RefCOCO, RefCOCO+, RefCOCOg (Kazemzadeh et al., 2014; Mao et al., 2016). The RefCOCO dataset, built upon the MS COCO dataset, consists of 142,209 referring expressions corresponding to 50,000 objects in natural images. The referring expressions in RefCOCO are generated with a focus on the target objects, providing a dataset suitable for studying object-centric referring expression comprehension. The dataset includes segmentation masks for the target objects, enabling pixel-level analysis and evaluation. RefCOCO serves as a valuable resource for developing and benchmarking algorithms that aim to accurately ground referring expressions to specific objects within an image.
    RefCOCO+ is an extension of the RefCOCO dataset, also utilizing images from MS COCO. It contains 141,564 referring expressions, corresponding to 49,856 objects. The key difference between RefCOCO and RefCOCO+ lies in the way the referring expressions are generated. In RefCOCO+, the expressions are collected with a focus on the entire image context, resulting in expressions that may include more background information and irrelevant details. This dataset challenges algorithms to handle referring expressions that are not solely centered around the target objects, requiring a deeper understanding of the image context. Like RefCOCO, RefCOCO+ provides segmentation masks for the target objects.
    RefCOCOg sets itself apart from RefCOCO and RefCOCO+ by utilizing images sourced from Google search results. It consists of 95,010 referring expressions, corresponding to 49,822 objects. The referring expressions in RefCOCOg are collected independently of the images, resulting in more natural and diverse expressions. Unlike the other two datasets, RefCOCOg includes both object categories and stuff categories, such as ”sky” and ”grass,” expanding the scope of referring expression comprehension. However, RefCOCOg only provides bounding box annotations for the target objects and does not include pixel-level segmentation masks.
    These three datasets offer distinct challenges and opportunities for researchers working on referring expression comprehension.
  • GRIT (Gupta et al., 2022). The GRIT dataset was meticulously designed with specific principles in mind to address key challenges in computer vision research. The dataset creation process focused on selecting unambiguous tasks with clear definitions and unambiguous ground truth, similar to the approach taken in the GLUE benchmark. By including a variety of tasks such as object categorization, object localization, referring expressions, visual question answering, semantic segmentation, human keypoint estimation, and surface normal estimation, GRIT covers a broad range of visual skills. One of the unique features of GRIT is its evaluation of generalization to new data sources and concepts, robustness to image perturbations, and calibration measures, which are crucial for developing more flexible and general computer vision systems. Additionally, GRIT offers two tracks, Restricted and Unrestricted, to support the development of large-scale models and fair comparisons between models with varying compute resources. Overall, the GRIT dataset serves as a unified platform for assessing the overall capability of computer vision systems across diverse concepts and data sources, aiming to drive the development of performant and robust general-purpose vision systems.
Instruction Following Evaluation Datasets
Instruction Following Evaluation Datasets assess models’ ability to understand and execute natural language instructions in various domains. These datasets consist of high-level instructions paired with corresponding actions, challenging models to interpret complex instructions, reason about necessary steps, and generate appropriate outputs. They advance research in natural language processing, robotics, and embodied AI, providing benchmarks for evaluating approaches and techniques. They contribute to developing intelligent agents that can assist humans, enhance productivity, and enable more natural human-machine interactions.
  • TouchStone (Bai et al., 2023b). The TouchStone dataset creation process involved curating a diverse visual dialogue dataset encompassing five major categories of abilities and 27 subtasks. This dataset was carefully constructed to cover not only basic recognition and comprehension but also extend to literary creation and analysis abilities. One key feature of the TouchStone dataset is the integration of detailed image annotations, which enable the transformation of multimodal input content into a form understandable by advanced language models. This innovative approach allows for the direct evaluation of the quality of multimodal dialogue without the need for human intervention. By providing a comprehensive assessment of the capabilities of multimodal language models, the TouchStone dataset addresses the challenge of evaluating large vision-language models in open-ended dialogues, offering a standardized and objective method for assessing performance.
  • SEED-Bench (Li et al., 2023i). SEED-Bench dataset was meticulously curated through a multi-step process involving data collection, human annotation, and automated filtering mechanisms. This dataset stands out due to its comprehensive coverage of 12 evaluation dimensions, encompassing both spatial and temporal understanding in images and videos. By providing 19K multiple-choice questions with accurate annotations, SEED-Bench offers a robust platform for evaluating the generative comprehension capabilities of Multimodal Large Language Models (MLLMs). Researchers can leverage this dataset to address critical issues related to model performance, comprehension accuracy, and the effectiveness of generative models in processing multimodal inputs.
  • MME (Fu et al., 2024). The MME dataset in the article was meticulously created through manual design of instruction-answer pairs to prevent data leakage and ensure fair comparison of MLLMs. The dataset features a total of 14 subtasks covering perception and cognition abilities, including tasks like text translation, code reasoning, and arithmetic problems. By providing concise instructions and utilizing real photographs for data collection, the MME dataset addresses the need for a comprehensive evaluation of MLLMs, highlighting areas for model improvement and guiding future development in the field.
  • LLaVAW (Liu et al., 2024b) The ”LLaVA-Bench (In-the-Wild)” dataset serves as an evaluation benchmark to test the capability of models on complex reasoning in diverse environments. It comprises 24 images with 60 questions, featuring a variety of visual content including indoor and outdoor scenes, memes, paintings, and sketches. Each image is paired with a detailed, manually-curated description and a set of carefully selected questions. This benchmark focuses on evaluating the generalizability and adaptability of models to novel domains and their ability to handle complex reasoning across varied visual contexts.
Other MM Evaluation Datasets
Evaluation Datasets are not limited to the aforementioned aspects. Other datasets can provide a more comprehensive understanding of the capabilities of large models from various perspectives.
  • HM (Kiela et al., 2020). The Hateful Memes dataset was created as a challenge set to detect hate speech in multimodal memes. The dataset was carefully constructed to be challenging for unimodal architectures by including ”benign confounders” that can flip the label of a hateful meme. It consists of examples with high-confidence ratings from trained annotators and is balanced to include various multimodal fusion problems. This dataset aims to address the societal problem of hate speech by requiring reasoning about subtle cues in memes. It provides insights into the distribution of protected categories, types of attacks, and lexical statistics, highlighting the need for improvements in machine learning systems to tackle hate speech effectively.
  • MMB (Liu et al., 2023a) The ”MMBench” dataset is a comprehensive benchmark designed for evaluating large vision-language models (VLMs), consisting of around 3,000 single-choice questions that assess 20 distinct abilities ranging from object localization to social reasoning. Questions and choices are drawn from a mix of custom-generated and ChatGPT-assisted content, with images sourced from datasets like COCO-Caption and Places. MMBench utilizes a ”Circular Evaluation” strategy, which tests models’ robustness by rotating answer choices across multiple rounds. This method enhances the reliability of assessing models’ perception, reasoning, and instruction-following abilities. MMBench aims to provide detailed insights into AI models’ capabilities, addressing the limitations of traditional benchmarks by focusing on fine-grained skills and employing AI for precise answer matching, thereby advancing the field of AI by delineating the strengths and weaknesses of current multimodal models.
A.3.2. Video Data Multimodal Evaluation
In this subsubsection, we summarize commonly used evaluation datasets for video-text multimodal large language models (LLMs). Since many datasets are versatile enough to handle multiple evaluation tasks, our classification is provided for reference purposes only.
Video Question Answering
Video question answering evaluation can evaluate the model’s video understanding and question answering ability.
  • MSVD (Chen and Dolan, 2011). The MSVD dataset consists of 2,089 video segments and 85,550 English descriptions, with an average of 41 descriptions produced for each video, and at least 27 descriptions for over 95% of the videos. Even when focusing on descriptions from Tier-2 tasks, there are still 16 descriptions on average for each video, with at least 12 descriptions for over 95% of the videos. This dataset was constructed by asking annotators to watch short video clips and describe the main action or event in one sentence. The data collection framework used in this study involved deploying tasks on Amazon’s Mechanical Turk platform, where video segments were selected from YouTube. The highly parallel nature of the data allows for the creation of monolingual parallel training data, which can be beneficial for tasks such as paraphrase evaluation.
  • MSR-VTT (Xu et al., 2016). The MSR-VTT dataset is a significant contribution to the field of video understanding and language translation. It comprises 10K web video clips totaling 41.2 hours of content, with each clip annotated with about 20 natural sentences. The dataset was constructed by collecting 257 popular queries from a commercial video search engine, with 118 videos for each query, and annotating the clips with diverse and comprehensive categories. This large-scale dataset addresses the limitations of existing benchmarks by providing a wide variety of video content and associated language descriptions, enabling researchers to explore and develop more advanced algorithms for tasks such as video-to-text translation and video understanding.
  • TGIF-QA (Jang et al., 2017). The TGIF-QA dataset comprises 165,165 QA pairs extracted from 71,741 animated GIFs. The dataset was constructed by employing a combination of crowdsourcing and template-based approaches, generating questions based on video content. It introduces unique tasks for video VQA, such as counting repetitions of actions, identifying repeated actions, and recognizing state transitions in videos. These tasks require spatio-temporal reasoning and offer a new perspective on visual question answering in the video domain.
  • ActivityNet-QA (Yu et al., 2019). The ActivityNet-QA dataset consists of 58,000 QA pairs derived from 5,800 complex web videos. The dataset was constructed by sampling videos from the larger ActivityNet dataset and generating QA pairs for each video. This dataset can be utilized to address challenges in VideoQA, enabling researchers to explore various video representation strategies and improve the performance of question answering systems for long web videos.
  • LSMDC (Rohrbach et al., 2017). Large Scale Movie Description Challenge The LSMDC dataset consists of 118,114 sentences aligned with video clips from 202 movies. This dataset was constructed by collecting transcribed Audio Descriptions (ADs) and movie scripts, providing a rich source of visual and textual information for research purposes. Researchers can leverage this dataset to address challenges in generating video descriptions, understanding the relationship between visual content and language, and developing technologies to assist visually impaired individuals in accessing movie content.
  • MoVQA (Rohrbach et al., 2017). The MoVQA dataset consists of 21,953 manually annotated question-answer pairs sourced from 100 movies of diverse genres. The dataset is designed to assess model capabilities in long-form video understanding across different temporal lengths, with questions categorized into six types: information synopsis, temporal perception, spatial perception, causal reasoning, hypothetical reasoning, and external knowledge. By segmenting movies into single-scene, multi-scene, and full-scene levels, MoVQA provides a comprehensive platform for integrating multimodal information and addressing complex cognitive challenges in movie comprehension.
Video Captioning and Video Retrieval
Video captioning and video retrieval datasets serve as benchmarks for evaluating a model’s understanding of video content.
  • DiDeMo (Anne Hendricks et al., 2017). The DiDeMo dataset consists of over 10,000 unedited personal videos with 3-5 pairs of descriptions and distinct moments per video. The dataset was constructed in two phases: first, annotators watched videos, selected moments, and described them to ensure another user could select the same moment based on the description. Second, collected descriptions were validated by annotators marking corresponding moments in the videos. This dataset addresses the challenge of localizing moments in video with natural language by providing pairs of localized video segments and referring expressions that uniquely identify specific moments in the videos.
  • VATEX (Wang et al., 2019). The VATEX dataset comprises over 41,250 videos paired with 825,000 captions in both English and Chinese, including more than 206,000 English-Chinese parallel translation pairs. The dataset was constructed by collecting English and Chinese descriptions for video clips from the Kinetics-600 validation and holdout test sets using Amazon Mechanical Turk. This dataset enables researchers to address various challenges in multilingual visual understanding, such as multilingual video captioning, enhancing video dynamics comprehension, and exploring the interaction between videos and multilingual knowledge. Additionally, the dataset supports tasks like visual question answering, natural language moment retrieval, and multilingual visual concept clustering, facilitating large-scale training for multilingual, multimodal research endeavors.
Other Video Evaluation Datasets
Here, we summarize other evaluation datasets. Some are very comprehensive, while others offer unique video evaluations.
  • MVBench (Li et al., 2023h). The MVBench dataset comprises 20 challenging video tasks that require both spatial and temporal understanding. It was constructed using a static-to-dynamic method, which adapts static image tasks with dynamic goals to systematically design temporal tasks for video comprehension. This dataset can be utilized to evaluate the comprehension capabilities of Multi-modal Large Language Models (MLLMs) in open-world temporal understanding, ranging from perception to cognition. It also includes 1.9 million diverse instructional data points, which can be utilized to enhance a model’s ability to follow instructions.
  • EgoSchema (Mangalam et al., 2024). The EgoSchema dataset comprises over 5000 very long-form video language understanding questions derived from 250 hours of diverse egocentric video data. This dataset was manually curated to include questions that require choosing the correct answer from five choices based on a three-minute video clip, with a median certificate length of about 100 seconds. EgoSchema stands out from other video understanding datasets by featuring significantly longer temporal certificates, challenging state-of-the-art models that struggle in zero-shot evaluation scenarios. The dataset serves as a benchmark for evaluating and improving the capabilities of modern vision and language systems in comprehending complex and extended video content.
  • VideoChatGPT (Maaz et al., 2023). The dataset used for evaluating VideoChatGPT consists of 100,000 video-instruction pairs. This dataset was constructed through a combination of human-assisted annotation, where expert annotators provided detailed descriptions of video content, and semi-automatic annotation, which generated high-volume annotations using state-of-the-art vision-language models. The dataset is diverse and comprehensive, covering a wide range of data types such as detailed descriptions, question-answer pairs, and tasks stimulating creativity. This extensive dataset is crucial for training VideoChatGPT to comprehend video content effectively, integrating spatial and temporal cues into its understanding. It can be utilized to address challenges in video-based conversation models, such as improving temporal understanding, spatial consistency, and contextual comprehension.
  • Charade-STA (Gao et al., 2017). The Charades-STA dataset consists of around 10,000 videos, each with temporal activity annotations from 157 activity categories and multiple video-level descriptions. To address the need for clip-level sentence annotations in Temporal Activity Localization via Language Query (TALL), the Charades-STA dataset was constructed by generating sentence temporal annotations for the original Charades dataset. This involved connecting consecutive sub-sentences within each video to create complex queries, ensuring the time span of each query was less than half of the video length. The dataset contains 13,898 clip-sentence pairs in the training set, 4,233 pairs in the test set, and 1,378 complex sentence queries. This dataset enables researchers to explore the potential of the CTRL framework in handling longer and more complex sentences for temporal activity localization tasks.
  • QVHighlight (Lei et al., 2021). The QVHIGHLIGHTS dataset consists of over 10,000 YouTube videos that cover a wide range of topics, such as everyday activities, travel, social and political events. Each video in the dataset is annotated with a human-written natural language query, relevant moments in the video corresponding to the query, and saliency scores for all query-relevant clips. This comprehensive annotation allows for the development and evaluation of systems that can detect relevant moments and salient highlights for diverse user queries. The dataset was constructed to address the challenge of the lack of annotated data for detecting moments and highlights in videos, providing a valuable resource for research in this area.
A.3.3. Audio Evaluation Datasets
In this section, we give a comprehensive summary of commonly used evaluation datasets for audio-text multimodal large language models (LLMs).
Audio-Text Retrieval
Audio retrieval datasets serve as benchmarks for evaluating a model’s understanding of audio content.
  • AudioCaps (Kim et al., 2019). The AudioCaps dataset consists of 46,000 pairs of audio clips and human-written text descriptions. This dataset was constructed by collecting data through crowdsourcing on the AudioSet dataset, ensuring that the captions are faithful to the audio inputs. The dataset can be utilized to address the task of audio captioning for sound in the wild, providing a valuable resource for developing and evaluating models that generate natural language descriptions for audio content.
  • Clotho (Drossos et al., 2020). The Clotho dataset for audio captioning consists of 4981 audio samples ranging from 15 to 30 seconds in duration and 24,905 captions with lengths between eight to 20 words. The dataset was constructed by collecting audio samples from the Freesound platform, processing them to meet specific criteria, and crowdsourcing captions using Amazon Mechanical Turk and annotators from English-speaking countries. Clotho aims to address the task of general audio content description through free text, providing a diverse set of audio samples and captions for research in audio captioning methods.
Audio Question Answering
Audio question answering evaluation can evaluate the model’s audio understanding and question answering ability.
  • ClothoAQA (Lipping et al., 2022). The Clotho-AQA dataset consists of 1991 audio files, each lasting between 15 to 30 seconds, selected from the Clotho dataset. For each audio file, six different questions were collected through crowdsourcing using Amazon Mechanical Turk, with corresponding answers provided by different annotators. Two of the questions for each audio file are designed to have ’yes’ or ’no’ answers, while the remaining questions have single-word answers. This dataset was created to address the task of audio question answering, where a system analyzes both the audio signal and a natural language question to generate a relevant natural language answer.
  • Audio-MusicAVQA (Li et al., 2022c). The MUSIC-AVQA dataset in the article comprises over 45,000 question-answer pairs derived from more than 9,000 videos, totaling over 150 hours of content. The dataset was constructed by collecting real musical performance videos from YouTube, covering solo performances, ensembles of the same instrument, and ensembles of different instruments. The dataset includes 33 question templates across 9 question types, facilitating spatio-temporal reasoning over audio-visual scenes. Researchers can leverage this dataset to explore multimodal scene understanding, study audio-visual interactions, and advance spatio-temporal reasoning in dynamic audio-visual scenarios.