### v2

# A Survey of Multimodal Large Language Model from A Data-centric Perspective

Abstract.
Multimodal large language models (MLLMs) enhance the capabilities of standard large language models by integrating and processing data from multiple modalities, including text, vision, audio, video, and 3D environments. Data plays a pivotal role in the development and refinement of these models. In this survey, we comprehensively review the literature on MLLMs from a data-centric perspective. Specifically, we explore methods for preparing multimodal data during the pretraining and adaptation phases of MLLMs. Additionally, we analyze the evaluation methods for the datasets and review the benchmarks for evaluating MLLMs. Our survey also outlines potential future research directions. This work aims to provide researchers with a detailed understanding of the data-driven aspects of MLLMs, fostering further exploration and innovation in this field.
Large Language Models; Generative Models; Multimodal Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Computer vision
1. introduction
In recent years, we have witnessed rapid advancement of large language models (LLMs) and multimodal large language models (MLLMs) (Zhao et al., 2023c; Wu et al., 2023b). MLLMs, such as GPT-4 (OpenAI, 2023), Flamingo (Alayrac et al., 2022), LLaVA (Liu et al., 2024b), BLIP2 (Li et al., 2023e), Cambrian-1 (Tong et al., 2024) and X-InstructBLIP (Panagopoulou et al., 2023), integrate multiple modality information, demonstrating impressive comprehension and generation capabilities. These models achieve competitive performance in traditional multimodal tasks, such as visual recognition (Zhang et al., 2024a), video understanding (Xu et al., 2021; Tang et al., 2023), speech recognition (Min and Wang, 2023) and 3D understanding (Guo et al., 2023; Hong et al., 2023). Moreover, their excellent language understanding capacity enables strong performance in text-rich tasks, such as question answering (Hu et al., 2024), multi-dialog conversation and logical reasoning (Xu et al., 2024; Li et al., 2023k).
Most existing MLLMs focus on modifying model architecture to explore the use of information from modalities (Xu et al., 2024; Panagopoulou et al., 2023; Alayrac et al., 2022). While model effectiveness is crucial, data also significantly impacts the success of MLLMs. For example, Hoffmann et al. (2022) shows that in order to scale up models, it is necessary to increase the scale of the training data. Beyond data volume, data quality is equally important. Previous research (Sorscher et al., 2022) indicates that carefully curated datasets can enable smaller models to achieve comparable performance to larger ones. However, comprehensive studies on data curation and utilization for MLLMs are still lacking. Therefore, this study aims to provide a comprehensive understanding of MLLMs from a data-centric perspective.
In contrast to model-centric approaches that prioritize architectural enhancements while relying on fixed datasets, data-centric perspectives emphasize the comprehensive impact of the training corpus datasets on the model performance. Within the scope of data-centric MLLMs, our focus lies on leveraging the heterogeneous nature of data modalities, enhancing data structure, increasing data quantity, and elevating data quality to augment MLLMs (Zha et al., 2023). Our survey discusses three key questions from a data-centric perspective at different stages of MLLMs:
  • Q1: How to collect, select, and manage data for MLLMs? The substantial data volume requirements and the heterogeneity of multimodal data pose challenges in gathering, selecting, and effectively managing data for model training. Different training stages of MLLMs also lead to varying data type requirements.
  • Q2: How data affects the performance of MLLMs? Understanding the relationship between data characteristics and the performance of MLLMs is crucial for optimizing datasets and enhancing model capabilities.
  • Q3: How to evaluate the data for MLLMs? It is necessary to develop comprehensive evaluation benchmarks to assess the performance and robustness of MLLMs across diverse tasks.
Unique contribution of this survey. Several existing works have focused on LLMs (Zhao et al., 2023c; Hadi et al., 2023; Naveed et al., 2023) and MLLMs (Wu et al., 2023b; Zhang et al., 2024c) from a model-centric perspective, but lack an in-depth analysis from the data-centric perspective. Recently, some work has started to focus on data preparation for LLMs, such as data management methods (Wang et al., 2023b), data selection methods (Albalak et al., 2024), and comprehensive reviews of LLM datasets (Liu et al., 2024a). However, such work primarily focuses on data management and selection methods for text-only LLMs and does not provide a detailed analysis of the data processing pipeline for MLLMs. The most closely related work to ours is data-centric artificial intelligence (DCAI) (Zha et al., 2023; Jarrahi et al., 2022; Jakubik et al., 2022; Polyzotis and Zaharia, 2021; Whang et al., 2023), which also focuses on data-centric views of AI research but does not specifically analyze LLMs and MLLMs.
With the rapid growth of MLLMs and the increasingly important role of data in this large model era, we believe it is crucial to provide a comprehensive overview of data-centric approaches for MLLMs. This survey aims to thoroughly review the literature on the advances of MLLMs from a data-centric perspective and discusses open issues or future directions in this field.
In this survey, we reviewed the literature on the advances of MLLMs from data-centric perspectives. We provide researchers and developers with a general and comprehensive understanding of the latest developments in the field of MLLM data side. The key contributions of this survey are summarized as follows:
  • New data-centric perspective. We provide a comprehensive review of MLLMs from a data-centric perspective, considering modalities such as text, image, video, and audio.
  • Data preparation and management pipeline. We summarize the data preparation and management pipeline for MLLMs in both the pre-training and adaptation phases.
  • Data evaluation methods and evaluation benchmarks. We outline commonly used methods to evaluate datasets as well as MLLMs evaluation benchmarks from a data-centric perspective.
  • Open issues and future directions. We discuss open issues in current research on data-centric LLMs and propose several future research directions.
Figure 1. Overview of the data pipeline for MLLMs.
[…]
[Overview of data pipeline for MLLMs]Overview of data pipeline for MLLMs
The rest of this survey is organized as follows: Section 2 introduces the preliminaries for LLMs and MLLMs, and discusses the motivations to analyze them from a data-centric perspective. Sections 3 to 5 summarize the main stages of collecting, processing, and selecting data for MLLMs training. Section 6 enumerates the data evaluation methods and existing evaluation datasets for MLLMs according to evaluation tasks. Section 7 discusses open issues and highlights several future research directions in this field. Finally, we conclude the survey in Section 8.
2. backgrounds and categorization
2.1. Large Language Models
The field of natural language processing (NLP) has witnessed a remarkable evolution in language modeling techniques, culminating in the development of LLMs. The journey began with statistical methods (Jelinek, 1998; Gao and Lin, 2004; Rosenfeld, 2000), which laid the foundation for language modeling by capturing the probabilistic distributions of words and phrases. Subsequently, neural network approaches emerged (Bengio et al., 2000; Mikolov et al., 2010; Kombrink et al., 2011), leveraging the power of deep learning to learn more complex and abstract representations of language. However, such models were typically designed for specific tasks and encountered challenges related to scalability and generalization.
A significant breakthrough came with the introduction of pre-trained language models, which aimed to capture general language representations by training on vast amounts of unlabeled text data (Kenton and Toutanova, 2019; Radford et al., [n. d.]). Building upon these advancements, LLMs have taken the concept of pre-training to an unprecedented scale, with models containing billions of parameters and trained on massive corpora spanning hundreds of gigabytes to terabytes of text data (Brown et al., 2020; Raffel et al., 2020).
LLMs exhibit several notable properties that distinguish them from previous language models. One fundamental characteristic is the scaling law, which describes how the performance of large language models improves as they scale in terms of model size, training data, and computational resources (Kaplan et al., 2020; Hoffmann et al., 2022). This power-law relationship suggests that larger models trained on more data tend to capture more complex patterns and generalize better to new tasks. Another intriguing property of LLMs is the emergence of abilities that were not explicitly trained for, often referred to as emergent abilities (Wei et al., 2022). This ability suggests that LLMs can capture and leverage complex linguistic patterns and knowledge from the pre-training data, enabling them to perform tasks beyond their original training objective. Furthermore, LLMs have demonstrated the capability of in-context learning (Brown et al., 2020), where they can perform tasks based on a few examples provided in the input prompt, without the need for explicit fine-tuning. This highlights the models’ ability to rapidly adapt to new tasks and generalize from limited examples.
The evolution of LLMs typically involves three key stages: training, adaptation, and evaluation (Zhao et al., 2023c). The training stage focuses on learning general language representations from large-scale unlabeled corpora, capturing the underlying patterns and structures of natural language. The adaptation stage involves guiding the pre-trained model to specific tasks and human preference, often through supervised finetuning and human preference alignment (Raffel et al., 2020; Ouyang et al., 2022). This step is crucial for optimizing the model’s performance on the target application and ensuring its outputs are aligned with the desired objectives and human preference. Finally, the evaluation stage involves assessing the performance of the adapted model through various metrics and benchmarks to ensure that it meets the required standards and performs effectively in real-world scenarios.
2.2. Multimodal Large Language Models
Multimodal large language models (MLLMs) extend the capabilities of traditional LLMs by leveraging their comprehension, generation, and task-solving abilities across different modalities, including natural language (text and audio) and visual information (videos, images, and 3D models). The development of MLLMs is driven by the need to address real-world problems that often involve multiple modalities. For instance, in robotics, MLLM can process visual input from cameras, interpret natural language instructions, and analyze structured data from sensors to perform complex tasks (Zeng et al., 2023). Similarly, in healthcare, MLLMs can analyze medical images, process electronic health records, and interpret patient-doctor conversations to aid in diagnosis and treatment planning.
Multimodal large language models (MLLMs) typically consist of three primary components: a modality encoder, a projector, and a large language model (LLM). The modality encoder is responsible for encoding data from different modalities, the projector aligns the data from various modalities with the LLM, and the LLM serves as the backbone, providing comprehension capabilities. By integrating LLMs with multimodal encoders and projectors, MLLMs can process various types of information, demonstrating strong understanding and analytical capabilities to address downstream tasks across various modalities (Wu et al., 2023b).
MLLMs represent a significant advancement in AI by extending the capabilities of traditional LLMs to process and analyze multiple modalities. MLLMs can solve complex real-world problems that require understanding and reasoning across different types of information. As research in this field progresses, we can expect MLLMs to play an increasingly important role in various domains, from robotics and healthcare to education and entertainment.
Data-Centric MLLMsData Collectingand ProcessingData CollectingSourcesCommonWebpages CommonCrawl Project (Raffel et al., 2020), Self-Crawled Data: General Approches (Sharma et al., 2018; Changpinyo et al., 2021), Focused Approaches (Wikipedia (Gao et al., 2020; Srinivasan et al., 2021; Schamoni et al., 2018), Flickr (Young et al., 2014), FreeSound (Mei et al., 2023; Fonseca et al., 2022)), Querying search engines-based approaches (Wu et al., 2017; Gu et al., 2022) Social Media Stack Exchange (Soboleva et al., 2023; Chen et al., 2023g), Reddit (Baumgartner et al., 2020; Desai et al., 2021; Overbay et al., 2023), YouTube (Miech et al., 2019; Zhou et al., 2017), X (Barbieri et al., 2020) Academic Papers ArXiv (Computer, 2023; Gao et al., 2020; Liu et al., 2023d), S2ORC (Lo et al., 2020; Soldaini and Lo, 2023; Soldaini et al., 2024) Books Project Gutenberg (Rae et al., 2019), Smashwords (Zhu et al., 2015), Bibliotik (Gao et al., 2020), Amazon.com (Mishra et al., 2019), Old Photo Books (Okamoto et al., 2023) Codes BigQuery (Nijkamp et al., 2022), GitHub (Kocetkov et al., 2022; Nijkamp et al., 2022; Gao et al., 2020), The Stack (Li et al., 2023a) ProfessionalSources Legal Domain: FreeLaw (Gao et al., 2020); Math Domain: Khan Academy, DeepMind (Saxton et al., 2018); Medical Domain: Online Medical Websites (Li et al., 2023i; Ben Abacha and Demner-Fushman, 2019), Knowledge Bases (Li et al., 2023i; Welbl et al., 2018), In-Hospital Database Systems (Johnson et al., 2023, 2019; Bustos et al., 2020); Financial Domain: Financial Platforms and Forums (Ok, 2023; Lu et al., 2023b) Data ProcessingMethodsFiltering Textual Filtering (Joulin et al., 2016; Grave et al., 2018; Raffel et al., 2020, 2020; Yuan et al., 2021; Gunasekar et al., 2023), Image-Level Filtering (Zhu et al., 2024; Guo et al., 2021; Yang et al., 2022), Video-Level Filtering (Wang et al., 2023a; Blattmann et al., 2023; Farnebäck, 2003; Baek et al., 2019) Deduplication Exact Duplication (Raffel et al., 2020; Suárez et al., 2019; Lee et al., 2022), Approximate Duplication (Computer, 2023; Lee et al., 2022; Penedo et al., 2023), Sentence-Level and Document-Level Semantic Duplication (Reimers and Gurevych, 2019; Song et al., 2020; Silcock et al., 2022; Abbas et al., 2023) DataEnhancement Enhance X Modality Data (Li et al., 2023j; Luo et al., 2024; Wei et al., 2023), Enhance Text Data (Nguyen et al., 2023; Fan et al., 2024; Lai et al., 2023a; Chen et al., 2023e) Commonly-Used DatasetsImage datasets Image-Caption Datasets (Schuhmann et al., 2021, 2022, 2022), Content Descriptive Image-Caption Dataset (Chen et al., 2015; Young et al., 2014; Chen et al., 2022), Interleaved Image-Text Documents (Alayrac et al., 2022; Zhu et al., 2024; Laurençon et al., 2024), Visual Question Answer Image-Caption Datasets (Goyal et al., 2019; Zhu et al., 2016; Biten et al., 2019) Video datasets Video Captioning Datasets (Xu et al., 2016; Bain et al., 2021; Alayrac et al., 2022), Video Question Answering Datasets (Chen and Dolan, 2011; Xu et al., 2016; Jang et al., 2017), Interleaved Video-Text Dataset (Sun et al., 2023b) Audio datasets Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022b), Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) 3D datasets 3D Captioning Datasets (Dai et al., 2017; Armeni et al., 2017; Zheng et al., 2020) Data-CentricPre-trainingDomain Mixture Pure Text Domain Mixture (Xie et al., 2024a, b; Du et al., 2022), Multimodal Domain Mixture (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a) Modality Mixture Image Caption Data Mixture (McKinzie et al., 2024; Lai et al., 2023b), Image and Video Caption Data Mixture (Lin et al., 2023b; Zhang et al., 2023; Luo et al., 2023; Jin et al., 2023; Li et al., 2023f; Han et al., 2023; Li et al., 2023c, g; Wang et al., 2023a; Chen et al., 2023c) Quality Selection Active Learning-Based Selection (Xu et al., 2023a), Selection Before Training (Gadre et al., 2024), Distribution-Agnostic Selection (Gadre et al., 2024; Wang et al., 2024c; Mahmoud et al., 2024), Distribution-Aware Selection (Wang et al., 2024a; Gadre et al., 2024) Data-Centric AdaptationSupervisedFinetuningData Processing SFT Data Processing (Zhu et al., 2023; Chen et al., 2023e) Data Generation Captioning Instruction-Response Datasets Generation (Zhu et al., 2023; Xu et al., 2024; OpenAI, 2023), Question Answer Instruction-Response Datasets Generation (Liu et al., 2023c; Singh et al., 2019; Li et al., 2023k), Reasoning Instruction-Response Datasets Generation (Chen et al., 2023j, h, m), Other Instruction-Response Datasets Generation (Li et al., 2023k; Panagopoulou et al., 2023; Dai et al., 2024) Data Selection Coreset-Based Selection (Chen et al., 2023k; Das and Khetan, 2023; Sorscher et al., 2022), LLMs-Based Selection (Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e), gradient-Based Selection (Paul et al., 2021; Attendu and Corbeil, 2023; Xia et al., 2024), Self-Instruction-Based Selection Methods (Li et al., 2023l, d; Kung et al., 2023; Liu et al., 2024c) Alignment Human Preference Alignment: Pure Text (Ouyang et al., 2022; Li et al., 2023f), Multimodal (Liu et al., 2023b; Sun et al., 2023a; Chen et al., 2023i; Ouyang et al., 2022) Data-CentricEvaluationData Evaluation Dataset Diversity (Heusel et al., 2017; Friedman and Dieng, 2023); Dataset Quality (Honovich et al., 2022a; Rohrbach et al., 2018; Jing et al., 2023); Dataset Similarity (Sun and Saenko, 2016; Jiang et al., 2022; Pillutla et al., 2023) Evaluation DatasetsCaptioningTasks Image Captioning Datasets (Goyal et al., 2019; Hudson and Manning, 2019; Marino et al., 2019), Video Captioning Datasets (Xu et al., 2017; Chen and Dolan, 2011; Yu et al., 2019), Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022b) QuestionAnswering Tasks Image Question Answering Datasets (Chen et al., 2015; Young et al., 2014; Karpathy and Fei-Fei, 2015), Video Question Answering Datasets (Chen and Dolan, 2011; Venugopalan et al., 2015; Xu et al., 2016), Audio Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) Perception &Reasoning Tasks Perception and Reasoning Datasets (Singh et al., 2019; Masry et al., 2022; Kembhavi et al., 2016), Spatial Reasoning Datasets (Liu et al., 2023b; Kazemzadeh et al., 2014; Mao et al., 2016) Other NoteworthyTasks Classification Datasets (Kiela et al., 2020; Deng et al., 2009; Piczak, 2015; Wu et al., 2015; Goyal et al., 2017), Multimodal Dialogue Datasets (Bai et al., 2023b)
Figure 2. Overview of Data-Centric MLLMs
[…]
[Overview of Data-Centric MLLMs]Categorization of Data-Centric MLLMs work
2.3. Data-Centric AI and Why Data-Centric MLLMs
The field of artificial intelligence (AI) has experienced a paradigm shift towards data-centric approaches in recent years (Zha et al., 2023). Data-centric AI emphasizes the critical role of data quality, diversity, and representativeness in building robust and effective AI systems. This shift acknowledges that the performance of AI models is heavily based on the quality and characteristics of the data used for training, rather than solely focusing on algorithmic improvements.
In the context of LLMs and MLLMs, data play a pivotal role in determining their capabilities and limitations. LLMs, such as GPT-3 (Brown et al., 2020) and BERT (Kenton and Toutanova, 2019), are trained on vast amounts of textual data to learn the intricacies of language and generate coherent, contextually relevant outputs. Similarly, MLLMs, which integrate multiple modalities such as text, images, and speech, require diverse and well-aligned datasets in different modalities to learn meaningful cross-modal representations (Gadre et al., 2024). The quality, diversity, and representativeness of the training data directly impact the models’ ability to understand and generate language, as well as their capacity to reason and perform tasks across multiple modalities (meta llama, 2024; Liu et al., 2023c; Chen et al., 2023e).
Analyzing LLMs and MLLMs from a data-centric perspective offers several advantages. First, it enables researchers to identify and address potential biases and limitations in the training data that can propagate to the models’ outputs. By carefully curating and augmenting datasets to ensure diversity and representativeness, researchers can mitigate biases and improve the fairness and generalizability of the models (Chen et al., 2024a). Second, a data-centric approach allows for a deeper understanding of the models’ capabilities and limitations based on the characteristics of the training data (McKinzie et al., 2024). By systematically varying data properties and evaluating model performance, researchers can gain insights into the specific data attributes that contribute to the models’ success or failure in various tasks. This understanding can guide the development of more efficient and effective data collection and curation strategies. Moreover, adopting a data-centric perspective in the development of LLMs and MLLMs opens up opportunities for data-efficient training and adaptation (Liu et al., 2024b, 2023c; Chen et al., 2023e). Additionally, data-centric approaches facilitate the development of more interpretable and explainable models by linking model behaviors to specific data characteristics.
2.4. Categories of Data-Centric MLLMs
In this survey, we review previous work on MLLMs from a data-centric perspective. Our taxonomy is primarily based on different stages of MLLM development, as illustrated in Figure 2. Specifically, we organize our article around a data pipeline, shown in Figure 1, where the data pipeline for MLLMs comprises three stages: pre-processing (including data collection and data processing in Section 3), pre-training data processing in Section 4), and adaptation data processing in Section 5. Along this pipeline, we discuss data-related model performance and analyze how dataset curation affects model performance. During the pre-processing stage, large-scale multimodal data are collected from various sources and constructed into datasets. Processing steps such as filtering, deduplication, and data enhancement are applied to improve data quality. In the pre-training stage, data selection, domain mixing and modality mixing are performed to choose appropriate data for model training. In the adaptation stage, we focus on data processing and generation for supervised fine-tuning and human preference alignment. Furthermore, we discuss data evaluation metrics and evaluation datasets for MLLMs in Section 6, examining their construction and main characteristics. Finally, we suggest possible future directions for further research on data-centric MLLMs in Section 7.
3. Data collecting and processing
The first and fundamental step in training MLLMs is to collect and process sufficient data from various sources. In this section, we first introduce the sources for data collection. Next, we summarize the processing steps for multimodal data, including commonly used filtering, deduplication, and data enhancement methods. Finally, we provide an overview of commonly used datasets for MLLMs.
3.1. Data Collecting Sources
We first introduce the sources where researchers typically find raw data for MLLM pre-training. Commonly used data sources can be divided into six categories: common webpages, social media, academic papers, books, code repositories, and professional sources. Data from these different sources have various characteristics, each contributing uniquely to improving different abilities of the models.
Common Webpages.
Webpages are the main source for collecting large-scale training corpus. CommonCrawl project serves as the most commonly used start-point for large-scale webpages. It is instrumental in generating large-scale pre-training corpora like C4 (Raffel et al., 2020) for LLM pre-training. Beyond textual content, CommonCrawl’s vast archive of web pages, which includes numerous image-text pairs, has also become a vital resource for constructing multimodal pre-training datasets such as LAION-5B (Schuhmann et al., 2022).
Apart from CommonCrawl, several works focus on selecting specific webpages and crawling data independently. For textual datasets, initiatives like WuDaoCorpora (Yuan et al., 2021) crawl thousands of web pages to construct vast textual corpora. In the multimodal dataset sector, there are three main self-crawling approaches: general approach, focused approach on specific platforms, and querying search engines for images. Examples include AI Challenger Captions (Wu et al., 2017) and Wukong (Gu et al., 2022). General approaches, such as CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021), involve crawling billions of web pages without targeting specific platforms. Focused approaches, on the other hand, involve selecting specific webpages to crawl. For example, Wikipedia is a commonly used resource that serves as a rich and accurate source for text-based single-modal pre-training datasets like the Pile (Gao et al., 2020). It also plays a crucial role in creating multimodal pre-training datasets such as WIT (Srinivasan et al., 2021) and WikiCaps (Schamoni et al., 2018). For the vision modality, Flickr (Young et al., 2014) serves as an important source, containing photos and videos shared by online users. This contributes to image-text datasets such as Flickr30k (Young et al., 2014). For the audio modality, platforms such as BBC Sound Effects, FreeSound, and SoundBible are commonly used, contributing to datasets such as WavCaps (Mei et al., 2023) and FSD50K (Fonseca et al., 2022).
Social Media.
Social media platforms significantly enhance the training datasets for MLLMs by providing real-time, varied, and colloquial text data, as well as multimodal content that captures human expression and interaction. However, questions regarding the ownership and copyright of social media content are on the rise. MLLM researchers need to be aware of these issues and understand the rights and licensing of the data sources they use. Stack Exchange offers high-quality textual and visual content, hosting extensive forums with rich dialogues and images. In addition to Stack Exchange, Reddit offers user-generated content with discussions and voting on a wide array of topics, contributing to both textual and visual content. For video data, YouTube is a standout platform, with a vast amount of videos uploaded every minute. X, formerly known as Twitter, offers great amount of sharing text messages, images, audio, and videos.
Academic Papers.
Academic papers offer high-quality, authoritative content, enriching pre-training datasets with specialized knowledge and formal language, which is critical for developing language models with expertise in academic and professional domains. Academic papers are primarily utilized as sources for text-based pre-training datasets. arXiv offers a vast array of prepublished scientific research for training specialized language models. This is evidenced by its role in creating datasets such as RedPajama-Data-1T (Computer, 2023), Pile-arXiv (Gao et al., 2020), and image-text datasets such as MMC (Liu et al., 2023d). Another academic platform semantic scholars contribute to the S2ORC (Lo et al., 2020) corpus, which contains 81.1 million academic papers, offering metadata, abstracts, references, and full texts for 8.1 million open-access works.
Books.
Books provide a rich source of high-quality textual content for language model training and serve as fertile ground for multimodal pre-training datasets. They offer diverse visual data from book covers and illustrations, enriching models with a blend of literary depth and visual context.
Project Gutenberg acts as a pivotal source for language modeling and analysis with its extensive library of more than 70,000 free eBooks, fueling datasets such as PG-19 (Rae et al., 2019). Smashwords contributes to the well-known BookCorpus (Zhu et al., 2015) dataset, which encompasses a diverse collection of 11,038 self-published novels that span genres such as romance, science fiction, and fantasy. This extensive corpus has been instrumental in training landmark models such as GPT (Radford et al., [n. d.]) and BERT (Kenton and Toutanova, 2019). Bibliotik that consists of a mix of fiction and non-fiction books contributes to Books3 (Gao et al., 2020), part of the Pile dataset, which contains about 197,000 books. This dataset is essential for context modeling and narrative research, accounting for 2.1% of the RedPajama-Data-1T dataset.
Meanwhile, books also serve as an important source for constructing multimodal datasets. Book covers hold vast multimodal information, combining visual elements with metadata such as author names, titles, and genres. This fusion of data supports tasks like genre classification and visual question-answering. The OCR-VQA dataset (Mishra et al., 2019), sourced from Amazon.com, exemplifies how book covers can be leveraged to build rich pre-training datasets for MLLMs, demonstrating the value of covers in enhancing model understanding of visual and textual content. Old photo books offer a unique advantage for constructing multimodal pre-training datasets by providing historical visual and textual contexts, enriching models’ understanding of past societies and landscapes. For example, previous work collected 9,516 image-text pairs from 175 Japanese old photo books (Okamoto et al., 2023).
Domain-Specific Sources.
To enhance the performance of MLLMs in specific domains, general pre-training is supplemented with incremental, domain-specific pre-training. In the legal domain, pre-training data is primarily textual. FreeLaw stands out by offering free access to judgments from India’s higher courts, exemplified by the Pile-FreeLaw (Gao et al., 2020) dataset, which structures these judgments for machine learning. In the math domain, textual datasets primarily originate from Khan Academy’s exercises for foundational concepts and DeepMind’s Pile-DeepMind Mathematics (Saxton et al., 2018) for advanced, algorithmically generated problems. Additionally, DVQA (Kafle et al., 2018) introduces a multimodal aspect by combining bar charts with question-answer pairs generated through Matplotlib. In the medical domain, pre-training data comes from online medical websites, knowledge bases, and in-hospital database systems. Online medical websites such as Qianwen Health and PubMed contribute to datasets like Huatuo-26M (Li et al., 2023i) and MedQuAD (Ben Abacha and Demner-Fushman, 2019). Knowledge bases like Wikipedia also contribute to Huatuo-26M (Li et al., 2023i) and MedHop (Welbl et al., 2018). In-hospital database systems, including electronic health record systems (EHR) and ICU-specific clinical information systems, are used in MIMIC-IV (Johnson et al., 2023). Multimodal datasets in radiology, such as MIMIC-CXR-JPG (Johnson et al., 2019) and PADCHEST (Bustos et al., 2020), are generated from in-hospital radiology reports. In the financial domain, pre-training primarily relies on textual data sourced from both English and Chinese platforms. English contributions come from EDGAR and the SEC Financial Statement and Notes Data Sets, which enhance the FinTree project (Ok, 2023). Chinese sources include major sites like Sina Finance and Tencent Finance, as well as Eastmoney for specific documents, and forums like Guba and Xueqiu, culminating in the comprehensive BBT-FinCorpus (Lu et al., 2023b). This assortment of sources underscores the rich textual foundation for financial model pre-training.
3.2. Data Processing
3.2.1. Filtering
Data filtering in MLLM training is critical for enhancing model reliability and efficiency. Filtering out unwanted content is essential, as even minimal exposure to hate speech can negatively influence model behavior (Luccioni and Viviano, 2021; Gunasekar et al., 2023). For multimodal datasets, previous work has focused on filtering out unwanted data from different modalities separately. Specifically, for X-Text datasets, they consider filtering data based on both the text and the X modality independently.
Textual filters primarily encompass language and content filtering. For language filtering, documents or sentences that fall below a certain threshold for a specific language are removed. For English-only datasets, commonly used tools like Langdetect and FastText operate at the document level. For instance, C4 (Raffel et al., 2020) utilizes Langdetect to filter out any non-English pages with a probability below 0.99. Another renowned English language filter, FastText, is widely employed by datasets such as Dolma (Soldaini et al., 2024) and RefinedWeb (Penedo et al., 2023). CLUECorpus2020 (Xu et al., 2020) adopts a sentence-level filtering approach, selecting sentences whose language type is Chinese if a language is mentioned. For multilingual datasets, FastText, trained on Wikipedia data to classify 176 languages, remains a preferred choice, capable of processing 1,000 documents per second on a single CPU core. ROOTS (Laurençon et al., 2022) utilizes FastText for document-level language classification, resulting in a 1.6TB dataset containing 59 languages. For code datasets, detection methods are more elementary, often based on file extensions. For instance, in Palm (Chowdhery et al., 2023), files are filtered based on filename extensions to restrict to one of 24 common programming languages, resulting in 196GB of source code.
Another type of textual filtering is content filtering, which includes removing toxic and distracting content. Toxic content filtering targets text that is deemed rude, disrespectful, or unreasonable, employing heuristic and rule-based methods. For instance, in C4 (Raffel et al., 2020), pages containing words in a predefined ”Dirty, Naughty, Obscene, or Otherwise Bad Words” list are removed. Distracting content includes short or incomplete sentences, ”lorem ipsum” text, and useless identifiers like HTML, CSS, and JavaScript tags. Filtering methods usually involve multiple rule-based filters, such as punctuation-based segmentation in WudaoCorpora (Yuan et al., 2021), or machine learning approaches, as seen in phi-1 (Gunasekar et al., 2023), which use GPT-4 as an annotators and then train a high-quality content classifier based on the annotated results.
For image-level filtering, the most fundamental step is to remove images with excessively low resolution, as these images often fail to convey effective information. Additionally, it is necessary to filter out images with inappropriate aspect ratios, as such unconventional images frequently resemble banner-like advertisements (Zhu et al., 2024). Furthermore, similar to removing harmful content in text filtering, such as NSFW material, it is necessary to filter out such content from images. This can be achieved by training a binary classification model using appropriate datasets. Additionally, any potential occurrences of human faces or other sensitive elements in the images should be detected using a face detector and blurred accordingly (Guo et al., 2021). These filtering processes have a minimal effect on the overall model performance (Yang et al., 2022).
For video-level filtering, previous work has leveraged image or image-text filtering methods to handle static content and applied specific techniques for dynamic content. Video filtering involves four main components: scene transition detection, video quality and integrity improvement, refinement and coherence evaluation, and modality completeness. For scene transition detection, algorithms are used to remove abrupt scene changes, ensuring a cleaner dataset free from disruptive transitions (Wang et al., 2023a; Blattmann et al., 2023). To improve video quality and integrity, three aspects are considered: motion, textual information in the video, and resolution and frame rate. For motion analysis, dense optical flow (Farnebäck, 2003) is typically utilized to measure motion complexity and filter out static or repetitive sequences with limited educational value. To detect textual information in videos, OCR (Baek et al., 2019) is applied to identify and remove clips with substantial text, which can otherwise mislead the model’s visual interpretation. The resolution and frame rate of training video datasets are standardized, enabling model to learn from uniformly structured data (Wang et al., 2023a). In constructing video-text datasets, ensuring the relevance and clarity of annotations is paramount. Focused efforts are made to refine the textual content of video-text pairs, particularly subtitles, to ensure they are contextually meaningful and devoid of promotional or irrelevant material (Xu et al., 2023c). Models such as CLIP are employed to evaluate the coherence between video frames and the accompanying text (Chen et al., 2024b; Xu et al., 2023c; Wang et al., 2023a), ensuring that annotations accurately reflect the video content. In some studies, clips missing any modalities—vision, audio, or subtitles—are excluded to meet the requirements for comprehensive understanding across all modalities (Chen et al., 2024b).
3.2.2. Deduplications
Previous work has found there are great amount of duplicate data in various training datasets. For example, the BOOK Corpus (Zhu et al., 2015) contains thousands of duplicated books (Bandy and Vincent, 2021). The commonly used curated web crawl dataset C4 (Raffel et al., 2020) contains a single 61 word English sentence that repeated more than 6 hundred thousands times (Lee et al., 2022). More repeated data in training datasets can increase the rate of emitting memorized training data verbatim (Carlini et al., 2022; Kandpal et al., 2022), and deduplication of training datasets can prevent the memorization problem, thus alleviating privacy concerns (Kandpal et al., 2022). Moreover, duplication of training data can also cause the performance degradation of the pre-trained models (Hernandez et al., 2022). Training on deduplicated datasets can save the training cost while does not hurt the model perplexity (Lee et al., 2022).
Existing deduplication methods consider exact duplication, approximate duplication and semantic duplication in sentence-level (or sequence-level) and document-level. Exact deduplication is the most simple way to remove duplication. For sentence or sequence level, exact deduplication consider to remove exact string matching between sentence (or sequence) (Raffel et al., 2020; Suárez et al., 2019). To improve computational and memory efficiency, Suffix Arrays (Manber and Myers, 1993; Lee et al., 2022) and Bloom Filters (Computer, 2023; Soldaini et al., 2024) are used to achieve parallelized linear time complexity. For document level exact deduplication, URL deduplication is considered to remove exactly same web pages (Penedo et al., 2023). Approximate deduplication methods mainly focus on document-level duplicates. Usually, Locality Sensitive Hashing (LSH)-based MinHash (Broder, 1997) or SimHash (Charikar, 2002) methods are used to remove approximately duplicate documents (Computer, 2023; Lee et al., 2022; Penedo et al., 2023). These methods can achieve document-level deduplication with linear time and space complexity of the document number and can be implemented in a highly distributed setting (Lee et al., 2022). Apart from hashing-based approximate deduplication, recently there are some work consider to leverage pre-trained foundation models as semantic embedding metric for document-level embedding (Silcock et al., 2022; Kaddour, 2023; Abbas et al., 2023; Tirumala et al., 2023). These Semantic deduplication methods consider using Sentence-BERT (Reimers and Gurevych, 2019) MPNET (Song et al., 2020) for embedding (Silcock et al., 2022), E5-Large (Wang et al., 2022) for embedding (Kaddour, 2023), and OPT-125M (Zhang et al., 2022) for embedding (Abbas et al., 2023; Tirumala et al., 2023). These semantic deduplication processes are usually after the exact deduplication and approximate deduplication processes, removing the semantic duplicates by clustering the embedding points and keeping representative data in each clusters. However, how to choose an appropriate pre-trained foundation models and whether this models can embed semantic information efficiently are still open questions.
For image-text pairs or interleaved image-text documents, common methods of image deduplication include using the image URL (Zhu et al., 2024) or employing pHash (Zauner, 2010) algorithms. It’s worth noting that in the realm of interleaved image-text documents, there is currently no universally accepted, reliable method for deduplication based on both image and text elements simultaneously.
For Video multimodal data, video fingerprinting technology is uniquely applied to videos to identify and remove exact duplicates, addressing the challenge of video reuploads and mirrored content, which is more prevalent and complex in videos than in static images(Xu et al., 2023c).
3.2.3. Data Enhancement
Data enhancement of multimodal datasets usually focus on two aspects, enhancing the X modality data, and enhancing the text data. Traditional data augmentation methods for single modality data have been discussed thoroughly in previous work (Xu et al., 2023d; Cauli and Reforgiato Recupero, 2022; Ko et al., 2015; Wei et al., 2020; Zhao et al., 2021; Feng et al., 2021). In this work, we only consider data enhancement methods for MLLMs datasets.
For vision-language models, enhancing the quality of image-caption datasets is crucial for training MLLMs. Improving the captions not only enhances their alignment with the images but also prevents the discarding of highly informative images due to poor text quality. Rewriting text using BLIP2 has been shown to effectively improve training results (Nguyen et al., 2023). However, for large-scale image-text datasets (approximately 1.28 billion), this method experiences diminishing returns in terms of ImageNet accuracy, possibly due to the lack of diversity in the generated captions (Nguyen et al., 2023). Approaches like LaCLIP (Fan et al., 2024) and VeCLIP (Lai et al., 2023a) employ LLMs with carefully designed prompts to generate more diverse captions. ShareGPT4V (Chen et al., 2023e) employed the 100K high-quality captions generated by GPT4-Vision to fine-tune an alternative caption model and named it as Share-Captioner. The Share-Captioner is capable of generating highly content-related descriptions with unified instruction for the pre-train dataset. MLM (Wang et al., 2024c) used GPT-4 or GPT-4V to constructing multimodal instruction tuning data on proposed quality scoring tasks to fine-tune MLM to realize accurate quality assessment. Then they adopt the fine-tuned MLM Filter to generate quality scores for each data point in the data pool and then select the high-quality data.
Improving image resolution can enhance the performance of MLLMs. Initially, MLLMs primarily processed fixed, lower-resolution inputs, typically around 224 pixels (Liu et al., 2024b; Chen et al., 2023j; Zhu et al., 2023). Recent models such as LLaVA-1.5 (Liu et al., 2023c) and BLiVA (Hu et al., 2024) have improved performance by increasing the input resolution to 336 pixels and integrating task-specific global features. Furthermore, models like Qwen-VL (Bai et al., 2023a) and OtterHD (Li et al., 2023m) have pushed resolution support to 448 pixels, incorporating fine-tuning in the visual encoder during training, while maintaining the original image size during inference, leading to more precise segmentation recognition. Notably, Monkey (Li et al., 2023j) has significantly raised the resolution to 896 pixels by employing multiple visual encoders and leveraging fine-tuning techniques from Qwen-VL. Models like LLAVA HR (Luo et al., 2024) and Vary (Wei et al., 2023) have introduced additional visual encoders to capture more complex features, requiring extensive pre-training tasks. LLAVA UHD (Xu et al., 2024) and Ureader (Ye et al., 2023a) have improved the model’s capacity to understand detailed image features through adaptive region segmentation. However, a fundamental challenge remains within the MLLM architecture: models using lower-resolution inputs struggle to detect fine details, whereas those with higher resolutions may underperform in tasks requiring a broader global understanding.
3.3. Commonly-used Datasets for MLLMs
In this section we will briefly introduce commonly used multimodal datasets from different modality. These datasets are not utilized for this research. The comprehensive summary of text-only datasets has been elaborately discussed in previous literature (Liu et al., 2024a); therefore, this part will not be extensively covered within the scope of this discussion.
3.3.1. Image Datasets
For vision langauge models, the commonly used image-text datasets can be categorized into three distinct types: image-caption datasets, interleaved image-text datasets, and visual question answering datasets.
Image-Caption Datasets.
Image-caption pairs datasets are the most commonly used datasets in the pre-training stage of MLLMs. By leveraging the captions of each image, one can align the image modality with the text modality, enabling LLMs to understand image information.
General image-caption datasets contain images with short captions, typically one sentence or a few words, that describe the key features of the image. Notably, the LAION series (comprising LAION-400M (Schuhmann et al., 2021) and LAION-5B (Schuhmann et al., 2022)) represents some of the largest compilations, aggregating billions of pairs from the CommonCrawl. These datasets emphasize the maintenance of high quality through meticulous filtering based on the relevance of image-text pairs and the exclusion of inappropriate content. Derived from the LAION-5B subset, the LAION-COCO (Schuhmann et al., 2022) dataset is specifically designed to explore the impact of synthetic captions on model training. Also curated from CommonCrawl project, the COYO-700M (Minwoo Byeon, 2022) dataset emphasizes the informative correlation between images and alt-texts. The DataComp (Gadre et al., 2024) challenge introduces COMMONPOOL, a substantial multimodal dataset constructed from 12.8 billion image-text pairs that encourages innovation in dataset design. Apart from CommonCrawl project, some self-crawled datasets also play a vital role. Several datasets are curated from images with alt attributes from webpages. For example, Conceptual Captions (CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021)) use generalized alt-text from the web and are processed with multiple filters, containing 3.3 million and 12 million image-text pairs respectively. Also leveraging alt-text for images on webpages, ALT200M (Hu et al., 2022) dataset collects 200 million images with their alt attributes, aiming for understanding scaling of vision-language models. For multilingual image-caption datasets, a representative one is the Wukong (Gu et al., 2022) dataset, which addresses the scarcity of large-scale datasets in the Chinese language, offering 100 million quality-assured pairs that are instrumental for the development of Chinese vision-language pre-training models. Larger dataset such as Long text & image pairs (LTIP) (Alayrac et al., 2022) dataset provides a unique collection of 312 million images with lengthier textual descriptions, enhancing models like Flamingo that are tailored for complex multimodal tasks.
Content descriptive image-caption datasets contain longer caption with more descriptive information. Usually each image in content description image-caption datasets contain at least 5 sentences of captions. Content descriptive image-caption datasets are essential for advancing the interface between visual content and text descriptions, enhancing the ability of MLLMs to interpret and generate more detailed description from images. For example, MS-COCO (Chen et al., 2015) employs images with multiple reference captions to refine evaluation metrics, while Flickr30K (Young et al., 2014) extends the variety of descriptions to bolster semantic inference capabilities. In addition, datasets like Visual Genome (Krishna et al., 2017) provide comprehensive annotations beyond simple captions, including objects, attributes, and relationships, to facilitate complex scene understanding. While the AI Challenger Captions (Wu et al., 2017) dataset specifically addresses the need for non-English language representation in image captioning, offering extensive annotations in Chinese to bridge the semantic gap between low-level visual features and high-level conceptual descriptions. Narrative and textual comprehension are further introduced by datasets like VIST (Huang et al., 2016), which emphasizes storytelling through visual sequences. TextCaps (Sidorov et al., 2020) incorporates reading comprehension into captioning. Collectively, these datasets not only propel advancements in automated image captioning but also tackle broader challenges in AI’s capability to process and generate meaningful visual and textual data, significantly enhancing context-aware machine understanding.
Interleaved Image-Text Documents.
A fundamental contrast between image-text caption datasets and interleaved image-text document datasets lies in their composition: image-caption pair datasets typically consist of a single image accompanied by multiple closely related captions, whereas interleaved image-text datasets consist of a text document interspersed with several illustrative images. The correlation between the images and the accompanying text tends to be relatively lower, but these datasets usually provides more semantic information, which is crucial for maintaining models’ semantic understanding. M3W (Alayrac et al., 2022) leverages data from 43 million web pages, integrating images with text to train models in few-shot learning scenarios. MMC4 (Zhu et al., 2024), an extension of the C4 corpus, aligns images with texts using CLIP features, resulting in a dataset containing over 101.2 million documents. OBELICS (Laurençon et al., 2024), sourced from CommonCrawl, comprises 141 million web pages. OmniCorpus (Li et al., 2024) scales the interleaved image-text datasets to 10 billion-level, providing much richer image and text information from diverse sources.
Visual Question Answer (VQA) Datasets
Visual question answer (VQA) datasets facilitate advanced research in MLLMs by allowing models to interpret images and answer related questions. These datasets typically consist of images paired with corresponding questions and answers, aiming to develop and evaluate models’ ability to understand visual content and provide accurate responses. For example, VQAv2.0 (Goyal et al., 2019) expands on its forerunner by introducing complementary images that prompt different answers to the same question, effectively addressing biases and increasing the dataset’s diversity. Visual-7W (Zhu et al., 2016) extends the scope by incorporating questions across multiple dimensions—what, where, when, who, why, how, and which—linked to specific objects within images. ST-VQA (Biten et al., 2019) integrates scene text into the VQA framework, making it possible to answer text-based questions from visual data. Shikra-RD (Chen et al., 2023j) leverages advanced language models to annotate images with relational descriptions, enhancing image comprehension. OCR-VQA focuses on reading text from book covers to answer related questions, combining elements of optical character recognition and VQA. DocVQA (Mathew et al., 2021) targets document images for extractive question answering, emphasizing precision in answers derived from visible text. A-OKVQA (Marino et al., 2019) introduces questions that demand commonsense and world knowledge, pushing the envelope on reasoning capabilities required from AI systems. TextVQA (Singh et al., 2019) specifically challenges models to read and understand text within images to respond accurately, and GQA (Hudson and Manning, 2019) promotes advanced reasoning over detailed visual scenes annotated with comprehensive scene graphs. These datasets collectively advance the field by enabling more nuanced interactions between AI models and the rich content within images, aiming for greater depth in the understanding and contextual integration of visual and textual data.
3.3.2. Video Datasets
Video datasets are comprehensive collections of video clips accompanied by associated annotations or labels. These datasets are meticulously designed to facilitate the training and evaluation of models for a wide range of video-related tasks, including action recognition, various video-text tasks, as well as video-centric dialogue.
Commonly used video-text dataset such as the MSR-VTT (Bain et al., 2021) features 10,000 video clips totaling 41.2 hours, each annotated with multiple descriptive sentences. Similarly, the WebVid series offers millions of web-sourced video clips with accompanying captions, expanding from 2.5 million pairs in WebVid-2M to 10 million pairs in WebVid-10M (Bain et al., 2021), covering a vast array of 13,000 hours of video content. The VTP (Alayrac et al., 2022) dataset further enriches the field with 27 million short video-text pairs sourced from a select few high-quality websites. Moreover, the recently introduced Panda-70M (Chen et al., 2024c) dataset contains a massive 70 million high-quality video-caption pairs, designed to enhance the training of high-performance MLLMs. Collectively, these datasets are integral to advancing the capabilities of MLLMs in understanding and interacting with video content. InternVid (Wang et al., 2023a) scales video-text dataset to 230M annotated video-text pairs, with every video lasting 351.9s on average.
3.3.3. Audio Datasets
Audio datasets are curated collections of sound recordings along with associated annotations or labels. These datasets are specifically designed to support the training and evaluation of models for various audio-related tasks, such as speech recognition, sound event detection, music classification, and speaker identification. Audio datasets vary significantly in size, quality, and scope, with some focusing on specific languages or acoustic environments, while others strive for diversity to ensure generalizability across different audio processing scenarios. Notable examples include AISHELL-2 (Du et al., 2018), a Mandarin Chinese speech corpus with over 1000 hours of data from multiple regions; WavCaps (Mei et al., 2023), an extensive English audio captioning dataset featuring approximately 400,000 clips; and VSDial-CN (Chen et al., 2023b), a multi-modal dataset derived from VisDial, encompassing visual data alongside related dialogues and captions, tailored for Automatic Speech Recognition (ASR) systems.
3.3.4. 3D Datasets
The development of MLLMs benefits from the utilization of diverse 3D datasets, which provide comprehensive environmental and structural data critical for tasks like scene understanding and semantic segmentation. Notably, the ScanNet (Dai et al., 2017) dataset offers an extensive collection of RGB-D video data across 1,513 indoor scenes, annotated with 3D camera poses, surface reconstructions, and semantic segmentations, totaling over 2.5 million views. Similarly, the S3DIS (Armeni et al., 2017) dataset includes point clouds from six large indoor areas, encompassing 271 rooms, with detailed semantic annotations for each point. The Structured3D (Zheng et al., 2020) dataset provides a vast repository of 3,500 home designs, encompassing 21,835 rooms detailed with elements such as object geometry, materials, and textures, tailored for analyses in 3D reconstruction and interior design.
4. Data-centric pre-training
The pre-training stage of multimodal large language models (MLLMs) is crucial for developing models that can effectively process and generate information across multiple modalities. This stage can be divided into two distinct phases, each focusing on specific aspects of the model’s architecture and training objectives.
The first phase involves two parallel processes: pre-training the LLM backbone (such as Vicuna (Chiang et al., 2023) and LLaMA2 (Touvron et al., 2023)) using text-only datasets, and pre-training the modality encoders (such as ViT (Dosovitskiy et al., 2020) and CLIP-ViT (Radford et al., 2021) for visual encoding, C-Former (Chen et al., 2023b) and HuBERT (Hsu et al., 2021) for audio encoding, and ULIP-2 (Xue et al., 2022, 2023) for 3D point cloud encoding) using pairs of data from different modalities, such as image-text or video-text pairs. This phase aims to establish a strong foundation for the model’s understanding of both textual and non-textual information.
The second phase builds upon the knowledge acquired in the first phase by further training the LLM and modality encoders using a mixture of multimodal data. During this phase, the input projector (such as a linear projector, cross-attention mechanism, Q-Former (Li et al., 2022a), or P-Former (Jian et al., 2024)) is trained to effectively map the features extracted by the modality encoders into the LLM’s embedding space. This mapping allows for a unified representation of textual and non-textual information within the model’s architecture. Some researchers consider employing a selective training approach during the second phase, which involves keeping certain components of the model frozen while training specific parts. This approach helps to preserve the knowledge acquired during the first phase, allows for targeted adaptation to specific multimodal tasks, and reduces computational costs.
4.1. Domain Mixture
The performance of a language model (LM) is significantly influenced by the composition of its pre-training data, which often includes sources such as Wikipedia, books, and web text (Xie et al., 2024a, b).
Previous methods typically select domain weights heuristically or optimize them using downstream tasks (Du et al., 2022). However, these approaches can be suboptimal or costly, requiring the training of LMs for different sets of domain weights and potentially leading to overfitting to specific downstream tasks. DoReMi (Xie et al., 2024a) offers a solution by optimizing domain weights using proxy models that are 30 times smaller than the target LLM, without requiring knowledge of downstream tasks. Recent work (Liu et al., 2024e) formulate domain mixture problem as a regression problem. They assume the rank invariance of domain mixture between small models and large models, thus developing domain mixture regression function to predict the optimal domain mixture for model training. There are also some work discuss this problem through domain mixture scaling law. These work extend the original scaling law (Kaplan et al., 2020; Hoffmann et al., 2022) to domain data mixing law (Que et al., 2024; Ye et al., 2024; Ge et al., 2024), discussing the relationship between domain mixture rate and loss through scaling models.
In the realm of video understanding, there is a consensus on the necessity for diverse datasets (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a). Xu et al. (2023c) highlight the use of hierarchical multi-label classification models (Giunchiglia and Lukasiewicz, 2020) to ensure dataset balance and comprehensiveness across multiple dimensions, thereby optimizing the model’s ability to understand and process various video domains, which is instrumental in training more versatile applicable video-language models.
4.2. Modality Mixture
When pre-training MLLMs, determining the optimal proportions of multimodal data is crucial for enhancing the model’s performance across different tasks. MM1 (McKinzie et al., 2024) explores the balance between image-caption pairs, interleaved image-caption documents, and text-only data for vision-language pre-training. They found that a ratio of 5:5:1 for caption/interleaved/text data yields the best overall performance on text-only tasks as well as zero-shot and few-shot image-text tasks. Additionally, incorporating synthetic data (Lai et al., 2023b) improves the model’s few-shot learning capabilities for image-text tasks.
For the development of MLLMs tailored for video understanding, the strategic integration of both video-text and image-text pairs during the pre-training phase is pivotal. This approach, highlighted in recent research (Lin et al., 2023b; Zhang et al., 2023; Jin et al., 2023; Han et al., 2023; Li et al., 2023c; Wang et al., 2023a), enhances the model’s capabilities by leveraging the complementary nature of these data types. Integrating labeled image datasets like COCO with video data broadens the variety of training datasets and addresses the scarcity of high-quality video-text resources. Studies (Lin et al., 2023b; Jin et al., 2023) indicate that training on both images and videos enhances models’ ability to understand static and dynamic visual information without needing task-specific adaptations. Adding a temporal modeling module to the vision encoder bridges the temporal dynamics of videos with the static nature of images, fostering cohesive visual understanding across different types of visual media. However, treating images as single-frame videos (Zhang et al., 2023; Luo et al., 2023) or pseudo-videos may weaken the model’s ability to grasp the temporal aspects of video sequences. An imbalanced mix could also bias the model towards one modality, complicating its learning process. Therefore, determining the optimal mix ratio of image-text pairs and video-text pairs is crucial for improving the model’s comprehensive understanding capabilities, ensuring a balanced approach to learning from both static images and dynamic videos. Additionally, video LLMs are increasingly being trained through multiple branches (Chen et al., 2024b; Zhang et al., 2023; Lyu et al., 2023), such as vision-language, audio-language, and subtitle-language. These methods leverage the correlations among different modalities to broaden the models’ understanding and learning capabilities. Recent studies (Shu et al., 2023; Chen et al., 2024b; Han et al., 2023) have shown that combining audio, subtitles, and visual data during training enhances performance across various video understanding benchmarks. This modular training approach also offers the flexibility to pre-train with partial data (Han et al., 2023; Sun et al., 2023c), such as visual-only or audio-only, particularly when comprehensive multimodal data is not available.
4.3. Quality Selection
Due to the diverse distribution of data, training a large model with all available data is not always optimal. Therefore, data selection becomes essential, offering benefits such as reduced training time and energy consumption (Gadre et al., 2024). Previous work has studied data selection methods using n-gram similarity with high quality datasets (Xie et al., 2023; Gao et al., 2020; Chowdhery et al., 2023), perplexity (Marion et al., 2023; Wenzek et al., 2020), influence functions (Park et al., 2023) and llm-based classifier (Wettig et al., 2024). Unlike pure text datasets, data selection for multimodal datasets must consider the alignment between different modalities. Data selection methods can be categorized into two types: active learning-based and pre-training selection. Active learning-based methods, such as CiT (Xu et al., 2023a), use data proxies to dynamically select training data during the training process, achieving significant acceleration effects. Pre-training selection methods, like Datacomp (Gadre et al., 2024) and Bunny (He et al., 2024), evaluate and select all data before training begins.
Regarding data selection criteria, methods can be distribution-agnostic, focusing solely on individual data point quality, or distribution-aware, considering the overall data distribution. Distribution-agnostic methods include training a model with the top 30% of data ranked by CLIP score, which can significantly improve results (Gadre et al., 2024). More comprehensive metrics have also been developed, showing better performance than using CLIP score alone (Wang et al., 2024c). For instance, Mahmoud et al. (2024) use the difference between original and synthetic captions generated by a small model to evaluate image-text pair alignment. Given the limited scope of purely distribution-agnostic work, combining distribution-agnostic and distribution-aware methods is often more effective. For example, using CLIP-score-based filters alongside image-based filters can outperform either method alone on large datasets (Wang et al., 2024a; Gadre et al., 2024).
5. data-centric adaptation
Adaptation is crucial for aligning pre-trained multimodal large language models (MLLMs) with specific tasks and user preferences. Self-supervised pre-training provides LLMs with a broad understanding of textual and multimodal information, while supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) are essential for adapting these models to excel in targeted applications and adhere to user preferences and societal values. Data-centric SFT involves training models on carefully curated datasets across multiple modalities. In contrast, data-centric RLHF focuses on collecting human judgments to guide the model toward generating desirable and ethically aligned responses. Emphasizing data-centric approaches highlights the importance of high-quality, domain-specific, and multimodal datasets in the successful adaptation of MLLMs.
5.1. Data-Centric Supervised Finetuning
Supervised fine-tuning (SFT) has become an indispensable technique for adapting MLLMs to specific domains and tasks. By leveraging carefully curated instruction-formatted datasets that encompass multimodal information, these models can be guided to acquire the necessary knowledge and abilities to excel in targeted applications.
During the SFT stage, various strategies have been explored, including fine-tuning one or all of the three primary pre-trained components: the LLM backbone, the modality encoders, and the input projector. This selective approach allows for a balance between adapting the model to specific tasks and preserving the generalized knowledge acquired during pre-training.
To facilitate effective supervised fine-tuning, diverse datasets that align with the target domain and task are used. These datasets typically include a combination of multimodal instruction-response pairs and text-only SFT data. By exposing the MLLMs to this rich variety of data, the model can develop a comprehensive understanding of the target domain and acquire the necessary skills to perform the desired tasks effectively.
The curation of self-supervised fine-tuning datasets usually includes four steps:
  • Collect X-text pairs from various data sources, similar to the pre-training stage as introduced in Section 3.1.
  • Process the data for further use, including filtering, deduplication, translation, etc.
  • Construct the processed X-text pairs into instruction-response form, which involves designing the instruction and converting captions into answers.
  • Select high-quality instruction-response pairs based on the fine-tuning needs. This step is not always necessary but can enhance the dataset’s quality.
5.1.1. SFT Data Collection and Processing
After collecting these data, several general approaches are employed to improve data quality, including filtering, deduplication, and data enhancement. These steps are similar to the data processing methods used in the pre-training stage described in Section 3.2.
In addition to these general data processing approaches, specific methods aim to improve data quality for supervised fine-tuning tasks. The main goal is to add more textual information to the datasets or improve the quality of captions. These methods, while akin to the data enhancement techniques mentioned in Section 3.2, are more focused on SFT data generation. For example, MiniGPT-4 (Zhu et al., 2023) curated detailed image description datasets for vision-language adaptation by leveraging their pre-trained model with instructions such as ”Describe this image in detail,” ensuring generated sentences contained more than 80 tokens. They also used ChatGPT to refine the descriptions and manually checked for quality, obtaining 3,500 high-quality image-text pairs. Similarly, ShareGPT4V (Chen et al., 2023e) used GPT-4 Vision to generate 100,000 high-quality captions that included world knowledge, spatial information, aesthetic evaluation, and more. These captions were directly used to generate SFT instruction-response pairs.
5.1.2. Instruction-Response Pairs Generation
After collecting and processing the original X-Text datasets, the next step is to generate the instruction-response pairs datasets based on the original datasets. The supervised fine-tuning stage is typically aiming to enhance the model’s ability on downstream tasks. For this reason, previous work consider to generate instruction-response pairs for different downstream tasks, including descriptive tasks, question-answer tasks, reasoning tasks, and classification tasks. Based on task categorization, previous work selects different source datasets and uses various methods to generate the instruction-response pairs. In this section, we first introduce the differences between each downstream task, and then we summarize representative methods of generating instruction-response datasets for each type of downstream task.
Captioning Instruction-Response Datasets.
The simplest tasks are the captioning ones. For this type of task, images (or other modality data) are usually provided, and the MLLMs are asked to give out description of the data. The instruction-response datasets designed to enhance the model’s captioning ability are relatively easy to construct. These kinds of instructions are usually one question asking for the description of the X modality data. Specifically, for image data, MiniGPT-4 (Zhu et al., 2023) construct description instruction tuning by randomly sample instructions such as ”Describe this image in detail”, and their modified high-quality captions described in Section. 5.1.1 are directly used as the response. While LLaVA (Xu et al., 2024) considers designing two types of description instructions, brief description and detailed description and uses GPT-4 (OpenAI, 2023) to generate the answers. For brief descriptions, they add keywords such as ”concisely”, ”brief”, and ”short” in their questions, and for detailed descriptions, they add keywords such as ”descriptive” and ”comprehensive”. ShareGPT4V (Chen et al., 2023e) replaces the detailed description data in LLaVA with its GPT4-Vision generated high-quality captions. X-InstructBLIP (Panagopoulou et al., 2023) further uses these descriptive prompts not only on image-text datasets but also on audio-text, video-text, and 3D-text datasets as a description instruction.
Question Answer Instruction-Response Datasets.
Question answer task requires the model to answer questions based on the given information on the X modality. This kind of task includes simple questions answers, and multiple choices. The Most well-known question-answer task is the visual question answer (VQA) task for visual modality. For VQA tasks, previous research on computer vision has developed a large amount of VQA datasets, and recent work on MLLMs leverages these VQA datasets for constructing instruction-response datasets. For example, LLaVA-1.5 (Liu et al., 2023c) leverage GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), OKVQA (Marino et al., 2019) and VQAv2 (Goyal et al., 2019) to enrich their instruction response datasets. They format these datasets in a uniform template as ”Answer the question using a single word or phrase.” TextVQA (Singh et al., 2019) is another source used for constructing instruction-response datasets which contain text information in images. M3IT (Li et al., 2023k) utilize VQA datasets, such as VQA-v2 (Goyal et al., 2019), Shapes VQA (Andreas et al., 2016) and DocVQA (Mathew et al., 2021), to develop a multimodal instruction tuning dataset that contains 2.4 million instances. In addition to visual question-answer tasks, instruction tuning to improve model performance in the video question answer task also leverages previous video question answer datasets, such as MSRVTT-QA (Xu et al., 2017), iVQA (Yang et al., 2021), MSVD-QA (Xu et al., 2017), and ActivityNet-QA (Yu et al., 2019).
Reasoning Instruction-Response Datasets
Shikra-RD (Chen et al., 2023j) focuses on enhancing the models’ ability for referential dialogue problems, i.e., users point to specific areas and ask questions. They collect public VQA datasets, image captioning data, and several datasets with existing positional annotations to build instruction tuning datasets. Leveraging captions from the Flickr30K (Young et al., 2014) image-caption dataset and using GPT-4, they obtain high-quality RD annotations. PVIT (Chen et al., 2023h) converts the traditional VQA datasets such as GQA and OCR into instruction-response datasets with region-level information. MiniGPT-V2 (Chen et al., 2023m) designs a multi-task instruction template, following the conversation template from LLaMA-2 (Li et al., 2023f). They add task identifier tokens and spatial location representations to identify the task and spatial information in the image. They build five types of instruction-response datasets, including LLaVA-Instruct-150K, grounded image captions from Flickr30K (Young et al., 2014) with direct instruction and object parsing, multi-round conversations, and text-only unnatural instruction datasets (Honovich et al., 2022b). CogVLM (Wang et al., 2024b) provides an IT dataset called CogVLM-SFT-311K. To construct this dataset, they manually select high-quality IT data from MiniGPT-4 and integrate it with LLaVA-Instruct-150K, translating it into Chinese. They then manually correct any noise in their collected instruction tuning datasets and translate them back into English. In their model training, they also leverage VQA datasets and bounding box instruction tuning datasets generated from four types of datasets: grounded captioning (GC) datasets with box-noun phrase pairs in each image, referring expression generation (REG) datasets with box and explanation, referring expression comprehension (REC) datasets with boxes annotated for text, and grounded visual question answering (GroundedVQA) datasets. LLaVA1.5 (Liu et al., 2023c) leverage region level VQA datasets such as Visual Genome (Krishna et al., 2017) and RefCOCO (Kazemzadeh et al., 2014; Mao et al., 2016) in their IT datasets construction.
Other Instruction-Response Datasets
Another type of downstream task is the classification task, where the data is classified into either a label from a candidate pool or an open-label. M3IT (Li et al., 2023k) leverages various datasets for image classification, including ImageNet (Russakovsky et al., 2015), Grounded Object Identification (COCO-GOI)(Lin et al., 2014), COCO-Text(Veit et al., 2016), Image Text Matching (COCO-ITM)(Lin et al., 2014), e-SNLI-VE(Kayser et al., 2021), Multi-modal Fact Checking (Mocheg)(Yao et al., 2023), and IQA(Duanmu et al., 2021). Additionally, for image classification tasks, InstructBLIP (Dai et al., 2024) uses HatefulMemes (Kiela et al., 2020), a binary classification dataset for detecting hateful content in memes, to generate an instruction-response dataset. X-InstructBLIP (Panagopoulou et al., 2023) further extends this approach to audio classification by using AudioSet (Gemmeke et al., 2017). The instructions used are sentences asking for the classification of the audio, such as ”Classify the following audio”. And the response is the class of the original data.
5.1.3. SFT Data Selection
Supervised fine-tuning is crucial for adapting LLMs with downstream applications. Zhou et al. (2024) found that the primary knowledge of LLMs is acquired during the pre-training stage, and the purpose of instruction tuning is to enable LLMs to learn how to perform well on specific tasks and interact with humans. They determined that only a small set of carefully crafted, high-quality instructions is sufficient to endow LLMs with powerful instruction-following capabilities. Subsequently, various methods have been proposed for data selection to identify high-quality data, improving performance and reducing training costs. These methods are categorized into four types: coreset-based, LLMs-based, gradient-based, and self-instruction-based methods.
Coreset-Based Methods
Coreset provides a compact representation of a larger dataset while preserving its essential characteristics. From a geometric perspective, similar data points in the feature space are close to each other. For example, Sener and Savarese (2017) employs the greedy k-center algorithm to select a coreset and applies it to CNN image classification. In addressing the multi-class classification task, MODERATE CORESET (Xia et al., 2022) scores data points based on their distance to the class center. Data points with scores close to the score median are selected as the coreset, which achieves a balance between discriminative, compressible, and diverse. SIMILAR (Kothawade et al., 2021) picks data points according to the submodular information measures (SMI). A properly selected SMI function can maintain the diversity of selected data and handle imbalance classes, out-of-distribution data, and redundancy. Some algorithms select data points based on model performance, focusing on ”important” samples. For example, Bachem et al. (2017) perform importance sampling based on the upper bound of the sensitivity score to generate a coreset. Generally speaking, data points with high costs are more likely to be selected. Other mertics of performance include GraNd (Paul et al., 2021), least confidence (Wang and Shang, 2014), and etc.
Given that Coreset techniques have demonstrated promising results both theoretically and empirically, it’s feasible to apply these algorithms to the data used for training LLMs. Chen et al. (2023k) employ the K-greedy algorithm to identify just 0.5% of the core data for fine-tuning a pre-trained language model, achieving performance only 1-2% lower than that obtained using the entire dataset. Similarly, the approach proposed by Das and Khetan (2023) involves selecting data that represents the easiest and most challenging examples, a strategy proven effective in (Sorscher et al., 2022). By utilizing only 32.5% of carefully selected data, they achieved state-of-the-art results.
LLMs-Based Methods
External model-based methods frequently utilize external models to evaluate the quality, diversity, and complexity of data. These models serve various functions, from scoring data to enhancing its characteristics. For instance, Du et al. (2023) leverage DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023f) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023e) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023c) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster. Zheng et al. (2024) and Sun et al. (2024) leverage LLMs to automatically select high-quality domain data, achieving superior performance. For multimodal instruction tuning datasets, MLLMs are typically used to obtain or select high-quality instruction tuning data. LLaVA1.5 (Liu et al., 2023c) pioneers the use of text-only GPT-4 to expand the COCO (Lin et al., 2014) bounding box and caption dataset into a multimodal instruction-following dataset. Following LLaVA, ShareGPT-4v (Chen et al., 2023e) utilizes GPT-4v to enhance a portion of LLaVA’s instruction tuning data. For Video Large Language Models (VideoLLMs), Liang et al. (2024) pioneered the use of CLIP score for video keyframe data selection, achieving SoTA performance.
Gradient-Based Methods
Neural Networks are typically trained through gradient-based optimization techniques. Consequently, researchers have been exploring gradient-based methods to use a subset of data to approximate the gradient effectively. The EL2U score was introduced as a method to evaluate how removing a single data point affects the gradient (Paul et al., 2021). However, the EL2U score is limited to single-task learning scenarios. To address multi-task learning, particularly in NLP tasks, another approach has been developed (Attendu and Corbeil, 2023). This method applies the EL2U score to each individual task and then combines these scores using the L2 norm to accommodate the complexities of multi-task learning. Another notable work is (Xia et al., 2024), which adapts existing influence formulations to work with the Adam optimizer. It only uses a few steps of gradients to compute gradient features efficiently and then stores the compressed features in a gradient datastore for efficient data selection.
Self-Instruction-Based Methods
For self-instruction-based methods, the evaluation and selection of data do not require the involvement of any external models. Li et al. (2023l) proposes a self-instruction method for LLMs to autonomously identify and select challenging data. Li et al. (2023d) identifies high-quality data as those that, when included as part of a few-shot example, can enhance the model’s performance. Kung et al. (2023) proposes a novel task-level uncertainty metric that measures the sensitivity of LLMs to instruction perturbations for a task, to identify high-quality data, and combines this with the concept of active learning to iteratively select data. Liu et al. (2024c) scores data at three granular levels: token, sentence, and model. Each level’s scoring is calculated based on the previous level, with the model-level score serving as the final score.
5.2. Data-Centric Human Preference Alignment
To align human preferences with language models, previous work has considered constructing instruction-response datasets that respect human preferences. These datasets usually involve three main elements. RLHF aims to use reinforcement learning (RL) to align the model with human preferences (Ouyang et al., 2022). After training the LLM, RLHF collects human feedback to rank Q&A text data, trains a reward model (RM), and then uses RL to fine-tune the LLMs.
The most impactful work of RLHF is InstructGPT (Ouyang et al., 2022). To collect human feedback, InstructGPT first designs prompts manually and from users who use the InstructGPT playgrounds. Then, they process the prompts using heuristic de-duplication methods, such as checking for prompts that share a long common prefix and limiting the number of prompts to 200 per user ID, to protect the user’s sensitive information and improve the quality of prompts.
These prompts encompass multiple tasks such as text generation, question answering, dialogue, summarization, and more. Human taggers are then asked to rank the responses to each prompt, constructing a Q&A dataset used to train the reward model (RM). Subsequently, the fine-tuning stage of the LLM is framed as a reinforcement learning (RL) problem using the trained reward model. In LLaMa2 (Li et al., 2023f), human preferences are further divided into fine-grained aspects, with two separate reward models trained to rank helpfulness and safety, respectively. Further work also aims to improve the models’ generation abilities from these aspects.
For MLLMs, several works have focused on aligning models with human preferences. Based on LLaVA (Liu et al., 2023b), LLaVA-RLHF (Sun et al., 2023a) designed a reinforcement learning from human feedback (RLHF) method to align LLaVA with human feedback. DRESS (Chen et al., 2023i) aims to improve models’ response quality from the perspective of human values, particularly the 3H criteria (Ouyang et al., 2022) (helpfulness, honesty, and harmlessness). They provide two datasets: the Large Vision Language Model with Natural Language Feedback (LVLM_NLF) dataset and the Vision-Language Safety (VLSafe) dataset. To construct the VLSafe dataset, they adopt the LLM-Human-in-the-Loop approach, iteratively creating and filtering the data based on the COCO dataset. The LLM they used is GPT-3.5 Turbo. The final dataset contains 5,874 samples, meeting most requirements for harmlessness alignment and evaluation.
6. evaluation
To offer a comprehensive data-centric viewpoint for assessing MLLMs, this section begins with two key aspects. First, we review widely-used methods for evaluating data quality. Then, we provide an overview of MLLMs evaluation datasets, highlighting their collection, processing, and characteristics.
6.1. Data Evaluation
Evaluating datasets is essential to ensure their quality and reliability. This process helps identify and correct errors and biases, enhancing the accuracy of model training and predictions. In this section, we summarize recent dataset evaluation metrics from the following perspectives.
6.1.1. Dataset Diversity
Dataset diversity is crucial for developing robust and generalizable machine learning models. A diverse dataset enables the model to manage various scenarios and minimizes the risk of bias. Some studies (Heusel et al., 2017; Sajjadi et al., 2018) require a reference distribution or a dataset. For instance, Heusel et al. (2017) measures the Wasserstein-2 distance between two Gaussian distributions: one fitted to the embeddings of the reference sample and the other to the embeddings of the sample being evaluated for diversity. Other work use similarity scores to define diversity. For example, Shen et al. (2019); Fomicheva et al. (2020) utilize the average pairwise similarity score or its complement, the average dissimilarity. Fomicheva et al. (2020) propose a two-metric evaluation paradigm using precision and recall, with precision measuring quality and recall assessing diversity in terms of coverage of the reference distribution. The Vendi score (Friedman and Dieng, 2023; Pasarkar and Dieng, 2023; YEH et al., 2023) stands out from existing diversity evaluation metrics as a reference-free, flexible, and interpretable metric that measures internal diversity without comparing it to a reference distribution. Its reliance on a user-defined similarity function allows for broad applicability across domains, incorporating correlations between features while being computationally efficient and unsupervised, which makes it a valuable tool for diverse machine learning applications. Oppositely, Lee et al. (2023) use diversity score called Task2Vec diversity coefficient (Miranda et al., 2022) to evaluate the diversity of publicly available datasets.
6.1.2. Dataset Quality
Dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enables models to learn effectively and make precise predictions. Ensuring dataset quality is essential for the accuracy and reliability of machine learning models. High-quality data enable effective learning and precise predictions. There are several approaches to evaluate data quality. TRUE (Honovich et al., 2022a) offers a thorough evaluation of factual consistency metrics in grounded text generation systems. By standardizing datasets and introducing a meta-evaluation protocol, it highlights the robust performance of large-scale NLI and QG-QA methods across different tasks. Object Hallucination (Rohrbach et al., 2018) introduces CHAIR (Caption Hallucination Assessment with Image Relevance), which measures the proportion of words in a generated caption that correspond to objects present in the image, using ground truth sentences and object segmentations. FAITH SCORE (Jing et al., 2023) assesses the faithfulness of generated answers from large vision-language models (LVLMs). It identifies descriptive statements, extracts atomic facts, and checks their consistency with input images. This structured process ensures that generated answers align closely with the visual content, providing a comprehensive evaluation of faithfulness in LVLM outputs.
6.1.3. Dataset Similarity
In machine learning theory, performance improves when the training data distribution closely matches the testing data distribution. Therefore, it is essential to evaluate the distances of data distribution both theoretically and empirically. Common metrics for measuring distribution similarity include Euclidean distance, KL division, CORAL loss (Sun and Saenko, 2016), Wasserstein distance, and MMD distance (Jiang et al., 2022). The MAUVE scores (Pillutla et al., 2023) serve as a comparison measure between pairs of distributions, such as those encountered in generative modeling of text or images. This metric provides statistical bounds and extensive experiments to demonstrate its effectiveness.
6.2. Evaluation Datasets for MLLMs
In this section, we will briefly summarize different types of evaluation datasets according to downstream tasks.
Common evaluation tasks for MLLMs can be divided into captioning tasks, question answering tasks, perception and reasoning tasks, and other noteworthy tasks such as classification tasks. Captioning datasets (Chen et al., 2015; Venugopalan et al., 2015; Kim et al., 2019) act as benchmarks to evaluate a model’s fundamental ability to understand information from different modalities. They can be used to assess the zero-shot or few-shot learning capabilities of MLLMs in multimodality tasks, offering insights into their ability to generalize and adapt to new scenarios with limited training examples. While evaluation datasets for question answering (QA) are crucial for assessing the performance of MLLMs that handle diverse data types like text, images, and audio (Goyal et al., 2019; Xu et al., 2017; Lipping et al., 2022). These datasets require models to analyze and combine information from multiple sources to respond to relevant questions accurately. Perception and reasoning are crucial tests for assessing the general abilities of machine learning language models (MLLM). These tasks requires model to reason for complicated logic and capture structured or spatial information (Singh et al., 2019; Masry et al., 2022; Liu et al., 2023b; Mangalam et al., 2024). There are also several noteworthy tasks not mentioned above, including classification task that require model to classify examples for right labels (Kiela et al., 2020; Piczak, 2015, 2015; Wu et al., 2015; Goyal et al., 2017), and multimodal dialog task that require for model’s in-context learning abilities to answer questions during multiple rounds of conversations (Bai et al., 2023b).
7. future direction
In this section, our attention is directed towards research avenues for MLLMs from a data-centric viewpoint. We categorize forthcoming directions based on the structure of our article. We begin by addressing future directions concerning data collecting and processing, followed by pre-training, adaptation, and evaluation aspects.
7.1. Data Processing System for MLLMs.
The processing of data for MLLMs entails numerous intricate steps across various modalities, as outlined in Section 3.2. Recent efforts in preparing data for MLLM training involve processing data autonomously through specialized data pipelines and processing operators (Alayrac et al., 2022; Changpinyo et al., 2021; Bain et al., 2021; Du et al., 2018). While previous work has focused on designing data processing systems for LLMs, such as Data-Juicer (Chen et al., 2023d) and Oasis (Zhou et al., 2023), there remains a gap in the availability of data processing systems tailored specifically for MLLM data. Such systems have to be equipped to handle multi-modal data types, encompassing not only textual data but also images, videos, audio, and 3D data formats.
7.2. Data Quantity Analysis for MLLM Pre-training.
Large language models have shown specific properties such as Large language models exhibit distinct characteristics such as emergent abilities and scaling laws. Previous research has examined emergent phenomena concerning model scale, measured by factors like training compute and the number of model parameters, as well as in relation to evaluation metrics (Wei et al., 2022; Schaeffer et al., 2024). Scaling laws have been investigated concerning model size, data scale, and data quality. OpenAI initially explored the power-law relationship of pre-trained loss with respect to model size, dataset size, and training compute in neural language models (Kaplan et al., 2020). Subsequently, DeepMind introduced a new scaling law demonstrating the optimal allocation of compute resources (Hoffmann et al., 2022). Building on this, Goyal et al. (2024) investigated the trade-offs between dataset quality and quantity, while Ye et al. (2024) examined scaling laws associated with data mixtures. Furthermore, Zhang et al. (2024b) analyzed the impact of these properties, particularly focusing on text-only LLMs. However, there remains a gap in understanding how data quantity influences emergent abilities in MLLMs. Exploring the scaling laws of MLLMs, particularly regarding the evaluation of data quantity, represents an area that warrants further investigation.
7.3. Data Quality Analysis for MLLM Pre-training.
Adjusting the ratio and quantity of pre-training data holds significant potential for enhancing model performance. However, with the abundance of pre-training data available, there arises a critical necessity for efficient data selection algorithms. Evaluating data for large models typically demands the utilization of the models themselves, which can impose significant computational burdens. Hence, the development of Proxy Models becomes crucial, as they offer substantial reductions in computational costs. Moreover, Proxy Models prove instrumental in optimizing the mixing ratio of data from diverse sources. While several approaches like Doremi (Xie et al., 2024a) and Doge (Fan et al., 2023) have been proposed in this regard, there remains a need for the refinement of proxy models and further exploration into the relationship between proxy models and their original counterparts.
7.4. MLLM Data Evaluation.
As elaborated in Section 6.1, despite the introduction of various data evaluation metrics, there remains a notable absence of comprehensive metrics tailored specifically for evaluating multimodal data. The assessment of multimodal data presents heightened challenges owing to its diverse array of data types, often encompassing multiple tasks and modalities (Liu et al., 2023c; Chen et al., 2024c). The quality of such data can be impacted by the individual models corresponding to each modality and their alignments. Hence, it becomes imperative to scrutinize the quality of multimodal data from multiple dimensions. In addition to relying solely on human-defined features, enhancing data quality can be achieved by leveraging statistical metrics and methodologies, such as scrutinizing distributions. The notion of domain adaptation offers theoretical underpinnings for matching distributions (Jiang et al., 2022). Consequently, assessing distributions emerges as a promising avenue, both in terms of theoretical frameworks and empirical validation.
7.5. Data Quality Improving for MLLM Supervised Fine-Tuning.
Harnessing Instruction Tuning data can substantially bolster a model’s proficiency in adhering to instructions and enhancing its performance across specific tasks. Given the pivotal role of both quality and quantity in optimizing model performance, the exploration of this area emerges as a crucial research endeavor. The evaluation of data for large-scale models necessitates metrics that are model-agnostic, effectively capturing the distinct characteristics of the data. As highlighted in the preceding paragraph, there is a pressing need to devise metrics tailored specifically for instruction tuning data. An alternative pragmatic approach involves employing Large Language Models (LLMs) for data evaluation, leveraging the extensive knowledge accrued during their pre-training phase (Du et al., 2023; Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e; Lu et al., 2023c; Parkar et al., 2024). However, utilizing LLMs for data assessment proves to be cost-effective. Despite the success of GPT-based methods for automated data quality evaluation, such approaches often lack interpretability. Hence, further investigation is warranted to gain a deeper understanding of GPT’s efficacy and to delineate the boundaries of employing GPT for data evaluation.
7.6. MLLM Lifelong Learning.
MLLMs often learn from diverse data types sequentially across various training phases (Li et al., 2022a; Liu et al., 2024b, 2023c). Typically, this sequential training starts with initializing the model using pre-trained weights for each modality, followed by pre-training, instruction tuning, and concluding with reinforcement learning from human feedback. It is critical throughout these stages to prevent catastrophic forgetting, ensuring the model retains its foundational language capabilities while acquiring multimodal skills. Therefore, integrating lifelong learning strategies with MLLMs represents a significant and necessary research direction.
8. conclusion
In this survey, we review recent advancements in data-centric multimodal large language models (MLLMs), introducing key concepts, findings, and techniques essential for processing training data for these models. Specifically, our discussion focuses on three crucial aspects of data handling: pre-training, adaptation, and evaluation. For each aspect, we highlight pivotal techniques and insights for data processing that are critical for effectively training MLLMs. Additionally, we provide a comprehensive summary of available data resources for training these models and discuss strategies for utilizing machine learning to enhance the performance of MLLMs. This survey aims to encapsulate the most recent literature on data-centric MLLMs, serving as a valuable reference for both researchers and engineers in the field.